### v1

Abstract
This paper investigates regularized stochastic gradient descent (SGD) algorithms for estimating nonlinear operators from a Polish space to a separable Hilbert space. We assume that the regression operator lies in a vector-valued reproducing kernel Hilbert space induced by an operator-valued kernel. Two significant settings are considered: an online setting with polynomially decaying step sizes and regularization parameters, and a finite-horizon setting with constant step sizes and regularization parameters. We introduce regularity conditions on the structure and smoothness of the target operator and the input random variables. Under these conditions, we provide a dimension-free convergence analysis for the prediction and estimation errors, deriving both expectation and high-probability error bounds. Our analysis demonstrates that these convergence rates are nearly optimal. Furthermore, we present a new technique for deriving bounds with high probability for general SGD schemes, which also ensures almost-sure convergence. Finally, we discuss potential extensions to more general operator-valued kernels and the encoder-decoder framework.
Keywords: Nonlinear operator learning, Operator-valued kernel, Regularized stochastic gradient descent, Convergence analysis
1 Introduction
In this paper, we consider a general model abstracted from nonlinear operator learning problems:
$y=h^{\dagger}(x)+\epsilon.$ (1.1)
Here, $\mathcal{X}$ is a Polish space and $\left(\mathcal{Y},\langle\cdot,\cdot\rangle_{\mathcal{Y}},\|\cdot\|_{\mathcal{Y}}\right)$ is a separable Hilbert space. The pair $(x,y)$ satisfying (1.1) is a random variable taking values in $\mathcal{X}\times\mathcal{Y}$, distributed according to an unknown probability measure $\rho$. The operator $h^{\dagger}:\mathcal{X}\rightarrow\mathcal{Y}$ is a measurable (possibly nonlinear) mapping defined by the conditional expectation $h^{\dagger}(x):=\mathbb{E}[y|x]$. The noise term $\epsilon$ is a centered $\mathcal{Y}$-valued random variable, assumed to be independent of $x$ and to have finite variance, i.e., $\sigma^{2}:=\mathbb{E}[\|\epsilon\|^{2}_{\mathcal{Y}}]<\infty$.
The model (1.1) has been widely employed in surrogate approaches for structured output prediction [46, 21, 26, 7, 14, 4]. In practice, many applications involve inputs or outputs with explicit or implicit discrete structures. Examples of implicitly structured data used in predictions include text, images, and videos in document processing and retrieval, as well as genes and proteins in computational biology. To learn models that predict outputs with structured components, surrogate methods embed the structured output into a Hilbert space, which formulates the task as operator regression with an infinite-dimensional output space. During prediction, a decoding step maps the output from the Hilbert space back to the original structured output space. Structured prediction tasks such as image completion [46], label ranking [29], and graph prediction [6] can thus be addressed through operator learning using surrogate approaches. Another important application of model (1.1) is functional output regression [24, 28, 25, 41]. These problems have become increasingly relevant with the growing capacity to collect functional data, motivating a shift toward a functional perspective in modeling [25]. This has led to the development of the now-thriving field of operator learning [1, 30], which aims to approximate operators between Hilbert (or more generally, Banach) spaces using data. A prominent example is the learning of solution operators for partial differential equations (PDEs), where the goal is to approximate the operator that maps a parameter space—describing the physical and geometrical constraints of the PDE—to its solution space [33, 3, 36]. In this paper, we study operator learning algorithms designed to act directly on functions rather than on high-dimensional vectors. This functional perspective allows us to capture the intrinsic properties of the problem, avoiding reliance on specific discretizations or pixelizations.
We introduce a supervised learning framework based on model (1.1). Consider a data set $\{(x_{t},y_{t})\}_{t=1}^{T}$ generated by model (1.1), or equivalently, drawn independently from the distribution $\rho$. To estimate $h^{\dagger}$, we minimize the regularized functional $\mathcal{E}(h)+\lambda\|h\|_{\mathcal{H}}^{2}$ over all $h\in\mathcal{H}$, where $\mathcal{H}$ is some Hilbert space, $\mathcal{E}(h):=\mathbb{E}\left[\|h(x)-y\|^{2}_{\mathcal{Y}}\right]$ denotes the mean squared error, and $\lambda>0$ is a regularization parameter. In this paper, we adopt a non-parametric approach to solve the nonlinear model (1.1), assuming that $\mathcal{H}$ is a vector-valued reproducing kernel Hilbert space (RKHS) induced by an operator-valued kernel $K$ [28, 25, 7, 4].
To illustrate our algorithm, we introduce some notations along with basic concepts from operator theory [15]. Consider a linear operator $A:\mathcal{H}_{1}\to\mathcal{H}_{2}$, where both $(\mathcal{H}_{1},\langle\cdot,\cdot\rangle_{\mathcal{H}_{1}},\|\cdot\|_{\mathcal{H}_{1}})$ and $(\mathcal{H}_{2},\langle\cdot,\cdot\rangle_{\mathcal{H}_{2}},\|\cdot\|_{\mathcal{H}_{2}})$ are Hilbert spaces. The set of bounded linear operators from $\mathcal{H}_{1}$ to $\mathcal{H}_{2}$ forms a Banach space under the operator norm $\|A\|=\sup_{\|f\|_{\mathcal{H}_{1}}=1}\|Af\|_{\mathcal{H}_{2}}$, denoted by $\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$, or simply $\mathcal{B}(\mathcal{H}_{1})$ when $\mathcal{H}_{1}=\mathcal{H}_{2}$. We call an operator $A\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$ Hilbert-Schmidt if it holds $\sum_{k\geq 1}\left\|Ae_{i}\right\|_{\mathcal{H}_{2}}^{2}<\infty$ for some (equivalently, any) orthonormal basis $\{e_{k}\}_{k\geq 1}$ of $\mathcal{H}_{1}$. The set of Hilbert-Schmidt operators from $\mathcal{H}_{1}$ to $\mathcal{H}_{2}$ forms a Hilbert space under the Hilbert-Schmidt inner product $\langle A,B\rangle_{\mathrm{HS}}=\sum_{k\geq 1}\langle Ae_{k},Be_{i}\rangle_{\mathcal{H}_{2}}$ and the induced norm $\|\cdot\|_{\mathrm{HS}}$, denoted by $\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{1},\mathcal{H}_{2})$. The adjoint of $A$, denoted by $A^{*}$, is the unique operator satisfying $\langle Af,f^{\prime}\rangle_{\mathcal{H}_{2}}=\langle f,A^{*}f^{\prime}\rangle_{\mathcal{H}_{1}}$ for all $f\in\mathcal{H}_{1}$ and $f^{\prime}\in\mathcal{H}_{2}$. If $A\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$, then $A^{*}\in\mathcal{B}(\mathcal{H}_{2},\mathcal{H}_{1})$ and $\|A\|=\|A^{*}\|$. An operator $A\in\mathcal{B}(\mathcal{H}_{1})$ is called self-adjoint if $A^{*}=A$, and positive if it is self-adjoint and satisfies $\langle Af,f\rangle_{\mathcal{H}_{1}}\geq 0$ for every $f\in\mathcal{H}_{1}$. Let $(\mathcal{H}_{\mathcal{K}},\langle\cdot,\cdot\rangle_{\mathcal{H}_{\mathcal{K}}},\|\cdot\|_{\mathcal{H}_{\mathcal{K}}})$ denote the RKHS generated by the scalar-valued kernel $\mathcal{K}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. Here, we say $K$ is a scalar-valued kernel if it is a real, symmetric, and positive-definite bivariate function. A mapping $K:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{B}(\mathcal{Y})$ is called an operator-valued kernel [40, 37] on $\mathcal{X}$ if:
  1. (1)
    For any $x,x^{\prime}\in\mathcal{X}$, $K(x,x^{\prime})$ is the adjoint operator of $K(x^{\prime},x)$, i.e., $K(x,x^{\prime})^{*}=K(x^{\prime},x)$;
  2. (2)
    For any $n\in\mathbb{N}$, $\{x_{i}\}_{i=1}^{n}\subset\mathcal{X}$ and $\{y_{i}\}_{i=1}^{n}\subset\mathcal{Y}$, it holds that $\sum_{i=1}^{n}\langle K(x_{i},x_{j})y_{i},y_{j}\rangle_{\mathcal{Y}}\geq 0$.
Note that the function $K(x,\cdot)y:\mathcal{X}\rightarrow\mathcal{Y}$ is well-defined for $x\in\mathcal{X}$ and $y\in\mathcal{Y}$. The vector-valued RKHS $\mathcal{H}$ is the completion of the linear span of $\left\{K(x,\cdot)y:x\in\mathcal{X},y\in\mathcal{Y}\right\}$ with inner product $\langle K(x,\cdot)y,K(x^{\prime},\cdot)y^{\prime}\rangle_{\mathcal{H}}=\langle K(x,x^{\prime})y,y^{\prime}\rangle_{\mathcal{Y}}$. Moreover, the reproducing property holds:
$\langle K(x,\cdot)y,f\rangle_{\mathcal{H}}=\langle y,f(x)\rangle_{\mathcal{Y}},\quad\forall(x,y)\in\mathcal{X}\times\mathcal{Y}\mbox{ and }f\in\mathcal{H}.$
More details about vector-valued RKHSs refer to [37, 11, 12]. Furthermore, when $\mathcal{Y}=\mathbb{R}$, $K$ reduces to a scalar-valued kernel.
The construction of operator-valued kernels plays an important role in our setting. A common choice is
$K(x,x^{\prime})=\mathcal{K}(x,x^{\prime})W,$ (1.2)
where $\mathcal{K}$ is a scalar-valued kernel and $W\in\mathcal{B}(\mathcal{Y})$ is a positive linear operator. In multi-task learning, $W$ is typically a finite-dimensional matrix that facilitates information sharing among tasks [19, 10]. For some functional output learning problems, $W$ is selected to be a multiplication or an integral operator [24, 27]. Additionally, some works on functional regression [34] and structured output learning [5, 13, 14, 4] directly construct operator-valued kernels by setting $W$ to be the identity operator. In [28], the kernels are taken to be the finite combinations of operator-valued kernels. For other constructions, see [25].
We briefly outline some algorithms for solving model (1.1). The work in [9] studies regularized least squares estimators in vector-valued RKHSs. In addition, [25] addresses model (1.1) using spectral decomposition of block operator matrices, while [28] proposes a block-coordinate descent method. To handle limited training data, [4] leverages the structure of the target output and proposes a reduced-rank method to solve model (1.1). Note that all these existing algorithms follow the batch learning paradigms. In this paper, we adopt a stochastic gradient descent (SGD) algorithm derived from the Tikhonov regularization scheme to solve the model (1.1), aiming to learn the nonlinear operators from streaming data. This algorithm is well-suited for real-time operator learning, enabling continuous adaptation without retaining historical data—a challenge also addressed in [47, 23] in the context of operator learning. The performance of the resulting estimator $h$ can be evaluated using the prediction error $\mathcal{E}(h)-\mathcal{E}(h^{\dagger})=\mathbb{E}\left[\left\|h(x)-h^{\dagger}(x)\right\|_{\mathcal{Y}}^{2}\right]$ and the estimation error $\|h-h^{\dagger}\|_{\mathcal{H}}^{2}$, where $\mathcal{E}(h):=\mathbb{E}[\|h(x)-y\|_{\mathcal{Y}}^{2}]$. To illustrate our algorithm, we define the minimizer of the regularized least squares problem as
$h_{\lambda}:=\mathop{\arg\min}_{h\in\mathcal{H}}\left\{\mathcal{E}(h)+\lambda\|h\|_{\mathcal{H}}^{2}\right\},$ (1.3)
where $\lambda>0$ is the regularization parameter. This paper focuses on two important settings of the SGD algorithm: one with constant step sizes and regularization parameters, and the other with decaying step sizes and regularization parameters. Hereinafter, we use $\mathbf{0}$ to denote the zero element in a Hilbert space.
The finite-horizon setting. In this setting, we assume access to finite i.i.d. samples $\{z_{t}=(x_{t},y_{t})\}_{t=1}^{T}$, where the sample size $T<\infty$ is known in advance. We aim to solve the regularized problem (1.3), where the parameter $\lambda$ depends on $T$. The SGD algorithm proceeds by updating the current estimator $h_{t}$ to $h_{t+1}$ using a single sample at the $t$-th iterate with a constant step size and regularization parameter. Specifically, the iteration begins with $h_{1}=\mathbf{0}$ and is recursively defined as
$h_{t+1}=h_{t}-\eta_{T}\left(K(x_{t},\cdot)(h_{t}(x_{t})-y_{t})+\lambda_{T}h_{t}\right),\quad t=1,\cdots T,$ (1.4)
[…]
The online setting. In this setting, the sample size $T$ may be unknown in advance or even infinite, which is well-suited for scenarios that require real-time iterative updates. To accommodate this setting, we update the regularization parameter $\lambda_{t}$ such that $h_{t+1}$ follows the regularization path [50] $h_{\lambda_{t}}$, [^1] Regularization path refers to the trajectory of solutions $h_{\lambda}$ as the regularization parameter $\lambda$ varies, characterizing how the learned model evolves under different levels of regularization. ensuring that $h_{t}-h_{\lambda_{t}}\rightarrow\mathbf{0}$ and $h_{\lambda_{t}}\rightarrow h^{\dagger}$ in the norm $\|\cdot\|_{\mathcal{H}}$ (or in the semi-norm associated with prediction error) as $t$ increases. This lead to the following iterative scheme, initialized with $h_{1}=\mathbf{0}$:
$h_{t+1}=h_{t}-\eta_{t}\left(K(x_{t},\cdot)(h_{t}(x_{t})-y_{t})+\lambda_{t}h_{t}\right),\quad t\geq 1.$ (1.5)
Morevover, we let both $\eta_{t}$ and $\lambda_{t}$ decay polynomially with respect to $t$, enabling stable and convergence of the solution while adapting to streaming data and mitigating overfitting.
In this paper, we study both settings of the SGD algorithm. We express the iterative forms (1.4) and (1.5) in a unified manner as the form given in (1.5). For decaying step sizes adopted in the online setting, we set the step size as $\eta_{t}=\bar{\eta}(t+t_{0})^{-\theta_{1}}$ for all $t\geq 1$, where $\theta_{1}\in(0,1)$, $\bar{\eta}>0$, and $t_{0}>0$. The regularization parameter is defined as $\lambda_{t}=\bar{\lambda}(t+t_{0})^{-\theta_{2}}$, where $\theta_{2}\in(0,1)$, $\bar{\lambda}>0$. We emphasize that in the online setting, both $\bar{\eta}$ and $\bar{\lambda}$ are constants independent of $t$ and the total number of iterations (e.g., the sample size) $T$. For the constant step sizes and regularization parameters adopted in the finite-horizon setting, we set $\eta_{t}=\eta_{1}T^{-\theta_{3}}$ and $\lambda_{t}=\lambda_{1}T^{-\theta_{4}}$ for $t=1,2,\dots,T$, where $\theta_{3}\in(0,1)$, $\theta_{4}>0$, and $\eta_{1},\lambda_{1}>0$. In this finite-horizon setting, the step sizes and regularization parameters explicitly depend on the total number of iterations $T$.
Throughout the paper, we impose the following assumption on the operator-valued kernels.
Assumption 1.
The vector-valued RKHS $\mathcal{H}$ is generated by the operator-valued kernel $K(x,x^{\prime})=\mathcal{K}(x,x^{\prime})I$, where $\mathcal{K}$ is the scalar-valued kernel with $\|\mathcal{K}\|_{\infty}\leq\kappa^{2}$ for some constant $\kappa>0$, and $I$ is the identity operator on $\mathcal{Y}$.
This simple construction of operator-valued kernels has been adopted in previous works, e.g., [4, 1]. Note that all elements $h$ in the vector-valued RKHS $\mathcal{H}$ are measurable. We also consider a more general class of kernels in Section 3.1. In particular, our analysis covers most operator-valued kernels, including those of the form (1.2) with a compact operator $W$. Furthermore, when choosing kernels as in (1.2), it follows from [12, Example 5] that $K$ is a Mercer [resp. $\mathcal{C}_{0}$] [^2] That is, the Banach space of continuous functions vanishing at infinity with the uniform norm. kernel if $\mathcal{K}$ is Mercer [resp. $\mathcal{C}_{0}$], implying that all operators in $\mathcal{H}$ are continuous.
The two types of step sizes considered in this paper have been extensively studied in the previous literature on SGD in various settings. The seminal work [42] shows that the step size serves as an implicit form of regularization, thus improving the algorithm’s generalization and robustness. Our recent work [41] investigates operator learning via the SGD algorithm between Hilbert spaces, deriving bounds for both prediction and estimation errors in expectation. However, the SGD algorithm in [41] does not incorporate the regularization term. On the other hand, in the context of operator learning, research on the almost-sure convergence of SGD algorithms is still scarce. Only a few works, including [43, 2, 44], have considered the almost-sure convergence in finite-dimensional output settings but either assume a noise-free scenario or provide convergence results that do not directly extend to operator learning problems. This clearly identifies a significant gap in the existing literature. Consequently, while advancements have occurred in finite-dimensional settings, substantial challenges related to operator learning and the fundamental role of regularization remain largely unexplored.
This paper aims to fill this gap by rigorously analyzing the regularized SGD algorithm applied to the nonlinear operator regression problem described in (1.1). Our main contributions are summarized as follows: First, we introduce specific regularity assumptions on $h^{\dagger}$ (or its associated Hilbert-Schmidt operator, as defined in Proposition 2.1), which effectively capture the intrinsic features of infinite-dimensional regression problems. Under these assumptions, we derive bounds in expectation for the prediction and estimation errors of the regularized SGD algorithm, showing improvements compared to the unregularized SGD algorithm studied in [41]. Second, we propose a novel technique for establishing high-probability bounds, ensuring almost-sure convergence via the Borel–Cantelli lemma. High-probability convergence provides a stronger guarantee than expectation-based bounds alone, moving beyond average-case performance. Crucially, we demonstrate that the introduction and careful tuning of regularization parameters are essential not only for achieving these high-probability bounds but also for significantly enhancing the convergence behaviors of the SGD algorithm. This underscores the superiority of our regularized approach over the unregularized framework considered in [41], highlighting the necessity of regularization for robust probabilistic guarantees. Lastly, the resulting convergence rates are demonstrated to be near-optimal, aligning closely with the minimax lower bounds established in [41], thus reinforcing the theoretical soundness and effectiveness of our proposed algorithm.
The rest of the paper is organized as follows. Section 2 introduces the main theoretical results and the required assumptions. In Section 3, we discuss several possible extensions of our framework, including general operator-valued kernels, structured output settings, and the encoder-decoder paradigm. Section 4 performs an error decomposition tailored to the regularized SGD algorithm. Building on this, Section 5 provides essential intermediate estimates used in subsequent analysis. Sections 6 and 7 are devoted to establishing bounds on the prediction and estimation errors—first in expectation, then with high probability. For clarity and conciseness, some technical proofs are presented in the appendix.
2 Main Results
This section introduces regularity conditions on the structure and smoothness of the target operator and the input random variables. We then present our main theorems. We begin with some notations for further statements. Denote $\mathbb{N}_{T}$ as the set $\{1,2,\cdots,T\}$. The rank-one operator $f\otimes g\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$ is defined by $f\otimes g(g^{\prime}):=\langle g,g^{\prime}\rangle_{\mathcal{H}_{1}}f$, where $g,g^{\prime}\in\mathcal{H}_{1}$ and $f\in\mathcal{H}_{2}$. We denote $\mathrm{Tr}(A)$ as the trace of a self-adjoint and compact operator $A\in\mathcal{B}(\mathcal{H}_{1})$. Let $\mathbb{E}$ and $\mathbb{E}_{z_{t}}$ denote the expectation with respect to the distribution $\rho$ and the sample $z_{t}:=(x_{t},y_{t})$, respectively. For $k\in\mathbb{N}_{T}$, let $\mathbb{E}_{z_{1},\cdots,z_{k}}$ denote the expectation with respect to $\{z_{i}\}_{i=1}^{k}$, abbreviated as $\mathbb{E}_{z^{k}}$. Recall that $\mathcal{K}$ is the scalar-valued kernel. Since $\mathcal{X}$ is separable, the RKHS $\mathcal{H}_{\mathcal{K}}$ induced by $\mathcal{K}$ is also separable. The operator $C=\mathbb{E}[\phi(x)\otimes\phi(x)]$, defined by $\phi(x):=\mathcal{K}(x,\cdot)\in\mathcal{H}_{\mathcal{K}}$, is self-adjoint, compact, and satisfies $\|C\|\leq\|C\|_{\mathrm{HS}}\leq\kappa^{2}$. Thus, for any $r>0$, the operator $C^{r}$ is also self-adjoint and compact. Moreover, it is straightforward to verify that
$\|C^{1/2}\|^{2}_{\mathrm{HS}}=\mathrm{Tr}(C)=\mathbb{E}\left[\|\phi(x)\|_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\leq\kappa^{2}.$
With the aid of the following proposition, the iterative process in the vector-valued RKHS $\mathcal{H}$ can be equivalently reformulated as an iterative process in $\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$.
Proposition 2.1.
The vector-valued RKHS $\mathcal{H}$, associated with the operator-valued kernel $K(x,x^{\prime})=\mathcal{K}(x,x^{\prime})W$, where $W$ is a positive operator and $\mathcal{K}$ is a scalar-valued kernel, is isometrically isomorphic to $\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\overline{W^{1/2}\mathcal{Y}})\subset\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$. Specifically, for each $h\in\mathcal{H}$, there exists a unique $H\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\overline{W^{1/2}\mathcal{Y}})$ such that
$h(x)=W^{1/2}H\phi(x),\quad\forall x\in\mathcal{X},$
and $\|h\|_{\mathcal{H}}=\|H\|_{\mathrm{HS}}$.
The proof of Proposition 2.1 is deferred to Appendix A.1. By applying Proposition 2.1 with $W=I$—in which case the kernel coincides with that specified by Assumption 1—the iteration (1.5) can be equivalently expressed as
$\begin{cases}H_{1}=\mathbf{0},\\ H_{t+1}=H_{t}-\eta_{t}\left(\left(H_{t}\phi(x_{t})-y_{t}\right)\otimes\phi(x_{t})+\lambda_{t}H_{t}\right),\\ h_{t}(\cdot)=H_{t}\left(\phi(\cdot)\right).\end{cases}$ (2.1)
Hereinafter, we assume that $h^{\dagger}(x)=H^{\dagger}\phi(x)$, where $H^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ is a Hilbert-Schmidt operator. Under this assumption, the nonlinear operator learning model (1.1) reduces to an infinite-dimensional linear model:
$y=H^{\dagger}\phi(x)+\epsilon,$ (2.2)
where the input and output are $\phi(x)$ and $y$, respectively. We define the prediction error of $H\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ as $\mathcal{E}(H)=\mathbb{E}\left[\|y-H\phi(x)\|_{\mathcal{Y}}^{2}\right]=\mathcal{E}(h)$, and the estimation error of $H$ as $\mathbb{E}[\|H-H^{\dagger}\|^{2}_{\mathrm{HS}}]$, where $h(x)=H\phi(x)$. According to Proposition 2.1, we have $\mathbb{E}[\|h-h^{\dagger}\|^{2}_{\mathcal{H}}]=\mathbb{E}[\|H-H^{\dagger}\|^{2}_{\mathrm{HS}}]$, which implies that the prediction error and estimation error for the estimator $h(\cdot)=H\phi(\cdot)$ in the original model (1.1) coincide with those of $H$ in the linearized model (1.1). Therefore, it suffices to analyze the convergence rates of the errors of $H_{t}$ associated with the SGD iteration (2.1) in $\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$. Although this is not directly required for our theoretical analysis, we emphasize for clarity that the iterative form of SGD derived from minimizing the regularized objective functional $\mathcal{E}(H)+\lambda\|H\|_{\mathrm{HS}}^{2}$ corresponds exactly to the iteration given in (2.1), which is equivalent to (1.5).
2.1 Assumptions
To conduct the convergence analysis, we need the following assumptions.
Assumption 2 (Regularity condition of $H^{\dagger}$).
There exists a Hilbert-Schmidt operator $S^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ and a positive parameter $r>0$, such that:
$H^{\dagger}=S^{\dagger}C^{r}.$
This assumption, introduced in [41], characterizes the regularity of the target operator $H^{\dagger}$ via its relation to the operator $C$. The parameter $r$ serves as a smoothness index—larger values of $r$ indicate higher regularity of $H^{\dagger}$. In the special case where $\mathcal{Y}=\mathbb{R}$, the Riesz representation theorem implies that $H^{\dagger}$ corresponds to an element $g^{\dagger}=C^{r}g$ in $\mathcal{H}_{\mathcal{K}}$ for some $g\in\mathcal{H}_{\mathcal{K}}$. This is exactly the regularity condition widely adopted in the convergence analysis of non-parametric regression in RKHS [48, 16, 2, 22].
Assumption 3 (Spectral decay condition of $C$).
There exists $s\in(0,1]$ such that:
$\mathrm{Tr}(C^{s})<+\infty.$
This condition is automatically satisfied for any $s\geq 1$ (as $\mathrm{Tr}(C)<\kappa^{2}$), and it imposes constraints on the decay rate of the eigenvalues of the operator $C$. Let $\{u_{k}\}_{k\geq 1}$ denote the non-increasing sequence of eigenvalues of $C$. Under this condition, the eigenvalues exhibit polynomial decay, specifically satisfying
$u_{k}\leq\mathrm{Tr}(C^{s})^{\frac{1}{s}}k^{-\frac{1}{s}}.$
A sufficient (though not necessary) condition for this assumption is that $u_{k}=O(k^{-\frac{1}{s}-\epsilon})$ for some $\epsilon>0$. For a detailed discussion on this condition, we refer the reader to [22]. In this paper, Assumption 3 is introduced to derive sharper error bounds. When combined with Assumption 3 for some $10<s<$ 1, it leads to improved convergence rates. This condition—commonly known as the capacity condition—was first introduced in [17] and has since been widely adopted in the literature, including [38, 4, 22, 41], as a way to capture the intrinsic complexity of infinite-dimensional learning problems. Assumptions 2 and 3 are essential to establish dimension-free convergence analysis. As we will show, the resulting convergence rates depend explicitly on the parameters $r$ and $s$, reflecting the regularity of the target operator and the capacity of the input random variables, respectively.
The following assumption is only required for establishing error bounds in expectation. Recall that $\phi(x):=\mathcal{K}(x,\cdot)\in\mathcal{H}_{\mathcal{K}}$ for some scalar-valued kernel $\mathcal{K}$.
Assumption 4 (Moment condition of $\phi(x)$).
There exists a constant $c>0$ such that for any compact linear operator $A\in\mathcal{B}(\mathcal{H}_{\mathcal{K}})$,
$\mathbb{E}\left[\left\|A\phi(x)\right\|^{4}_{\mathcal{H}_{\mathcal{K}}}\right]\leq c\left(\mathbb{E}\left[\|A\phi(x)\|^{2}_{\mathcal{H}_{\mathcal{K}}}\right]\right)^{2}.$
According to [41, Proposition 2.1], this assumption is equivalent to
$\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\right]\leq c\left(\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\right)^{2},\quad\forall f\in\mathcal{H}_{\mathcal{K}}.$ (2.3)
Condition (2.3) holds, for example, when $\phi(x)$ is strictly sub-Gaussian, implying that all linear functionals of $\phi(x)$ have bounded kurtosis. Similar assumptions have been adopted in several papers [49, 8, 22, 41]. To further deepen our understanding, we now present a novel characterization of Assumption 4, which is analogous to the idea discussed in [35].
Proposition 2.2.
Consider the principal component decomposition of $\phi(x)$:
$\phi(x)=\overline{\phi}+\sum_{k\geq 1}\sqrt{\lambda_{k}}\xi_{k}\phi_{k},$ (2.4)
where $\overline{\phi}:=\mathbb{E}[\phi(x)]$, and $\left\{\left(\lambda_{k},\phi_{k}\right)\right\}_{k\geq 1}$ are the eigenvalue-eigenvector pairs of the covariance operator $\Sigma:=\mathbb{E}\left[\left(\phi(x)-\overline{\phi}\right)\otimes\left(\phi(x)-\overline{\phi}\right)\right]$. The sequence $\{\xi_{k}\}_{k\geq 1}$ consists of zero-mean, uncorrelated real-valued random variables with $\mathbb{E}[\xi_{k}^{2}]=1$. If, in addition, $\{\xi_{k}\}_{k\geq 1}$ are independent, then Assumption 4 (or equivalently, (2.3)) holds provided that $\{\mathbb{E}[\xi_{k}^{4}]\}_{k\geq 1}$ are uniformly bounded. That is, there exists a constant $C>0$ such that
$\mathbb{E}\left[\xi_{k}^{4}\right]\leq C,\quad\forall k\geq 1.$
The proof of the above proposition is presented in Appendix A.2. The next assumption is used to derive high-probability error bounds.
Assumption 5 (Boundedness condition of $y$).
There exists some constant $M_{\rho}>0$ such that
$\|y\|_{\mathcal{Y}}\leq M_{\rho}$
almost surely.
2.2 Error Bounds in Expectation
In this subsection, we assume that Assumption 1 holds, Assumption 2 holds with $S^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ and $r>0$, Assumption 3 holds with $0<s\leq 1$, and Assumption 4 holds with $c>0$. Theorem 2.3 and Theorem 2.4 provide the convergence rates of prediction error and estimation error in expectation for the online setting. In contrast, Theorem 2.5 and 2.6 focus on the finite-horizon setting.
Theorem 2.3.
Suppose that Assumption 1, Assumption 2, Assumption 3 and Assumption 4 are satisfied. Define $\{h_{t}\}_{t\geq 1}$ through (2.1) with step sizes $\{\eta_{t}=\bar{\eta}(t+t_{0})^{-\theta_{1}}\}_{t\geq 1}$ and regularization parameters $\{\lambda_{t}=\bar{\lambda}(t+t_{0})^{-\theta_{2}}\}_{t\geq 1}$, where $10<\theta_{1}<$ 1, $10<\theta_{2}<$ 1 and $\bar{\eta}\bar{\lambda}>\theta_{2}\min\{r,1\}$. Additionally, let $t_{0}$ satisfy $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, $t_{0}\geq\exp\{\frac{1}{\theta_{1}}\}$, and
$c_{4}\sqrt{c}t_{0}^{-\theta_{1}}\log t_{0}<1,$
where $c_{4}$ is a constant independent of $t_{0}$, as specified in Proposition 5.9. Choose $\theta_{1}=\min\left\{\frac{2r+1}{2r+2},\frac{2}{3}\right\}$ and $\theta_{2}=1-\theta_{1}$. Then for any $T\geq 1$,
$\mathbb{E}_{z^{T}}\left[\mathcal{E}(h_{T+1})-\mathcal{E}(h^{\dagger})\right]\leq c_{1,1}\begin{cases}(T+t_{0})^{-\theta_{1}},&\text{when }s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0}),&\text{when }s=1.\\ \end{cases}$
Here the constant $c_{1,1}$ is independent of $T$, and will be given in the proof.
Remark 1.
In the theorem above, we set $\theta_{1}+\theta_{2}=1$, as this choice leads to the most favorable convergence rates achievable within our framework. The condition on $t_{0}$ is necessary for the proof, while the constraint on $\bar{\eta}\bar{\lambda}$ serves to accelerate convergence. When $\theta_{1}+\theta_{2}\neq 1$, the resulting rates are slower. In such cases, one can set $t_{0}=0$ and choose a small $\bar{\eta}\bar{\lambda}$; the corresponding analysis is similar and more straightforward, so we omit it for brevity.
In Theorem 2.3, since the constant $c_{4}$ is independent of $t_{0}$, one can choose $t_{0}$ sufficiently large to satisfy the required conditions. Compared to Assumption 3 with $s=1$, the stronger assumption with $10<s<$ 1 only removes a logarithmic factor in the convergence rate. It is also clear that the convergence rate saturates at $r=1/2$, i.e., increasing $r$ beyond $1/2$ does not yield further improvement. According to [41, Theorem 2.9], the result is minimax optimal (up to a logarithmic term) when $s=1$ and $r<1/2$. Compared to the unregularized SGD algorithm analyzed in [41, Theorem 2.4], adding a regularization term here leads to faster convergence. Specifically, while the prediction error rate of unregularized SGD in [41] saturates at $r=(1-s)/2$, the regularized SGD in our work improves the saturation level to $r=1/2$.
The following theorem provides the convergence rate for the estimation error.
Theorem 2.4.
Under the conditions of Theorem 2.3, choose $\theta_{1}=\min\left\{\frac{s+2r}{1+s+2r},\frac{2+s}{3+s}\right\}$ and $\theta_{2}=1-\theta_{1}$. Then for any $T\geq 1$,
$\mathbb{E}_{z^{T}}\left[\left\|h_{T+1}-h^{\dagger}\right\|_{\mathcal{H}}^{2}\right]\leq c_{1,2}(T+t_{0})^{-\min\left\{\frac{2r}{1+s+2r},\frac{2}{3+s}\right\}}.$
Here the constant $c_{1,2}$ is independent of $T$, and will be given in the proof.
The convergence rate of the estimation error saturates at $r=1$, which improves the convergence of unregularized SGD in [41, Theorem 2.6], where the rate saturates at $r=\frac{1-s}{2}$. Moreover, in [22, Theorem 3] and [41, Theorem 2.6], we cannot guarantee the convergence of the estimation error with decaying step sizes for $s=1$, whereas adding a regularization term addresses this issue. According to [41, Theorem 2.9], the convergence rate with decaying step sizes is minimax optimal when $r<1$.
Next, we present the convergence rates for prediction and estimation errors with constant step sizes and regularization parameters, where both depend on the total number of iterations $T$ (i.e., the total sample size).
Theorem 2.5.
Suppose that Assumption 1, Assumption 2, Assumption 3 and Assumption 4 are satisfied. Define $\{h_{t}\}_{t\in\mathbb{N}_{T}}$ through (2.1) with step sizes $\{\eta_{t}=\eta_{1}T^{-\theta_{3}}\}_{t\in\mathbb{N}_{T}}$ and regularization parameters $\{\lambda_{t}=\lambda_{1}T^{-\theta_{4}}\}_{t\in{\mathbb{N}_{T}}}$, where $T\geq 2$, $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$, and
$\eta_{1}<\frac{1}{6c\kappa^{2}\left(1+\frac{1}{2e\theta_{3}}\right)}.$
Choose $\theta_{3}=\frac{2r+1}{2r+2}$ and $\theta_{4}\geq\frac{2r+1}{(2r+2)\min\{2r+1,2\}}$. Then
$\mathbb{E}_{z^{T}}[\mathcal{E}(h_{T+1})-\mathcal{E}(h^{\dagger})]\leq c_{1,3}\begin{cases}T^{-\frac{2r+1}{2r+2}},&\text{ when }s<1,\\ T^{-\frac{2r+1}{2r+2}}\log T,&\text{ when }s=1.\end{cases}$
Here the constant $c_{1,3}$ is independent of $T$, and will be given in the proof.
Theorem 2.6.
Under the conditions of Theorem 2.5, choose $\theta_{3}=\frac{2r+s}{1+2r+s}$ and $\theta_{4}\geq\frac{2r}{(1+2r+s)\min\{2r,2\}}$. Then
$\mathbb{E}_{z^{T}}[\left\|h_{T+1}-h^{\dagger}\right\|_{\mathcal{H}}^{2}]\leq c_{1,4}T^{-\frac{2r}{1+2r+s}}.$
Here the constant $c_{1,4}$ is independent of $T$, and will be given in the proof.
In the case of constant step sizes and regularization parameters, the prediction error achieves the minimax optimal rate when $s=1$, and the estimation error performs so for any $s>0$, as established by the minimax lower bounds in [41]. Unlike the scenario with decaying step sizes and regularization parameters, no saturation occurs when these parameters are held constant. It is also noteworthy that the convergence rates and the choice of $\theta_{3}$ in the above two theorems align with those in Theorems 2.5 and 2.7 of [41], which analyze the unregularized SGD algorithm with constant step sizes. This contrasts with the case of decaying step size, where adding a regularization term leads to improved rates. When employing constant step sizes, introducing regularization does not improve the convergence rates; in fact, an improperly chosen $\theta_{4}$ (not sufficiently large) may degrade performance. The unregularized SGD can be viewed as the limiting case corresponding to $\theta_{4}=\infty$.
2.3 High-probability Error Bounds
In this subsection, we assume Assumption 1 holds, Assumption 2 holds with $S^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ and $r>0$, Assumption 3 holds with $0<s\leq 1$, and Assumption 5 holds with $M_{\rho}>0$. We derive high-probability error bounds for both the prediction and estimation errors in both the online and finite-horizon settings. These error bounds guarantee almost-sure convergence of the regularized SGD algorithm, providing a stronger guarantee than bounds in expectation, as convergence is ensured with high probability across all realizations. The notation $a\lesssim b$ denotes $a\leq Cb$ for some constant $C$ independent of $t$, $T$, and the confidence level $\delta$.
The following theorem establishes the prediction error bounds in the online setting.
Theorem 2.7.
Suppose that Assumption 1, Assumption 2, Assumption 3 and Assumption 5 are satisfied. Define $\{h_{t}\}_{t\geq 1}$ through (2.1) with step sizes $\{\eta_{t}=\bar{\eta}(t+t_{0})^{-\theta_{1}}\}_{t\geq 1}$ and regularization parameters $\{\lambda_{t}=\bar{\lambda}(t+t_{0})^{-\theta_{2}}\}_{t\geq 1}$, where $\bar{\eta}\bar{\lambda}>\max\{\theta_{2}\min\{r,1\},\theta_{1},2\theta_{1}-\frac{1}{2}\}$ and $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$. Choose
$\theta_{1}=\begin{cases}\frac{2r+1}{2r+2},&\text{ when }r<\frac{1}{2},\\ \frac{2}{3},&\text{ when }r\geq\frac{1}{2},\end{cases}$
and $\theta_{2}=1-\theta_{1}$. Then for any $T\geq 1$, with probability at least $1-2\delta$, the following holds:
  • (1)
    If $s<1$,
    $\mathcal{E}(h_{T+1})-\mathcal{E}(h^{\dagger})$ $\leq c_{2,1}\left((T+t_{0})^{-\theta_{1}}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}$
    $\lesssim(T+t_{0})^{-\theta_{1}}\log^{4}\frac{2}{\delta}.$
  • (2)
    If $s=1$,
    $\mathcal{E}(h_{T+1})-\mathcal{E}(h^{\dagger})$ $\leq c_{2,1}\left((T+t_{0})^{-\theta_{1}}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)\log(T+t_{0})\log^{2}\frac{2}{\delta}$
    $\lesssim(T+t_{0})^{-\theta_{1}}\log(T+t_{0})\log^{4}\frac{2}{\delta}.$
Here the constant $c_{2,1}$ is independent of $T$ and $\delta$, and will be given in the proof.
The following corollary, as a natural extension of Theorem 2.7, establishes a uniform high-probability bound that holds simultaneously for all $t\geq 1$.
Corollary 2.8.
Under the conditions of Theorem 2.7, choose
$\theta_{1}=\begin{cases}\frac{2r+1}{2r+2},&\text{ when }r<\frac{1}{2},\\ \frac{2}{3},&\text{ when }r\geq\frac{1}{2},\end{cases}$
and $\theta_{2}=1-\theta_{1}$. Then, with probability at least $1-2\delta$ , for all $1\leq t<\infty$, the following holds:
$\mathcal{E}(h_{t+1})-\mathcal{E}(h^{\dagger})\leq\widetilde{c}_{2,1}\begin{cases}(t+t_{0})^{-\theta_{1}}\log^{4}(t+t_{0})\log^{4}\frac{2}{\delta},&\text{when }s<1,\\ (t+t_{0})^{-\theta_{1}}\log^{5}(t+t_{0})\log^{4}\frac{2}{\delta},&\text{ when }s=1.\end{cases}$
Here the constant $\widetilde{c}_{2,1}$ is independent of $t$ and $\delta$.
In Theorem 2.9 and Corollary 2.10, we focus on the estimation error in the online setting.
Theorem 2.9.
Under the conditions of Theorem 2.7, choose
$\theta_{1}=\begin{cases}\frac{1+2r+s}{3+2r+s},&\text{ when }r<\frac{1-s}{2},\\ \frac{2\min\{r,1\}+s}{1+2\min\{r,1\}+s},&\text{ when }r\geq\frac{1-s}{2},\end{cases}$
and $\theta_{2}=1-\theta_{1}$. Then for any $T\geq 1$, with at least $1-2\delta$ probability, the following holds:
  • (1)
    If $r<\frac{1-s}{2}$,
    $\left\|h_{T+1}-h^{\dagger}\right\|_{\mathcal{H}}^{2}$ $\leq c_{2,2}(T+t_{0})^{-\frac{4r}{3+2r+s}}\log^{2}(T+t_{0})\log^{4}\frac{2}{\delta}.$
  • (2)
    If $r\geq\frac{1-s}{2}$,
    $\left\|h_{T+1}-h^{\dagger}\right\|_{\mathcal{H}}^{2}$ $\leq c_{2,2}\left((T+t_{0})^{-\frac{2\min\{r,1\}}{1+2\min\{r,1\}+s}}+(T+t_{0})^{-\frac{4\min\{r,1\}+s-1}{1+2\min\{r,1\}+s}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}$
    $\lesssim(T+t_{0})^{-\frac{2\min\{r,1\}}{1+2\min\{r,1\}+s}}\log^{4}\frac{2}{\delta}.$
Here the constant $c_{2,2}$ is independent of $T$ and $\delta$, and will be given in the proof.
Corollary 2.10.
Under conditions of Theorem 2.7, choose $\theta_{1}$ and $\theta_{2}$ as in Theorem 2.9. Then, with probability at least $1-2\delta$ , for all $1\leq t<\infty$, the following holds:
$\left\|h_{t+1}-h^{\dagger}\right\|_{\mathcal{H}}^{2}\leq\begin{cases}\widetilde{c}_{2,2}(t+t_{0})^{-\frac{4r}{3+2r+s}}\log^{6}(t+t_{0})\log^{4}\frac{2}{\delta},&\text{ when }r<\frac{1-s}{2},\\ \widetilde{c}_{2,2}(t+t_{0})^{-\frac{2\min\{r,1\}}{1+2\min\{r,1\}+s}}\log^{4}(t+t_{0})\log^{4}\frac{2}{\delta},&\text{ when }r\geq\frac{1-s}{2}.\end{cases}$
Here the constant $\widetilde{c}_{2,2}$ is independent of $t$ and $\delta$.
The following two theorems provide high-probability convergence rates for the prediction and estimation errors, respectively, in the finite-horizon setting.
Theorem 2.11.
Suppose that Assumption 1, Assumption 2, Assumption 3 and Assumption 5 are satisfied. Define $\{h_{t}\}_{t\in\mathbb{N}_{T}}$ through (2.1) with step sizes $\{\eta_{t}=\eta_{1}T^{-\theta_{3}}\}_{t\in\mathbb{N}_{T}}$ and regularization parameters $\{\lambda_{t}=\lambda_{1}T^{-\theta_{4}}\}_{t\in{\mathbb{N}_{T}}}$, where $T\geq 2$ and $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$. Choose $\theta_{3}=\frac{2r+1}{2r+2}$ and $\theta_{4}\geq\frac{2r+1}{(2r+2)\min\{2r+1,2\}}$. Then, with probability at least $1-2\delta$,
$\mathcal{E}(h_{T+1})-\mathcal{E}(h^{\dagger})$ $\leq c_{2,3}\begin{cases}T^{-\theta_{3}}\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{2}T\log^{4}\frac{2}{\delta},&\text{ when }s<1,\\ T^{-\theta_{3}}\log T\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{3}T\log^{4}\frac{2}{\delta},&\text{ when }s=1,\\ \end{cases}$
$\lesssim\log^{4}\frac{2}{\delta}\begin{cases}T^{-\theta_{3}},&\text{ when }s<1,\\ T^{-\theta_{3}}\log T,&\text{ when }s=1.\end{cases}$
Here the constant $c_{2,3}$ is independent of $T$ and $\delta$, and will be given in the proof.
Theorem 2.12.
Under the conditions of Theorem 2.11, choose
$\theta_{3}=\begin{cases}\frac{1+2r+s}{3+2r+s},&\text{ when }r<\frac{1-s}{2},\\ \frac{2r+s}{1+2r+s},&\text{ when }r\geq\frac{1-s}{2},\end{cases}$
and
$\theta_{4}\geq\begin{cases}\frac{2r}{(3+2r+s)r},&\text{ when }r<\frac{1-s}{2},\\ \frac{r}{(1+2r+s)\min\{r,1\}},&\text{ when }r\geq\frac{1-s}{2}.\end{cases}$
Then, with probability at least $1-2\delta$, the following holds:
  • (1)
    If $r<\frac{1-s}{2}$,
    $\left\|h_{T+1}-h^{\dagger}\right\|_{\mathcal{H}}^{2}$ $\leq c_{2,4}\left(T^{-\frac{1+2r-s}{3+2r+s}}+T^{-\frac{4r}{3+2r+s}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}$
    $\lesssim T^{-\frac{4r}{3+2r+s}}\log^{2}T\log^{4}\frac{2}{\delta}.$
  • (2)
    If $r\geq\frac{1-s}{2}$,
    $\left\|h_{T+1}-h^{\dagger}\right\|_{\mathcal{H}}^{2}$ $\leq c_{2,4}\left(T^{-\frac{2r}{1+2r+s}}+T^{-\frac{4r+s-1}{1+2r+s}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}$
    $\lesssim T^{-\frac{2r}{1+2r+s}}\log^{4}\frac{2}{\delta}.$
Here the constant $c_{2,4}$ is independent of $T$ and $\delta$, and will be given in the proof.
Building upon the results from the previous theorems, we note that Theorem 2.11 for $s=1$ and Theorem 2.12 for $r\geq\frac{1-s}{2}$ achieve the minimax lower bound derived in Theorem 2.9 of [41], up to a logarithmic factor.
3 Discussion
This section briefly illustrates how our results can be extended to the setting of general kernels and applied to structured prediction problems. Toward the end of the section, we also discuss how our results can be combined with the encoder-decoder framework via principal component analysis (PCA). Theoretical analysis of these topics will be presented in our future work.
3.1 Extension to General Kernel Setting
In the previous section, we established the error analysis on the regularized SGD for solving the nonlinear operator learning problem, considering both the online setting and the finite-horizon setting. The operator-valued kernel that induces the corresponding RKHS is assumed to be of the form $K=\mathcal{K}I$, as specified in Assumption 1. This choice of kernel has been employed in functional regression with structured output learning, as noted in Section 1. We now turn to extending the class of operator-valued kernels, further showing the generality and applicability of our analysis to a broader range of nonlinear operator learning problems.
We now consider an alternative setting for vector-valued RKHS and briefly list the conditions below.
  1. (1)
    Let $\mathcal{H}$ be separable, which is true if the spaces $\mathcal{X}$ and $\mathcal{Y}$ are separable and $K$ is a Mercer kernel [11, Corollary 5.2]. A kernel $K$ is Mercer if and only if the RKHS induced by $K$ is a subspace of the spaces of continuous operators from $\mathcal{X}$ to $\mathcal{Y}$, which in turn holds if and only if $K$ is locally bounded and $K(x,\cdot)$ is strongly continuous for any $x\in\mathcal{X}$ [11, Proposition 5.1].
  2. (2)
    We assume that the operator-valued kernel $K:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{B}(\mathcal{Y})$ satisfies that $K(x,x)$ is compact for any $x\in\mathcal{X}$.
  3. (3)
    We further assume that $K$ is strongly measurable. Under assumptions (1) and (2), this is equivalent to requiring that each element in $\mathcal{H}$ is a measurable function [11, Proposition 3.3].
  4. (4)
    By Corollary 4.6 and Proposition 4.8 in [11], if $K(x,x)$ is an operator of trace class (which implies compactness obviously) for almost all $x\in\mathcal{X}$ and
    $\mathbb{E}\left[\mathrm{Tr}\left(K(x,x)\right)\right]<\infty,$ (3.1)
    then the inclusion $\iota:\mathcal{H}\rightarrow L^{2}(\mathcal{X},\mathcal{Y})$ is well-defined and Hilbert-Schmidt. As a result, the operator $L_{K}:=\iota^{*}\iota$ is trace-class, which plays a role similar to that of $C$. And thus, we assume Assumption 3 holds for $L_{K}$ with $s\in(0,1]$, i.e., $\mathrm{Tr}(L_{K}^{s})<+\infty$.
The conditions listed above are required when conducting error analysis for kernels beyond the special case considered in Assumption 1. We remark that when the kernel is chosen as (1.2), i.e., $K(x,x^{\prime})=\mathcal{K}(x,x^{\prime})W$ with $W$ being a self-adjoint and positive operator, the above conditions are satisfied if $W$ is trace-class and the scalar-valued kernel $\mathcal{K}$ is a Mercer kernel such that $\sup_{x\in\mathcal{X}}\mathcal{K}(x,x)<\infty$.
We now outline the framework for generalizing the conclusions of this paper to the general scenario discussed above, while leaving the detailed proof to future work. It is straightforward to observe that for any $h\in\mathcal{H}$, the operator $L_{K}$ satisfies
$L_{K}h=\mathbb{E}\left[K(x,\cdot)h(x)\right].$
Moreover, for any $h\in\mathcal{H}$, since the noise $\epsilon$ is centered and independent of $x$, there holds
$\mathcal{E}(h)-\mathcal{E}(h^{\dagger})$ $=\mathbb{E}\left\|h(x)-h^{\dagger}(x)\right\|_{\mathcal{Y}}^{2}$
$=\mathbb{E}\left[\left\langle K(x,\cdot)(h(x)-h^{\dagger}(x)),h-h^{\dagger}\right\rangle_{\mathcal{H}}\right]$
$=\left\|L_{K}^{1/2}\left(h-h^{\dagger}\right)\right\|_{\mathcal{H}}^{2}.$
Our goal is to bound the prediction error $\left\|L_{K}^{1/2}\left(h-h^{\dagger}\right)\right\|_{\mathcal{H}}^{2}$ and estimation error $\left\|h-h^{\dagger}\right\|_{\mathcal{H}}^{2}$ for the regularized SGD estimator $h=h_{T+1}$. Similar to the approach in the proofs of our main results, we can derive analogs of equations (4.5), (4.6), Proposition 4.2, and Proposition 4.3. Next, under assumptions similar to those in Section 2, we can carry out the error analysis, which we leave to future work.
We point out that the framework discussed in this subsection does not cover the case considered in Assumption 1, as the kernel $\mathcal{K}(x,x)I$ is not a compact operator and thus fails to satisfy the assumptions required in the current setting. Hence, the framework developed here and the one based on Assumption 1 are mutually exclusive. Nevertheless, combining the kernel choices from Assumption 1 and those introduced in this subsection can significantly broaden the applicability of our analysis developed in this paper.
3.2 Application to Structured Prediction
Now, we formulate the surrogate approach for structured prediction as an application example of our model (1.1). In structured prediction, the input takes values in $\mathcal{X}$ and the output takes values in $\mathcal{Z}$, where $\mathcal{X}$ is a Polish space and $\mathcal{Z}$ represents the structured output space. A structured loss function $\mathscr{D}:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$ is defined on $\mathcal{Z}$ to measure the discrepancy between the true and the predicted outputs. Let $x$ denote the input random variable and $z$ the output random variable. Given a set of independent and identically distributed input-output samples, our goal is to learn a mapping from the inputs to structured outputs. To this end, we minimize the prediction error defined by
$\mathcal{R}(f):=\mathbb{E}\left[\mathscr{D}(f(x),z)\right],$
where $f$ is an estimator of $f^{\dagger}$. The function $f^{\dagger}:\mathcal{X}\to\mathcal{Z}$ is the minimizer of $\mathcal{R}$, i.e., $f^{\dagger}=\arg\min_{f}\mathcal{R}(f)$.
We focus on the case where the loss function $\mathscr{D}$ is induced by a scalar-valued kernel $k_{\mathcal{Z}}:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$. Specifically, denote the RKHS induced by $k_{\mathcal{Z}}$ by $\left(\mathcal{Y},\|\cdot\|{\mathcal{Y}}\right)$, and embed $\mathcal{Z}$ into $\mathcal{Y}$ via the canonical feature map $\phi(z):=k_{\mathcal{Z}}(z,\cdot)$. We then define the structured loss as $\mathscr{D}(z,z^{\prime})=\|\phi(z)-\phi(z^{\prime})\|_{\mathcal{Y}}^{2}$. Building on extensive research on kernels for structured objects [20], this class of loss functions addresses various structured prediction problems. Instead of directly learning $f^{\dagger}$, we adopt a surrogate model $h^{\dagger}:\mathcal{X}\rightarrow\mathcal{Y}$, where $h^{\dagger}(x):=\mathbb{E}[y|x]$ and $y:=\phi(z)$ is a random variable taking values in $\mathcal{Y}$. This reduces the original structured prediction task to the model (1.1). We then reformulate the original structured prediction problem as the following surrogate nonlinear operator learning problem:
$\min_{h:\mathcal{X}\rightarrow\mathcal{Y}}\mathbb{E}[\|h(x)-y\|_{\mathcal{Y}}^{2}].$
We solve this problem using the SGD algorithm presented in this paper, which yields an approximation of $h^{\dagger}$, denoted by $\hat{h}$. During prediction, we use a decoding operator $D$ defined as
$D(h)(\cdot):=\mathop{\arg\min}_{z\in\mathcal{Z}}\left\{\|h(\cdot)-\phi(z)\|_{\mathcal{Y}}\right\}$
for any estimator $h$, as detailed in [13, 4]. Let $\hat{f}=D(\hat{h})$, denote the estimator for $f^{\dagger}$ obtained via the algorithm. The surrogate approach for structured prediction is illustrated in Figure 1.
According to Ciliberto et al. [13], the following properties hold:
  • (1)
    Fisher Consistency: $D(h^{\dagger})=f^{\dagger}$ almost surely.
  • (2)
    Comparison Inequality:
    $\mathcal{R}(\hat{f})-\mathcal{R}(f^{\dagger})\lesssim\left(\mathbb{E}\left[\|\hat{h}(x)-h^{\dagger}(x)\|_{\mathcal{Y}}^{2}\right]\right)^{\frac{1}{2}}.$
Thus, to bound $\mathcal{R}(\hat{f})-\mathcal{R}(f^{\dagger})$, it suffices to bound $\mathbb{E}[\|\hat{h}(x)-h^{\dagger}(x)\|_{\mathcal{Y}}^{2}]=\mathcal{E}(\hat{h})-\mathcal{E}(h^{\dagger})$, as conducted in this paper. This guarantees decay rates of the prediction error under mild assumptions.
In many structured prediction tasks-such as those in natural language processing (e.g., sequence labeling, machine translation) or time series forecasting-data often arrive sequentially or in streams. In such cases, SGD is particularly well-suited, as it allows for incremental model updates with each new data point. This makes it an effective tool for structured prediction in streaming or time-dependent environments.
$\mathcal{X}$$\mathcal{Z}$$\mathcal{Y}$$\hat{h}$$\phi$$\hat{f}=D(\hat{h})$$D$
Figure 1: Surrogate approach for structured prediction
3.3 Combining with PCA Encoder-decoder Framework
In this subsection, we integrate the regularized SGD for solving the nonlinear operator learning problem, as developed in the previous section, with classical PCA to illustrate the adaptability of our approach within the encoder-decoder framework. Here, we only provide the core idea, and the detailed results and proofs will be presented in our future work. In the following, we outline the fundamental setup of the problem and offer essential clarifications regarding the relevant definitions and notations.
Let $\left(\mathcal{X},\langle\cdot\rangle_{\mathcal{X}},\left\|\cdot\right\|_{\mathcal{X}}\right)$ and $\left(\mathcal{Y},\langle\cdot\rangle_{\mathcal{Y}},\left\|\cdot\right\|_{\mathcal{Y}}\right)$ be two real separable Hilbert spaces. Suppose that
$h^{\dagger}:\mathcal{X}\to\mathcal{Y}$ (3.2)
is a potentially nonlinear operator. Given i.i.d. samples $\left\{x_{t},y_{t}\right\}_{t=1}^{T}\sim\rho$, where $y_{t}=h^{\dagger}(x_{t})+\epsilon_{t}$, and $\epsilon_{t}$ denotes centered i.i.d. noise independent of $x_{t}$, our goal is to solve the prediction problem, i.e., to minimize the prediction error $\mathcal{E}(h)=\mathbb{E}_{\rho}\left[\|h(x)-y\|_{\mathcal{Y}}^{2}\right]$ for some estimator $h$. To this end, we apply the PCA technique to project the input and output samples into points onto finite-dimensional Euclidean spaces. We then approximate the mapping between the finite-dimensional space using kernel methods. Below, we briefly review the PCA technique.
The primary function of PCA is to extract the principal features of the data. High-dimensional data often suffers from the curse of dimensionality, and dimensionality reduction—achieved by identifying and retaining the most significant information—serves as an effective remedy. This constitutes the central role of PCA. In our setting, we employ PCA to reduce the samples from an infinite-dimensional space to a finite-dimensional one. The PCA algorithm applied to a random input $x$ in $\mathcal{X}$ seeks to minimize the reconstruction error $\mathbb{E}\left[\left\|(I-P)x\right\|_{\mathcal{X}}^{2}\right]$ over $\Pi_{d_{\mathcal{X}}}$, the set of all orthogonal projections $P$ with rank $d_{\mathcal{X}}$. Given the covariance operator of $x$ defined as $\Sigma_{x}:=\mathbb{E}[x\otimes x]$, there exist eigenvalue-eigenvector pairs $\{\lambda_{i}^{d_{\mathcal{X}}},\phi_{i}^{d_{\mathcal{X}}}\}_{i\geq 1}$ satisfying $\langle\Sigma_{x}\phi_{i}^{d_{\mathcal{X}}},\phi_{j}^{d_{\mathcal{X}}}\rangle_{\mathcal{X}}=\delta_{ij}$ and $\lambda_{1}^{d_{\mathcal{X}}}\geq\lambda_{2}^{d_{\mathcal{X}}}\geq\cdots\geq 0$, where $\delta_{ij}=1$ if $i=j$, otherwise 00. It can be shown that the optimal PCA projection is given by
$P_{d_{\mathcal{X}}}^{\mathcal{X}}=\mathop{\arg\min}_{P\in\Pi_{d_{\mathcal{X}}}}\mathbb{E}\left[\left\|(I-P)x\right\|_{\mathcal{X}}^{2}\right]=\mathcal{D}_{d_{\mathcal{X}}}^{\mathcal{X}}\circ\mathcal{E}_{d_{\mathcal{X}}}^{\mathcal{X}},$
where the encoder $\mathcal{E}_{d_{\mathcal{X}}}^{\mathcal{X}}:\mathcal{X}\to\mathbb{R}^{d_{\mathcal{X}}}$ is defined as
$\mathcal{E}_{d_{\mathcal{X}}}^{\mathcal{X}}(x):=\left(\left\langle x,\phi_{i}^{\mathcal{X}}\right\rangle_{\mathcal{X}}\right)_{i=1}^{d_{\mathcal{X}}},$
and the decoder $\mathcal{D}_{d_{\mathcal{X}}}^{\mathcal{X}}:\mathbb{R}^{d_{\mathcal{X}}}\to\mathcal{X}$ is defined as
$\mathcal{D}_{d_{\mathcal{X}}}^{\mathcal{X}}(\eta):=\sum_{i=1}^{d_{\mathcal{X}}}\eta_{i}\phi_{i}^{\mathcal{X}}=\left(\mathcal{E}_{d_{\mathcal{X}}}^{\mathcal{X}}\right)^{*}(\eta).$
It then follows that $\mathbb{E}\left[\left\|(I-P_{d_{\mathcal{X}}}^{\mathcal{X}})x\right\|_{\mathcal{X}}^{2}\right]=\sum_{i>d_{\mathcal{X}}}\lambda_{i}^{d_{\mathcal{X}}}$, see [32, Theorem 3.8].
In practice, it is usually difficult to obtain $\Sigma_{x}$ directly, so we typically use the empirical covariance operator $\Sigma_{x}^{T}=\frac{1}{T}\sum_{i=1}^{T}x_{i}\otimes x_{i}$ as a substitute, thus deriving the empirical PCA. Following the same procedure, we naturally obtain the empirical encoder ^$\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$ and empirical decoder $\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$. Similarly, by replacing $\Sigma_{x}^{T}$, the input random variable $x$, and the rank $d_{\mathcal{X}}$ with $\Sigma_{y}^{T}:=\frac{1}{T}\sum_{i=1}^{T}y_{i}\otimes y_{i}$, the output random variable $y$ and $d_{\mathcal{Y}}$, respectively, we apply empirical PCA to $y$ in $\mathcal{Y}$ with rank $d_{\mathcal{Y}}$, and obtain the empirical encoder ^$\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$ and empirical decoder $\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$.
$\mathcal{X}$$\mathcal{Y}$$\mathbb{R}^{d_{\mathcal{X}}}$$\mathbb{R}^{d_{\mathcal{Y}}}$${\color[rgb]{0.82,0.01,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.01,0.11}h^{\dagger}}$^$\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$$\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$^$\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$$\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$$f^{\dagger}$
Figure 2: Commutative diagram of PCA enceoder-decoder framework
We now formulate the estimator as $h:=\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ f\circ\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$. It is then natural to choose $f^{\dagger}=\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ h^{\dagger}\circ\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$. This formulation naturally give rise to a commutative diagram, as illustrated in Figure 2. The works [3, 31] represent $f$ using neural networks. In constract, we represent $f$ in an RKHS induced by a matrix-valued kernel $k:\mathbb{R}^{d_{\mathcal{X}}}\times\mathbb{R}^{d_{\mathcal{X}}}\to\mathcal{B}(\mathbb{R}^{d_{\mathcal{Y}}})$. Specifically, we consider kernels of the form $k(u,v)=\phi(\|u-v\|_{\mathbb{R}^{d_{\mathcal{X}}}})$, where $\phi:[0,\infty)\to\mathbb{R}$ is a radial function, such that $k$ is positive definite for $d_{\mathcal{X}}>0$. This property can be equivalently characterized by requiring $\phi$ to be completely monotone. Notable examples satisfying this condition include the inverse multiquadrics $\phi(x)=\left(c^{2}+x^{2}\right)^{-\beta}$ and the Gaussian kernel $\phi(x)=e^{-\alpha|x|^{2}}$ for any $c>0$, $\beta>0$, and $\alpha>0$; see [45]. Let $\mathcal{H}_{k}$ denote the RKHS induced by the matrix-valued kernel $k$. With a slight abuse of notation, we define the prediction error as
$\mathcal{E}(f)=\mathcal{E}(h):=\mathbb{E}\left[\left\|h(x)-y\right\|_{\mathcal{Y}}^{2}\right]=\mathbb{E}\left[\left\|\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ f\circ\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}(x)-y\right\|_{\mathcal{Y}}^{2}\right],\quad\forall f\in\mathcal{H}_{K},$
[…]
$\nabla\mathcal{E}(f)=2\mathbb{E}\left[\phi\left(\left\|\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}(x)-\cdot\right\|\right)\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\left(\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ f\circ\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}x-y\right)\right].$
Based on samples, we derive the regularized SGD iteration with $f_{1}=0$, and
$f_{t+1}=f_{t}-\eta\left(\phi\left(\left\|\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}(x_{t})-\cdot\right\|\right)\left(f_{t}\left(\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}x_{t}\right)-\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}y_{t}\right)+\lambda f_{t}\right),$
where $\eta$ is the step size. This can be interpreted as an SGD scheme based on the samples {^$\{\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}x_{t},\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}y_{t}\}_{t=1}^{T}$ in $\mathcal{H}_{k}$. Accordingly, we define $h_{1}=0$ and
$h_{t+1}=h_{t}-\eta\left(\phi\left(\left\|\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}(x_{t}-\cdot)\right\|\right)\widehat{P}^{\mathcal{Y}}_{d_{\mathcal{Y}}}(h_{t}(x_{t})-y_{t})+\lambda h_{t}\right),$ (3.3)
where $\widehat{P}^{\mathcal{Y}}_{d_{\mathcal{Y}}}:=\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$ is the empirical projection operator.
Under suitable assumptions, we can assert that, $\phi\left(\left\|\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}(x-\cdot)\right\|\right)\widehat{P}^{\mathcal{Y}}_{d_{\mathcal{Y}}}$ converges to $\phi(\|x-\cdot\|_{\mathcal{X}})I_{\mathcal{Y}}$ as $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$ tend to $\infty$. Hence, the SGD iteration (3.3) can be approached by the following scheme with $\widetilde{h}_{1}=0$, and
$\widetilde{h}_{t+1}=\widetilde{h}_{t}-\eta\left(\phi(\|x_{t}-\cdot\|_{\mathcal{X}})(\widetilde{h}_{t}(x_{t})-y_{t})+\lambda\widetilde{h}_{t}\right),$
which is the setting analyzed in this paper.
To summarize, using the PCA encoder-decoder as a concrete example, we see that our analysis can seamlessly align with the encoder-decoder framework. Rigorous proofs will be provided in our future work.
4 Error Decomposition
In this section, we present the error decomposition employed in the convergence analysis of upper bounds. We begin with several useful observations.
For any $H\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$, by the definition of $\mathcal{E}(H)$,
$\mathcal{E}(H)-\mathcal{E}(H^{\dagger})$ $=\mathbb{E}\left[\|y-H\phi(x)\|_{\mathcal{Y}}^{2}\right]-\mathbb{E}\left[\|y-H^{\dagger}\phi(x)\|_{\mathcal{Y}}^{2}\right]$
$=\mathbb{E}\left[\|(H^{\dagger}-H)\phi(x)+\epsilon\|_{\mathcal{Y}}^{2}\right]-\sigma^{2}$
$=\mathbb{E}\left[\|(H-H^{\dagger})\phi(x)\|_{\mathcal{Y}}^{2}\right]+2\mathbb{E}\left[\langle\epsilon,(H^{\dagger}-H)\phi(x)\rangle_{\mathcal{Y}}\right].$
Since $\epsilon$ is a centered noise independent of $x$, we have $\mathbb{E}\left[\langle\epsilon,(H^{\dagger}-H)\phi(x)\rangle_{\mathcal{Y}}\right]=0$. Therefore,
$\mathcal{E}(H)-\mathcal{E}(H^{\dagger})=\mathbb{E}\left[\|(H-H^{\dagger})\phi(x)\|_{\mathcal{Y}}^{2}\right].$
Furthermore, suppose that $\{f_{j}\}_{j\geq 1}$ is an orthonormal basis of the separable Hilbert space $\mathcal{Y}$. We express $(H-H^{\dagger})\phi(x)$ using a Fourier expansion:
$\mathcal{E}(H)-\mathcal{E}(H^{\dagger})$ $=\mathbb{E}\left[\sum_{j\geq 1}\langle(H-H^{\dagger})\phi(x),f_{j}\rangle_{\mathcal{Y}}^{2}\right]$ (4.1)
$=\sum_{j\geq 1}\mathbb{E}\left[\left\langle(H-H^{\dagger})\phi(x)\otimes\phi(x)(H-H^{\dagger})^{*}f_{j},f_{j}\right\rangle_{\mathcal{Y}}\right]$
$=\left\|(H-H^{\dagger})C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}.$
Our goal in this paper is to estimate $\left\|(H-H^{\dagger})C^{\alpha}\right\|_{\mathrm{HS}}^{2}$ for $\alpha=0$ or $1/2$, corresponding respectively to the estimation error and the prediction error, for a given estimator $H$.
We define the regularizing operator as
$H_{\lambda}$ $:=\mathop{\arg\min}_{H\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})}\mathcal{E}(H)+\lambda\|H\|_{\mathrm{HS}}^{2}$ (4.2)
[…]
$H_{\lambda}=H^{\dagger}C(C+\lambda I)^{-1}=S^{\dagger}C^{1+r}(C+\lambda I)^{-1},$ (4.3)
where the final equality follows from Assumption 2.
As introduced in Section 1, we consider two types of step sizes and regularization parameters. Both can be uniformly expressed in the following form:
$\begin{cases}\eta_{t}=\bar{\eta}(t+t_{0})^{-\theta_{1}},&\\ \lambda_{t}=\bar{\lambda}(t+t_{0})^{-\theta_{2}},&\end{cases}$ (4.4)
where $t_{0}\geq 0$, $\eta_{t}$ is the step size with $\theta_{1}\in[0,1)$ and $\lambda_{t}$ denotes the regularization parameter with $\theta_{2}\in[0,1)$. To avoid confusion, we clarify the parameter settings below:
  1. 1.
    The online setting. In this setting, in (4.4) we require that $\theta_{1},\theta_{2}\in(0,1)$, $t_{0}>0$, and $\bar{\eta},\bar{\lambda}>0$ be constants independent of the current iteration $t$.
  2. 2.
    The finite-horizon setting. In this setting, we set $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ for $t=1,2,\dots,T+1$, where $t_{0}=\theta_{1}=\theta_{2}=0$, and $\theta_{3}\in(0,1)$, $\theta_{4}>0$. Unlike the decaying case, here $\bar{\eta}=\bar{\eta}(T)$ and $\bar{\lambda}=\bar{\lambda}(T)$ depend on $T$, while $\eta_{1}$ and $\lambda_{1}$ are constants independent of $T$.
Lemma 4.1.
Let $\{H_{t}\}_{t\geq 1}$ be defined as (2.1). Then, we have
$H_{t+1}-H_{\lambda_{t}}=$ $(H_{t}-H_{\lambda_{t-1}})(I-\eta_{t}(C+\lambda_{t}I))$ (4.5)
$+(H_{\lambda_{t-1}}-H_{\lambda_{t}})(I-\eta_{t}(C+\lambda_{t}I))+\eta_{t}\mathcal{B}_{t},$
where $I$ denotes the identity operator, and $\mathcal{B}_{t}$ is defined by
$\mathcal{B}_{t}=(H_{t}-H^{\dagger})C+(y_{t}-H_{t}\phi(x_{t}))\otimes\phi(x_{t}).$
Moreover, for any $t\in\mathbb{N}_{T}$, it holds that $\mathbb{E}_{z_{t}}[\mathcal{B}_{t}]=0.$
Proof.
From (4.3), we have $H^{\dagger}C=H_{\lambda_{t}}(C+\lambda_{t}I)$. Combining this with the update rule in algorithm (2.1), we obtain the equality in (4.5), which can be directly verified.
Note that $H_{t}$ depends on $z^{t-1}$ and is independent of $z_{t}$. Therefore, we have
$\mathbb{E}_{z_{t}}[\mathcal{B}_{t}]$ $=(H_{t}-H^{\dagger})C+\mathbb{E}_{z_{t}}[(y_{t}-H_{t}\phi(x_{t}))\otimes\phi(x_{t})]$
$=(H_{t}-H^{\dagger})C+\mathbb{E}_{z_{t}}[\left((H^{\dagger}-H_{t})\phi(x_{t})+\epsilon_{t}\right)\otimes\phi(x_{t})].$
Since $\epsilon_{t}$ is a centered noise independent of $x_{t}$, it follows that $\mathbb{E}_{z_{t}}[\epsilon_{t}\otimes\phi(x_{t})]=0$. Hence,
$\mathbb{E}_{z_{t}}[\mathcal{B}_{t}]=(H_{t}-H^{\dagger})C+(H^{\dagger}-H_{t})C=0.$
The proof is then completed. ∎
We set $\lambda_{0}=t_{0}^{-\theta_{2}}$ in the online setting, and $\lambda_{0}=\bar{\lambda}$ in the finite-horizon setting. Let $\prod_{j=T+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))=I$. By applying induction to the equality (4.5), we derive the following key identity used in the error decomposition:
$H_{T+1}-H_{\lambda_{T}}=$ $(H_{T}-H_{\lambda_{T-1}})(I-\eta_{T}(C+\lambda_{T}I))$ (4.6)
$+(H_{\lambda_{T-1}}-H_{\lambda_{T}})(I-\eta_{T}(C+\lambda_{T}I))+\eta_{T}\mathcal{B}_{T}$
$=$ $\cdots$
$=$ $-H_{\lambda_{0}}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)+\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))$
$+\sum_{t=1}^{T}\eta_{t}\mathcal{B}_{t}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I)).$
In the next proposition, we decompose the expectation of the prediction error (when $\alpha=1/2$) and estimation error (when $\alpha=0$), given by $\mathbb{E}_{z^{T}}\left[\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\alpha}\right\|_{\mathrm{HS}}^{2}\right]$, into four terms that can each be estimated individually.
Proposition 4.2.
Let $\{H_{t}\}_{t\in\mathbb{N}_{T}}$ be defined as (2.1). Suppose that Assumption 4 holds with some $c>0$. Then, for any $T\geq 1$ and $0\leq\alpha\leq\frac{1}{2}$, the following inequality holds:
$\mathbb{E}_{z^{T}}\left[\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\alpha}\right\|_{\mathrm{HS}}^{2}\right]\leq\mathcal{T}_{1}+\mathcal{T}_{2}+\mathcal{T}_{3}+\mathcal{T}_{4},$ (4.7)
where
$\begin{split}&\mathcal{T}_{1}:=2\left\|(H_{\lambda_{T}}-H^{\dagger})C^{\alpha}\right\|_{\mathrm{HS}}^{2},\\ &\mathcal{T}_{2}:=6\left\|H_{\lambda_{0}}C^{\alpha}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|^{2}_{\mathrm{HS}},\\ &\mathcal{T}_{3}:=6\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})C^{\alpha}\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|^{2}_{\mathrm{HS}},\\ &\mathcal{T}_{4}:=6\sqrt{c}\sum_{t=1}^{T}\eta_{t}^{2}\left(\sqrt{c}\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}+\sigma^{2}\right)\mathrm{Tr}\left(C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right).\end{split}$ (4.8)
Proof.
Since $(H_{T+1}-H^{\dagger})C^{\alpha}=(H_{T+1}-H_{\lambda_{T}})C^{\alpha}+(H_{\lambda_{T}}-H^{\dagger})C^{\alpha}$, we have
$\mathbb{E}_{z^{T}}\left[\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\alpha}\right\|_{\mathrm{HS}}^{2}\right]$ $=\mathbb{E}_{z^{T}}\left[\left\|(H_{T+1}-H_{\lambda_{T}})C^{\alpha}+(H_{\lambda_{T}}-H^{\dagger})C^{\alpha}\right\|_{\mathrm{HS}}^{2}\right]$
$\leq 2\mathbb{E}_{z^{T}}\left[\left\|(H_{T+1}-H_{\lambda_{T}})C^{\alpha}\right\|_{\mathrm{HS}}^{2}\right]+2\left\|(H_{\lambda_{T}}-H^{\dagger})C^{\alpha}\right\|_{\mathrm{HS}}^{2}.$
We aim to bound $\mathbb{E}_{z^{T}}\left[\left\|(H_{T+1}-H_{\lambda_{T}})C^{\alpha}\right\|_{\mathrm{HS}}^{2}\right]$. From the equality (4.6), it follows that
$(H_{T+1}-H_{\lambda_{T}})C^{\alpha}=$ $-H_{\lambda_{0}}C^{\alpha}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)+\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})C^{\alpha}\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))$
$+\sum_{t=1}^{T}\eta_{t}\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))=:J_{1}+J_{2}+J_{3}.$ (4.9)
Then,
$\mathbb{E}_{z^{T}}\left[\left\|(H_{T+1}-H_{\lambda_{T}})C^{\alpha}\right\|_{\mathrm{HS}}^{2}\right]\leq 3\|J_{1}\|_{\mathrm{HS}}^{2}+3\|J_{2}\|_{\mathrm{HS}}^{2}+3\mathbb{E}_{z^{T}}\left[\|J_{3}\|_{\mathrm{HS}}^{2}\right].$
We express $\mathbb{E}_{z^{T}}\left[\|J_{3}\|^{2}_{\mathrm{HS}}\right]=\mathbb{E}_{z^{T}}\left[\left\|\sum_{t=1}^{T}\eta_{t}\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}^{2}\right]$ as
$\sum_{t=1}^{T}\sum_{t^{\prime}=1}^{T}\eta_{t}\eta_{t^{\prime}}\mathbb{E}_{z^{T}}\left[\left\langle\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I)),\mathcal{B}_{t^{\prime}}C^{\alpha}\prod_{j=t^{\prime}+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\rangle_{\mathrm{HS}}\right].$
Using the property $\mathbb{E}_{z_{t}}[\mathcal{B}_{t}]=0$, for $t>t^{\prime}$, we obtain
$\mathbb{E}_{z^{T}}\left[\left\langle\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I)),\mathcal{B}_{t^{\prime}}C^{\alpha}\prod_{j=t^{\prime}+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\rangle_{\mathrm{HS}}\right]$
$=\mathbb{E}_{z^{t-1}}\mathbb{E}_{z_{t}}\left[\left\langle\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I)),\mathcal{B}_{t^{\prime}}C^{\alpha}\prod_{j=t^{\prime}+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\rangle_{\mathrm{HS}}\right]$
$=\mathbb{E}_{z^{t-1}}\left[\left\langle\mathbb{E}_{z_{t}}\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I)),\mathcal{B}_{t^{\prime}}C^{\alpha}\prod_{j=t^{\prime}+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\rangle_{\mathrm{HS}}\right]=0.$
Similarly, the above equality also holds for $t<t^{\prime}$. Consequently, there holds
$\mathbb{E}_{z^{T}}\left[\|J_{3}\|_{\mathrm{HS}}^{2}\right]=\sum_{t=1}^{T}\mathbb{E}_{z^{T}}\left[\left\|\eta_{t}\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}^{2}\right].$ (4.10)
Using the property $\mathbb{E}_{z_{t}}[\mathcal{B}_{t}]=0$ again, we have
$\mathcal{B}_{t}=-\mathbb{E}_{z_{t}}[(y_{t}-H_{t}\phi(x_{t}))\otimes\phi(x_{t})]+(y_{t}-H_{t}\phi(x_{t}))\otimes\phi(x_{t}).$
Denote $\eta_{t}\left[(y_{t}-H_{t}\phi(x_{t}))\otimes\phi(x_{t})\right]C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))$ by $\mathcal{A}$, then substituting $\mathcal{A}$ into (4.10) yields that
$\begin{gathered}\mathbb{E}_{z^{t}}\left[\left\|\eta_{t}\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}^{2}\right]=\mathbb{E}_{z^{t-1}}\mathbb{E}_{z_{t}}\left[\left\|-\mathbb{E}_{z_{t}}[\mathcal{A}]+\mathcal{A}\right\|_{\mathrm{HS}}^{2}\right]\\ \leq\mathbb{E}_{z^{t}}\left[\left\|\mathcal{A}\right\|_{\mathrm{HS}}^{2}\right]=\mathbb{E}_{z^{t}}\left[\left\|\eta_{t}\left[(y_{t}-H_{t}\phi(x_{t}))\otimes\phi(x_{t})\right]C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}^{2}\right].\end{gathered}$ (4.11)
Take $\{e_{i}\}_{i\geq 1}$ to be an orthonormal basis of Hilbert space $\mathcal{H}_{\mathcal{K}}$. Since $C$ is self-adjoint, by (4.10), (4.11) and the definition of the Hilbert-Schmidt norm, there holds
$\mathbb{E}_{z^{T}}\left[\|J_{3}\|_{\mathrm{HS}}^{2}\right]$ (4.12)
$\leq$ $\sum_{t=1}^{T}\mathbb{E}_{z^{t}}\left[\sum_{i\geq 1}\left\|\eta_{t}\left[(y_{t}-H_{t}\phi(x_{t}))\otimes\phi(x_{t})\right]C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))e_{i}\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}\right]$
$=$ $\sum_{t=1}^{T}\mathbb{E}_{z^{t}}\left[\sum_{i\geq 1}\left\|\eta_{t}(y_{t}-H_{t}\phi(x_{t}))\right\|_{\mathcal{Y}}^{2}\left\langle\phi(x_{t}),C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))e_{i}\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]$
$=$ $\sum_{t=1}^{T}\eta_{t}^{2}\mathbb{E}_{z^{t}}\left[\left\|y_{t}-H_{t}\phi(x_{t})\right\|_{\mathcal{Y}}^{2}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}\right]$
$=$ $\sum_{t=1}^{T}\eta_{t}^{2}\mathbb{E}_{z^{t-1}}\mathbb{E}_{x_{t}}\left[\mathbb{E}_{\epsilon_{t}}\left\|(H^{\dagger}-H_{t})\phi(x_{t})+\epsilon_{t}\right\|_{\mathcal{Y}}^{2}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}\right],$
where we use $y_{t}=H^{\dagger}\phi(x_{t})+\epsilon_{t}$ in the last equality. It is obvious that
$\mathbb{E}_{\epsilon_{t}}\left[\left\|(H^{\dagger}-H_{t})\phi(x_{t})+\epsilon_{t}\right\|_{\mathcal{Y}}^{2}\right]=\|(H_{t}-H^{\dagger})\phi(x_{t})\|_{\mathcal{Y}}^{2}+\sigma^{2},$
where $\sigma^{2}=\mathbb{E}[\|\epsilon\|_{\mathcal{Y}}^{2}]$ is the variance of $\epsilon$. Substitute it back into (4.12) and use the Cauchy-Schwartz inequality. Then we obtain
$\mathbb{E}_{z^{T}}\left[\|J_{3}\|_{\mathrm{HS}}^{2}\right]$ $\leq\sum_{t=1}^{T}\eta_{t}^{2}\mathbb{E}_{z^{t-1}}\mathbb{E}_{x_{t}}\left[\left(\left\|(H^{\dagger}-H_{t})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}+\sigma^{2}\right)\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}\right]$
$\leq\sum_{t=1}^{T}\eta_{t}^{2}\left(\mathbb{E}_{z^{t-1}}\sqrt{\mathbb{E}_{x_{t}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{4}}+\sigma^{2}\right)$
×$\qquad\times\left(\mathbb{E}_{x_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{4}\right)^{1/2}$
$\leq\sqrt{c}\sum_{t=1}^{T}\eta_{t}^{2}\left(\sqrt{c}\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}+\sigma^{2}\right)\mathbb{E}_{x_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2},$
where the last inequality is due to Assumption 4. Since
$\mathbb{E}_{x_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}$ $=\sum_{i\geq 1}\mathbb{E}_{x_{t}}\left\langle C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t}),e_{i}\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}$
$=\sum_{i\geq 1}\left\langle C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}e_{i},e_{i}\right\rangle_{\mathcal{H}_{\mathcal{K}}}$
$=\mathrm{Tr}\left(C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right),$
there holds
$\mathbb{E}_{z^{T}}\left[\|J_{3}\|_{\mathrm{HS}}^{2}\right]\leq\sqrt{c}\sum_{t=1}^{T}\eta_{t}^{2}\left(\sqrt{c}\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}+\sigma^{2}\right)\mathrm{Tr}\left(C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right),$
which finishes our proof. ∎
Hereafter, we refer to $\mathcal{T}_{1}$ as the approximation error, $\mathcal{T}_{2}$ as the initial error, $\mathcal{T}_{3}$ as the drift error, and $\mathcal{T}_{4}$ as the sample error, respectively.
We now present the error decomposition of $\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\alpha}\right\|_{\mathrm{HS}}^{2}$, which serves as for establishing a high-probability upper bound. For any random variable $\mu$ taking values in $\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$, we denote the $L^{\infty}$ norm of $\|\mu\|_{\mathrm{HS}}$ by $\|\mu\|_{L^{\infty}_{\mathrm{HS}}}$.
Proposition 4.3.
Let $\{H_{t}\}_{t\in\mathbb{N}_{T}}$ be defined as (2.1). Suppose that Assumption 5 holds for some $M_{\rho}>0$. Then, for any $T\geq 1$ and $0\leq\alpha\leq\frac{1}{2}$, the quantity $\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\alpha}\right\|_{\mathrm{HS}}^{2}$ admits the decomposition:
$\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\alpha}\right\|_{\mathrm{HS}}^{2}\leq\mathcal{T}_{1}+\mathcal{T}_{2}+\mathcal{T}_{3}+6\left\|\sum_{t=1}^{T}\chi_{t}\right\|^{2}_{\mathrm{HS}},$ (4.13)
where $\mathcal{T}_{1}$, $\mathcal{T}_{2}$, and $\mathcal{T}_{3}$ are defined in (4.8), and $\chi_{t}=\eta_{t}\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))$ satisfies
$\|\chi_{t}\|_{\mathrm{HS}}\leq 2\eta_{t}\kappa\left(M_{\rho}+\kappa\left\|H_{t}\right\|_{L^{\infty}_{\mathrm{HS}}}\right)\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|,\quad\forall t\in\mathbb{N}_{T}.$ (4.14)
Proof.
The proof follows a similar strategy to the previous proposition. As in the proof of Proposition 4.2, we readily obtain
$\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\alpha}\right\|_{\mathrm{HS}}^{2}\leq 2\left\|(H_{T+1}-H_{\lambda_{T}})C^{\alpha}\right\|_{\mathrm{HS}}^{2}+2\left\|(H_{\lambda_{T}}-H^{\dagger})C^{\alpha}\right\|_{\mathrm{HS}}^{2}$
and
$\left\|(H_{T+1}-H_{\lambda_{T}})C^{\alpha}\right\|_{\mathrm{HS}}^{2}\leq 3\|J_{1}\|_{\mathrm{HS}}^{2}+3\|J_{2}\|_{\mathrm{HS}}^{2}+3\|J_{3}\|_{\mathrm{HS}}^{2},$
where $J_{1}$, $J_{2}$, and $J_{3}$ are defined in (4.9). Defining $\chi_{t}=\eta_{t}\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))$, we then have $J_{3}=\sum_{t=1}^{T}\chi_{t}$. Since $\mathcal{B}_{t}$ can be expressed as
$\mathcal{B}_{t}=(y_{t}-H_{t}\phi(x_{t}))\otimes\phi(x_{t})-\mathbb{E}_{z_{t}}\left[(y_{t}-H_{t}\phi(x_{t}))\otimes\phi(x_{t})\right],$
it follow from Assumption 5 that
$\|\mathcal{B}_{t}\|_{\mathrm{HS}}\leq 2\left\|(y_{t}-H_{t}\phi(x_{t}))\otimes\phi(x_{t})\right\|_{L^{\infty}_{\mathrm{HS}}}\leq 2\kappa\left(M_{\rho}+\kappa\left\|H_{t}\right\|_{L^{\infty}_{\mathrm{HS}}}\right).$
Thus,
$\|\chi_{t}\|_{\mathrm{HS}}\leq 2\eta_{t}\kappa\left(M_{\rho}+\kappa\left\|H_{t}\right\|_{L^{\infty}_{\mathrm{HS}}}\right)\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|.$
The proof is then finished. ∎
5 Intermediate Estimates for Error Analysis
In this section, we derive bounds for $\mathcal{T}_{1}$, $\mathcal{T}_{2}$, $\mathcal{T}_{3}$, and $\mathcal{T}_{4}$. The bounds for $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are presented in a unified form encompassing the finite-horizon setting. The term $\mathcal{T}_{3}$ arises exclusively in the online setting and is therefore analyzed only within that context. For $\mathcal{T}_{4}$, we provide separate bounds corresponding to these two settings. These intermediate results play an important role in the subsequent analysis of prediction and estimation errors, both in expectation and with high probability.
5.1 Bounding Approximation Error
We bound $\mathcal{T}_{1}$ in the following proposition.
Proposition 5.1.
Under the Assumption 2 with $r>0$, there exists a constant $c_{1}$ independent of $t_{0}$, $T$, $\bar{\eta}$, and $\bar{\lambda}$, such that
$\mathcal{T}_{1}=2\left\|(H_{\lambda_{T}}-H^{\dagger})C^{\alpha}\right\|_{\mathrm{HS}}^{2}\leq c_{1}\lambda_{T}^{\min\{2(r+\alpha),2\}}.$
Proof.
We know from (4.3) that
$H_{\lambda_{T}}-H^{\dagger}$ $=H^{\dagger}C(C+\lambda_{T}I)^{-1}-H^{\dagger}$
$=-\lambda_{T}H^{\dagger}(C+\lambda_{T}I)^{-1}=-\lambda_{T}S^{\dagger}C^{r}(C+\lambda_{T}I)^{-1},$
where the last identity uses Assumption 2. It then follows that
$\left\|(H_{\lambda_{T}}-H^{\dagger})C^{\alpha}\right\|_{\mathrm{HS}}$ $=\|\lambda_{T}S^{\dagger}C^{r+\alpha}(C+\lambda_{T}I)^{-1}\|_{\mathrm{HS}}$ (5.1)
$\leq\lambda_{T}\|S^{\dagger}\|_{\mathrm{HS}}\left\|C^{r+\alpha}(C+\lambda_{T}I)^{-1}\right\|.$
Since
$\left\|C^{r+\alpha}(C+\lambda_{T}I)^{-1}\right\|\leq\sup_{0\leq x\leq\kappa^{2}}\frac{x^{r+\alpha}}{x+\lambda_{T}}$ (5.2)
$\leq\begin{cases}\kappa^{2(r+\alpha-1)},&\text{ when }r+\alpha\geq 1,\\ (r+\alpha)^{r+\alpha}(1-r-\alpha)^{1-r-\alpha}\lambda_{T}^{r+\alpha-1},&\text{when }r+\alpha<1,\end{cases}$
combining (5.1) with (5.2), there exists a constant $c_{1}$ such that
$\left\|(H_{\lambda_{T}}-H^{\dagger})C^{\alpha}\right\|_{\mathrm{HS}}^{2}\leq c_{1}\lambda_{T}^{\min\{2(r+\alpha),2\}}.$
It is clear that $c_{1}$ is independent of $t_{0}$, $T$, $\bar{\eta}$, and $\bar{\lambda}$, which completes this proof. ∎
5.2 Bounding Initial Error
The following two lemmas will be used repeatedly throughout our analysis.
Lemma 5.2.
Let $\beta>0$ and let $\eta_{t},\lambda_{t}$ be defined as (4.4). Let $l,m$ be integers satisfying $1\leq l\leq m$. Suppose that $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$. Then, the following estimates hold:
  • (1)
    $\left\|C^{\beta}\prod_{t=l}^{m}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|\leq\exp\left\{-\sum_{t=l}^{m}\eta_{t}\lambda_{t}\right\}\frac{2(\kappa^{2\beta}+(\beta/e)^{\beta})}{1+\left(\sum_{t=l}^{m}\eta_{t}\right)^{\beta}}.$
  • (2)
    $\left\|C^{\beta}\prod_{t=l}^{m}\left(I-\eta_{t}(C+\lambda_{t}I)\right)^{2}\right\|\leq\left(\frac{\beta}{2e}\right)^{\beta}\left(\sum_{t=l}^{m}\eta_{t}\right)^{-\beta}\exp\left\{-2\sum_{t=l}^{m}\eta_{t}\lambda_{t}\right\}$.
  • (3)
    $\left\|C^{\beta}\prod_{t=l}^{m}\left(I-\eta_{t}(C+\lambda_{t}I)\right)^{2}\right\|\leq\exp\left\{-2\sum_{t=l}^{m}\eta_{t}\lambda_{t}\right\}\frac{2(\kappa^{2\beta}+(\beta/(2e))^{\beta})}{1+\left(\sum_{t=l}^{m}\eta_{t}\right)^{\beta}}.$
Proof.
Recall that $C$ (defined in Section 2) is self-adjoint and compact. By the definition of the operator norm, we have
$\left\|C^{\beta}\prod_{t=l}^{m}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|$ $\leq\sup_{0\leq x\leq\kappa^{2}}x^{\beta}\prod_{t=l}^{m}\left(1-\eta_{t}(x+\lambda_{t})\right)$ (5.3)
$\leq\sup_{0\leq x\leq\kappa^{2}}x^{\beta}\exp\left\{-\sum_{t=l}^{m}\eta_{t}(x+\lambda_{t})\right\}$
$=\left(\frac{\beta}{e}\right)^{\beta}\left(\sum_{t=l}^{m}\eta_{t}\right)^{-\beta}\exp\left\{-\sum_{t=l}^{m}\eta_{t}\lambda_{t}\right\},$
where the first inequality follows from the fact $1-\eta_{t}(x+\lambda_{t})\geq 0$ for all $t\geq 1$ and $0\leq x\leq\kappa^{2}$, which is ensured by the condition $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$. On the other hand, we also have
$\left\|C^{\beta}\prod_{t=l}^{m}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|\leq\kappa^{2\beta}\prod_{t=l}^{m}(1-\eta_{t}\lambda_{t})\leq\kappa^{2\beta}\exp\left\{-\sum_{t=l}^{m}\eta_{t}\lambda_{t}\right\}.$ (5.4)
Applying the inequality $\min\{a,b\}\leq\frac{2}{1/a+1/b},\forall a,b>0$ and combining (5.3) with (5.4), we obtain
$\left\|C^{\beta}\prod_{t=l}^{m}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|\leq\exp\left\{-\sum_{t=l}^{m}\eta_{t}\lambda_{t}\right\}\frac{2(\kappa^{2\beta}+(\beta/e)^{\beta})}{1+\left(\sum_{t=l}^{m}\eta_{t}\right)^{\beta}}.$
Now, using (5.4) once more, there holds
$\left\|C^{\beta}\prod_{t=l}^{m}\left(I-\eta_{t}(C+\lambda_{t}I)\right)^{2}\right\|$ $=\left\|C^{\beta/2}\prod_{t=l}^{m}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|^{2}$
$\leq\left(\frac{\beta}{2e}\right)^{\beta}\left(\sum_{t=l}^{m}\eta_{t}\right)^{-\beta}\exp\left\{-2\sum_{t=l}^{m}\eta_{t}\lambda_{t}\right\}.$
Moreover, since $\left\|C^{\beta}\prod_{t=l}^{m}\left(I-\eta_{t}(C+\lambda_{t}I)\right)^{2}\right\|\leq\kappa^{2\beta}\exp\left\{-\sum_{t=l}^{m}\eta_{t}\lambda_{t}\right\}$, applying $\min\{a,b\}\leq\frac{2}{1/a+1/b}$ again yields
$\left\|C^{\beta}\prod_{t=l}^{m}\left(I-\eta_{t}(C+\lambda_{t}I)\right)^{2}\right\|\leq\exp\left\{-2\sum_{t=l}^{m}\eta_{t}\lambda_{t}\right\}\frac{2(\kappa^{2\beta}+(\beta/(2e))^{\beta})}{1+\left(\sum_{t=l}^{m}\eta_{t}\right)^{\beta}}.$
This completes the proof. ∎
The next lemma establishes lower bounds for $\sum_{t=l}^{m}\eta_{t}$ and $\sum_{t=1}^{T}\eta_{t}\lambda_{t}$.
Lemma 5.3.
Let $10\leq\theta_{1}<$ 1, $10\leq\theta_{2}<$ 1, and $\eta_{t},\lambda_{t}$ be defined as (4.4). Then the following bounds hold for $1\leq l\leq m$ with $l\in\mathbb{N}$:
  • (1)
    $\sum_{t=l}^{m}\eta_{t}\geq\frac{\bar{\eta}}{1-\theta_{1}}\left[(m+t_{0}+1)^{1-\theta_{1}}-(l+t_{0})^{1-\theta_{1}}\right].$
  • (2)
    $\sum_{t=l}^{m}\eta_{t}\lambda_{t}\geq\begin{cases}\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(m+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(l+t_{0})^{1-\theta_{1}-\theta_{2}}\right],&\text{ when }\theta_{1}+\theta_{2}\not=1,\\ \bar{\eta}\bar{\lambda}\log\left(\frac{m+t_{0}+1}{t_{0}+l}\right),&\text{ when}\theta_{1}+\theta_{2}=1.\end{cases}$
    In particular, when $l=1$ and $m=T$ with $T\geq t_{0}+1$, we have:
  • (3)
    $\sum_{t=1}^{T}\eta_{t}\geq\frac{1-2^{\theta_{1}-1}}{1-\theta_{1}}\bar{\eta}(T+t_{0})^{1-\theta_{1}}.$
  • (4)
    $\sum_{t=1}^{T}\eta_{t}\lambda_{t}\geq\begin{cases}\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}(1-2^{\theta_{1}+\theta_{2}-1})(T+t_{0})^{1-\theta_{1}-\theta_{2}},&\text{ when }0\leq\theta_{1}+\theta_{2}<1,\\ \bar{\eta}\bar{\lambda}\log\left(\frac{T+t_{0}}{t_{0}+1}\right),&\text{ when }\theta_{1}+\theta_{2}=1,\\ \frac{\bar{\eta}\bar{\lambda}}{\theta_{1}+\theta_{2}-1}(1-2^{1-\theta_{1}-\theta_{2}})(t_{0}+1)^{1-\theta_{1}-\theta_{2}},&\text{ when }\theta_{1}+\theta_{2}>1.\end{cases}$
Proof.
We bound the summation $\sum_{t=l}^{m}\eta_{t}$ using
[…]
$=\frac{\bar{\eta}}{1-\theta_{1}}\left[(m+t_{0}+1)^{1-\theta_{1}}-(l+t_{0})^{1-\theta_{1}}\right].$
For the specific case where $l=1$, $m=T$, and $T\geq t_{0}+1$, we obtain
$\sum_{t=1}^{T}\eta_{t}$ $\geq\frac{\bar{\eta}}{1-\theta_{1}}\left[(T+t_{0}+1)^{1-\theta_{1}}-(t_{0}+1)^{1-\theta_{1}}\right]$
$\geq\frac{1-2^{\theta_{1}-1}}{1-\theta_{1}}\bar{\eta}(T+t_{0})^{1-\theta_{1}}.$
Next, we analyze the summation involving $\lambda_{t}$ using the same estimate as before:
[…]
=$\begin{cases}\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(m+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(l+t_{0})^{1-\theta_{1}-\theta_{2}}\right],&\text{ when }\theta_{1}+\theta_{2}\not=1,\\ \bar{\eta}\bar{\lambda}\log\left(\frac{m+t_{0}+1}{t_{0}+l}\right),&\text{ when}\theta_{1}+\theta_{2}=1.\end{cases}$
For the case $l=1$, $m=T$, and $T\geq t_{0}+1$, we obtain:
$\sum_{t=1}^{T}\eta_{t}\lambda_{t}\geq\begin{cases}\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}(1-2^{\theta_{1}+\theta_{2}-1})(T+t_{0})^{1-\theta_{1}-\theta_{2}},&\text{ when }\theta_{1}+\theta_{2}<1,\\ \bar{\eta}\bar{\lambda}\log\left(\frac{T+t_{0}}{t_{0}+1}\right),&\text{ when }\theta_{1}+\theta_{2}=1,\\ \frac{\bar{\eta}\bar{\lambda}}{\theta_{1}+\theta_{2}-1}(1-2^{1-\theta_{1}-\theta_{2}})(t_{0}+1)^{1-\theta_{1}-\theta_{2}},&\text{ when }\theta_{1}+\theta_{2}>1.\end{cases}$
The proof is then finished. ∎
Next, based on the two lemmas above, we provide a unified upper bound for $\mathcal{T}_{2}$ under the following two settings:
  • (1)
    $10<\theta_{1}<$ 1, $10<\theta_{2}<$ 1, and $t_{0}>0$, corresponding to the online setting;
  • (2)
    $\theta_{1}=\theta_{2}=t_{0}=0$, corresponding to the finite-horizon setting.
The bound in both cases is established by the following proposition.
Proposition 5.4.
Suppose that Assumption 2 holds with $S^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ and $r>0$. Then, for any $T\geq t_{0}+1$, $t_{0}\geq 0$, $10\leq\theta_{1}<$ 1, and $10\leq\theta_{2}<$ 1, the quantity
$\mathcal{T}_{2}=6\left\|H_{\lambda_{0}}C^{\alpha}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|_{\mathrm{HS}}^{2}$
admits the following bound:
$\mathcal{T}_{2}\leq c_{2}\bar{\eta}^{-2(r+\alpha)}\begin{cases}(T+t_{0})^{-2(r+\alpha)(1-\theta_{1})}\exp\{-\tau\bar{\eta}\bar{\lambda}(T+t_{0})^{1-\theta_{1}-\theta_{2}}\},&\text{ when }0\leq\theta_{1}+\theta_{2}<1,\\ (t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(T+t_{0})^{-2(r+\alpha)(1-\theta_{1})-2\bar{\eta}\bar{\lambda}},&\text{ when }\theta_{1}+\theta_{2}=1,\\ (T+t_{0})^{-2(r+\alpha)(1-\theta_{1})},&\text{ when }\theta_{1}+\theta_{2}>1,\end{cases}$
where $c_{2}$ and $\tau$ are constants independent of $t_{0}$, $T$, $\bar{\eta}$, and $\bar{\lambda}$.
Proof.
According to equality (4.3), it follows that
$\left\|H_{\lambda_{0}}C^{\alpha}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|_{\mathrm{HS}}^{2}$ $=\left\|S^{\dagger}C^{r+\alpha+1}(C+\lambda_{0}I)^{-1}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|_{\mathrm{HS}}^{2}$
$\leq\|S^{\dagger}\|_{\mathrm{HS}}^{2}\left\|C^{2(r+\alpha)}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)^{2}\right\|.$ (5.5)
By applying (2) in Lemma 5.2 with $\beta=2(r+\alpha)$ and Lemma 5.3 (3), the following inequality holds:
$\left\|H_{\lambda_{0}}C^{\alpha}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|_{\mathrm{HS}}^{2}$ (5.6)
$\leq$ $\|S^{\dagger}\|_{\mathrm{HS}}^{2}\left(\frac{r+\alpha}{e}\right)^{2(r+\alpha)}\left(\sum_{t=1}^{T}\eta_{t}\right)^{-2(r+\alpha)}\exp\left\{-2\sum_{t=1}^{T}\eta_{t}\lambda_{t}\right\}$
$\leq$ $\|S^{\dagger}\|_{\mathrm{HS}}^{2}\left(\frac{(r+\alpha)(1-\theta_{1})}{e(1-2^{\theta_{1}-1})}\right)^{2(r+\alpha)}\bar{\eta}^{-2(r+\alpha)}(T+t_{0})^{-2(r+\alpha)(1-\theta_{1})}\exp\left\{-2\sum_{t=1}^{T}\eta_{t}\lambda_{t}\right\}.$
Next, using (4) in Lemma 5.3, the exponential term can be bounded as:
$\exp\left\{-2\sum_{t=1}^{T}\eta_{t}\lambda_{t}\right\}\leq\begin{cases}\exp\{-\tau\bar{\eta}\bar{\lambda}(T+t_{0})^{1-\theta_{1}-\theta_{2}}\},&\text{ when }0\leq\theta_{1}+\theta_{2}<1,\\ (t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(T+t_{0})^{-2\bar{\eta}\bar{\lambda}},&\text{ when }\theta_{1}+\theta_{2}=1,\\ 1,&\text{ when }\theta_{1}+\theta_{2}>1,\end{cases}$
where $\tau=\frac{2}{1-\theta_{1}-\theta_{2}}(1-2^{\theta_{1}+\theta_{2}-1})$. Therefore, the bound for $\mathcal{T}_{2}$ becomes:
$\mathcal{T}_{2}$ $=6\left\|H_{\lambda_{0}}C^{\alpha}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|_{\mathrm{HS}}^{2}$ (5.7)
$\leq c_{2}\bar{\eta}^{-2(r+\alpha)}\begin{cases}(T+t_{0})^{-2(r+\alpha)(1-\theta_{1})}\exp\{-\tau\bar{\eta}\bar{\lambda}(T+t_{0})^{1-\theta_{1}-\theta_{2}}\},&\text{ when }0\leq\theta_{1}+\theta_{2}<1,\\ (t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(T+t_{0})^{-2(r+\alpha)(1-\theta_{1})-2\bar{\eta}\bar{\lambda}},&\text{ when }\theta_{1}+\theta_{2}=1,\\ (T+t_{0})^{-2(r+\alpha)(1-\theta_{1})},&\text{ when }\theta_{1}+\theta_{2}>1,\end{cases}$
where $c_{2}=6\|S^{\dagger}\|_{\mathrm{HS}}^{2}\left(\frac{(r+\alpha)(1-\theta_{1})}{e(1-2^{\theta_{1}-1})}\right)^{2(r+\alpha)}$ is independent of $t_{0}$, $T$, $\bar{\eta}$, and $\bar{\lambda}$.
The desired result is established and the proof is complete. ∎
5.3 Bounding Drift Error
In the finite-horizon setting, where $\lambda_{t}=\bar{\lambda}$ is fixed depending on $T$, we have $\mathcal{T}_{3}=0$. Therefore, it is sufficient to bound $\mathcal{T}_{3}$ under the regime of decaying step sizes and regularization parameters. In what follows, we focus on the setting where $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1.
Lemma 5.5.
Suppose that Assumption 2 holds with $S^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ and $r>0$ and let $t_{0}\geq 1$. Then, for any $t\geq 1$, the following bound holds:
$\left\|H_{\lambda_{t-1}}-H_{\lambda_{t}}\right\|_{\mathrm{HS}}\leq\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1},$
where $\widetilde{c}_{3}$ is a constant independent of $t_{0}$, $t$, $\bar{\eta}$, and $\bar{\lambda}$.
Proof.
Based on the expression for $H_{\lambda}$ in (4.3) and under Assumption 2, we deduce that
$\left\|H_{\lambda_{t-1}}-H_{\lambda_{t}}\right\|_{\mathrm{HS}}$ $=\left\|H^{\dagger}C(C+\lambda_{t-1}I)^{-1}-H^{\dagger}C(C+\lambda_{t}I)^{-1}\right\|_{\mathrm{HS}}$ (5.8)
$=\lvert\lambda_{t}-\lambda_{t-1}\rvert\left\|S^{\dagger}C^{r+1}(C+\lambda_{t}I)^{-1}(C+\lambda_{t-1}I)^{-1}\right\|_{\mathrm{HS}}$
$\leq\|S^{\dagger}\|_{\mathrm{HS}}\left\lvert\lambda_{t}-\lambda_{t-1}\right\rvert\left\|C^{r}(C+\lambda_{t}I)^{-1}\right\|$
$\leq\|S^{\dagger}\|_{\mathrm{HS}}\bar{\lambda}\left\lvert(t+t_{0}-1)^{-\theta_{2}}-(t+t_{0})^{-\theta_{2}}\right\rvert\begin{cases}\kappa^{2r-2},&\text{ when }r\geq 1,\\ r^{r}(1-r)^{1-r}\lambda_{t}^{r-1},&\text{ when }r<1,\end{cases}$
where the last inequality uses the fact that
$\left\|C^{r}(C+\lambda_{t}I)^{-1}\right\|\leq\sup_{0\leq x\leq\kappa^{2}}\left\{x^{r}(x+\lambda_{t})^{-1}\right\}\leq\begin{cases}\kappa^{2r-2},&\text{ when}r\geq 1,\\ r^{r}(1-r)^{1-r}\lambda_{t}^{r-1},&\text{ when }r<1.\end{cases}$
Applying the mean value theorem, there exists $\xi\in(0,1)$ such that
$\lvert(t+t_{0}-1)^{-\theta_{2}}-(t+t_{0})^{-\theta_{2}}\rvert$ $=\theta_{2}(t+t_{0}-\xi)^{-(\theta_{2}+1)}\leq\theta_{2}(t+t_{0}-1)^{-(\theta_{2}+1)}$
$\leq 2^{\theta_{2}+1}\theta_{2}(t+t_{0})^{-(\theta_{2}+1)},$ (5.9)
where the last inequality uses $t+t_{0}-1\geq(t+t_{0})/2$. Substituting (5.9) into (5.8), we arrive at
$\left\|H_{\lambda_{t-1}}-H_{\lambda_{t}}\right\|_{\mathrm{HS}}$ $\leq 2^{\theta_{2}+1}\theta_{2}\|S^{\dagger}\|_{\mathrm{HS}}\bar{\lambda}(t+t_{0})^{-(\theta_{2}+1)}\begin{cases}\kappa^{2r-2},&\text{ when }r\geq 1,\\ r^{r}(1-r)^{1-r}\lambda_{t}^{r-1},&\text{ when }r<1,\end{cases}$
$\leq\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1},$
where $\widetilde{c}_{3}$ is a constant independent of $t_{0}$, $t$, $\bar{\eta}$, and $\bar{\lambda}$.
The proof is complete. ∎
Now, we derive an error bound for $\mathcal{T}_{3}$ in the case $\alpha=0$, applied to the analysis of estimation error.
Proposition 5.6.
Suppose that Assumption 2 holds with $S^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ and $r>0$. Set $\alpha=0$ in $\mathcal{T}_{3}$ and assume $t_{0}\geq 1$. Let $T\geq t_{0}+1$ when $\theta_{1}+\theta_{2}<1$, and $T\geq 1$ otherwise. Additionally, assume $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$ and $\bar{\eta}\bar{\lambda}>\theta_{2}\min\{r,1\}$. Then, the following bound holds for $\mathcal{T}_{3}$:
$\mathcal{T}_{3}=6\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|^{2}_{\mathrm{HS}}\leq c_{3}\begin{cases}1,&\text{ when }\theta_{1}+\theta_{2}>1,\\ (T+t_{0})^{-2\theta_{2}\min\{r,1\}},&\text{ when }\theta_{1}+\theta_{2}\leq 1,\end{cases}$
where $c_{3}=c_{3}(t_{0},\bar{\lambda},\bar{\eta})$ is a constant independent of $T$.
Proof.
By Lemma 5.5, we deduce that
$\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}$ (5.10)
$\leq$ $\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}\sum_{t=1}^{T}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1}\left\|\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|.$
Since $C$ is self-adjoint and compact, with the operator norm $\|C\|\leq\kappa^{2}$ (see Section 2), it follows that
$\left\|\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|\leq\sup_{0\leq x\leq\kappa^{2}}\prod_{j=t}^{T}(1-\eta_{j}(x+\lambda_{j}))$ (5.11)
$\leq\sup_{0\leq x\leq\kappa^{2}}\left\{\exp\left\{-\sum_{j=t}^{T}\eta_{j}(x+\lambda_{j})\right\}\right\}\leq\exp\left\{-\sum_{j=t}^{T}\eta_{j}\lambda_{j}\right\}.$
Hence, we have the following bound for $\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}$:
$\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}\sum_{t=1}^{T}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1}\exp\left\{-\sum_{j=t}^{T}\eta_{j}\lambda_{j}\right\}.$
From (2) in Lemma 5.3, we know that $\exp\left\{-\sum_{j=t}^{T}\eta_{j}\lambda_{j}\right\}$ is bounded by
$\begin{cases}\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(T+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(t+t_{0})^{1-\theta_{1}-\theta_{2}}\right]\right\},&\text{ when }\theta_{1}+\theta_{2}\not=1,\\ \exp\left\{-\bar{\eta}\bar{\lambda}\log\left(\frac{T+t_{0}+1}{t+t_{0}}\right)\right\},&\text{ when }\theta_{1}+\theta_{2}=1.\end{cases}$
Thus, we obtain the following bound for the exponential term:
$\exp\left\{-\sum_{j=t}^{T}\eta_{j}\lambda_{j}\right\}\leq\begin{cases}\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(T+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(t+t_{0})^{1-\theta_{1}-\theta_{2}}\right]\right\},&\text{ when }\theta_{1}+\theta_{2}<1,\\ (t+t_{0})^{\bar{\eta}\bar{\lambda}}(T+t_{0})^{-\bar{\eta}\bar{\lambda}},&\text{ when }\theta_{1}+\theta_{2}=1,\\ 1,&\text{ when }\theta_{1}+\theta_{2}>1.\end{cases}$
We next consider the three cases corresponding to $\theta_{1}+\theta_{2}>1$, $\theta_{1}+\theta_{2}=1$, and $\theta_{1}+\theta_{2}<1$.
Case 1: When $\theta_{1}+\theta_{2}>1$, we obtain
$\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}$
$\leq$ $\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}\sum_{t=1}^{T}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1}\leq\widetilde{c}_{3}\frac{t_{0}^{-\theta_{2}\min\{r,1\}}}{\theta_{2}\min\{r,1\}}\bar{\lambda}^{\min\{r,1\}}.$
Case 2: When $\theta_{1}+\theta_{2}=1$, we derive the following bound:
$\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}$
$\leq$ $\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}(T+t_{0})^{-\bar{\eta}\bar{\lambda}}\sum_{t=1}^{T}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1+\bar{\eta}\bar{\lambda}}$
$\leq$ $\frac{\widetilde{c}_{3}}{\bar{\eta}\bar{\lambda}-\theta_{2}\min\{r,1\}}\bar{\lambda}^{\min\{r,1\}}(T+t_{0})^{-\theta_{2}\min\{r,1\}},$
where the last inequality uses the condition that $\bar{\eta}\bar{\lambda}>\theta_{2}\min\{r,1\}$.
Case 3: When $\theta_{1}+\theta_{2}<1$, there holds that
$\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}\leq\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}\sum_{t=1}^{T}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1}$
×$\times\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(T+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(t+t_{0})^{1-\theta_{1}-\theta_{2}}\right]\right\}.$
Now, we estimate the summation in the last inequality. Since $T\geq t_{0}+1$, we have $t+t_{0}\leq\frac{3}{4}(T+t_{0}+1)$ when $t\leq\frac{T}{2}$. By splitting the summation into two parts, from $1$ to $T/2$ and from $T/2$ to $T$, we deduce that
$\sum_{t=1}^{T}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1}\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(T+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(t+t_{0})^{1-\theta_{1}-\theta_{2}}\right]\right\}$
$\leq$ $\sum_{t=1}^{T/2}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1}\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left(1-\left(3/4\right)^{1-\theta_{1}-\theta_{2}}\right)(T+t_{0})^{1-\theta_{1}-\theta_{2}}\right\}$
$+\sum_{t=T/2}^{T}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1}$
$\leq$ $\frac{t_{0}^{-\theta_{2}\min\{r,1\}}}{\theta_{2}\min\{r,1\}}\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left(1-\left(3/4\right)^{1-\theta_{1}-\theta_{2}}\right)(T+t_{0})^{1-\theta_{1}-\theta_{2}}\right\}$
$+\frac{4^{\theta_{2}\min\{r,1\}}-1}{\theta_{2}\min\{r,1\}}(T+t_{0})^{-\theta_{2}\min\{r,1\}}.$
Using the fact that for any constants $k,\gamma>0$, there exists a constant $m$ such that $\exp\left\{-k(T+t_{0})^{1-\theta_{1}-\theta_{2}}\right\}\leq m\left(T+t_{0}\right)^{-\gamma}$, we conclude from this that
$\mathcal{T}_{3}$ $=6\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|^{2}_{\mathrm{HS}}$
$\leq c_{3}\begin{cases}1,&\text{ when }\theta_{1}+\theta_{2}>1,\\ (T+t_{0})^{-2\theta_{2}\min\{r,1\}},&\text{ when }\theta_{1}+\theta_{2}\leq 1,\end{cases}$
where $c_{3}$ is a constant independent of $T$.
We then finish the proof. ∎
Remark 2.
We will only use the bound for the case $\theta_{1}+\theta_{2}=1$ in the above proposition, as it provides better convergence rates than the other cases. Note that we cannot guarantee convergence of the estimation error when $\theta_{1}+\theta_{2}>1$. However, convergence of the prediction error is essured when $\theta_{1}+\theta_{2}>1$ with $10<\theta_{1},\theta_{2}<$ 1, and both the prediction and estimation errors converge when $\theta_{1}+\theta_{2}\leq 1$. The proofs for the remaining cases are similar and are omitted here to avoid repetition.
The following proposition plays a key role in deriving upper bounds for the drift and sample errors. Its technical proof is provided in Appendix A.3.
Proposition 5.7.
Let $v>0$, $\theta\in\mathbb{R}$, $t_{0}\geq 1$, and $T\geq t_{0}+1$. The step size $\eta_{t}$ is defined as (4.4). Suppose $\bar{\eta}\bar{\lambda}>\theta-1$ and $\theta_{1}+\theta_{2}=1$. Then,
$\sum_{t=1}^{T}\exp\left\{-\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{(t+t_{0})^{-\theta}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}\leq\delta_{1}\begin{cases}(T+t_{0})^{-\theta+\theta_{1}},&\text{ when }v>1,\\ (T+t_{0})^{-\theta+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{-\theta+1-v(1-\theta_{1})},&\text{ when }v<1,\end{cases}$
where $\delta_{1}=\delta_{1}(\bar{\lambda},\bar{\eta})$ is a constant independent of $T$ and $t_{0}$.
Remark 3.
The inequality in Proposition 5.7 remains valid for all $T\geq 1$, not only when $T\geq t_{0}+1$, provided that the constant $\delta_{1}$ is allowed to depend on $t_{0}$ and is chosen sufficiently large. More specifically, there exists a constant $\delta_{2}=\delta_{2}(t_{0},\bar{\lambda},\bar{\eta})$ such that, for any $T\geq 1$,
$\sum_{t=1}^{T}\exp\left\{-\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{(t+t_{0})^{-\theta}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}\leq\delta_{2}\begin{cases}(T+t_{0})^{-\theta+\theta_{1}},&\text{ when }v>1,\\ (T+t_{0})^{-\theta+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{-\theta+1-v(1-\theta_{1})},&\text{ when }v<1.\end{cases}$
Note that $\delta_{2}$ depends on $t_{0}$, whereas $\delta_{1}$ does not.
The following proposition establishes a bound on $\mathcal{T}_{3}$ for the case $\alpha=\frac{1}{2}$, which is instrumental in analyzing the prediction error.
Proposition 5.8.
Suppose that Assumption 2 holds with $S^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ and $r>0$. Set $\alpha=1/2$ in $\mathcal{T}_{3}$. Let $\theta_{1}+\theta_{2}=1$, $t_{0}\geq 1$ and $T\geq t_{0}+1$. Suppose that $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$ and $\bar{\eta}\bar{\lambda}>\theta_{2}\min\{r,1\}$. Then, there holds
$\mathcal{T}_{3}=6\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})C^{1/2}\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|^{2}_{\mathrm{HS}}\leq c_{3}(T+t_{0})^{-2\theta_{2}\min\{r,1\}+\theta_{1}-1},$ (5.12)
where $c_{3}$ is a constant independent of $T$ and $t_{0}$.
Proof.
Applying Lemma 5.5, we have
$\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})C^{1/2}\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}$
$\leq$ $\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}\sum_{t=1}^{T}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1}\left\|C^{1/2}\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|.$
Using Lemma 5.2 (1) with $\beta=\frac{1}{2}$, the above inequality is further bounded as
$\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})C^{1/2}\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}$
$\leq$ $2(\kappa+(1/2e)^{1/2})\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}\sum_{t=1}^{T}\exp\left\{-\sum_{j=t}^{T}\eta_{j}\lambda_{j}\right\}\frac{(t+t_{0})^{-\theta_{2}\min\{r,1\}-1}}{1+\left(\sum_{j=t}^{T}\eta_{j}\right)^{1/2}}$
$\leq$ $2(\kappa+(1/2e)^{1/2})\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}\delta_{1}(T+t_{0})^{-\theta_{2}\min\{r,1\}-(1-\theta_{1})/2},$
where in the last inequality we use Proposition 5.7 with $\theta=\theta_{2}\min\{r,1\}+1$ and $v=1/2<1$. As a consequence,
$\mathcal{T}_{3}=6\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})C^{1/2}\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|^{2}_{\mathrm{HS}}\leq c_{3}(T+t_{0})^{-2\theta_{2}\min\{r,1\}+\theta_{1}-1},$
where $c_{3}$ is a constant independent of $T$ and $t_{0}$.
The proof is then finished. ∎
5.4 Bounding Sample Error
Let $\mathbb{E}_{z^{0}}[\xi]=\xi$ for any random variable $\xi$. The next proposition applies to the online setting.
Proposition 5.9.
Suppose that Assumption 3 holds with $0<s\leq 1$. If $t_{0}\geq 1$, $T\geq t_{0}+1$, $10<\theta_{1}<$ 1, $10<\theta_{2}<$ 1, $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, and the following condition holds for any $t\in\mathbb{N}_{T}$:
$\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}\left(=\mathbb{E}_{z^{t-1}}\left[\left\|\left(H_{t}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\right]\right)\leq M,$ (5.13)
where $M$ is independent of $T$. Then, the following bound holds for $\mathcal{T}_{4}$:
$\mathcal{T}_{4}$ $=6\sqrt{c}\sum_{t=1}^{T}\eta_{t}^{2}\left(\sqrt{c}\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}+\sigma^{2}\right)\mathrm{Tr}\left(C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right)$
$\leq c_{4}\left(\sqrt{c}M+\sigma^{2}\right)\begin{cases}(T+t_{0})^{-\theta_{1}},&\text{ when }2\alpha>s,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0}),&\text{ when }2\alpha=s,\\ (T+t_{0})^{-(1+s-2\alpha)\theta_{1}+s-2\alpha},&\text{ when }2\alpha<s,\end{cases}$
where $c_{4}=c_{4}(\bar{\lambda},\bar{\eta})$ is a constant independent of $T$, $t_{0}$, and $M$.
Proof.
Assumption 3 on $C$ guarantees that
$\mathrm{Tr}\left(C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right)\leq\mathrm{Tr}\left(C^{s}\right)\left\|C^{1+2\alpha-s}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right\|.$
Then we use (3) in Lemma 5.2 with $\beta=1+2\alpha-s$ to bound the operator norm as
$\sum_{t=1}^{T}\eta_{t}^{2}\left(\sqrt{c}\mathbb{E}_{Z^{t}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}+\sigma^{2}\right)\mathrm{Tr}\left(C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right)$
$\leq$ $\sum_{t=1}^{T}\eta_{t}^{2}\left(\sqrt{c}M+\sigma^{2}\right)\mathrm{Tr}\left(C^{s}\right)\left\|C^{1+2\alpha-s}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right\|$
$\leq$ $2\left(\kappa^{2+4\alpha-2s}+((1+2\alpha-s)/(2e))^{1+2\alpha-s}\right)\left(\sqrt{c}M+\sigma^{2}\right)\mathrm{Tr}\left(C^{s}\right)$
×$\times\sum_{t=1}^{T}\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{\eta_{t}^{2}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{1+2\alpha-s}}.$
Now, applying Proposition 5.7 with $\theta=2\theta_{1}$ and $v=1+2\alpha-s$, we get
$\mathcal{T}_{4}$ $=6\sqrt{c}\sum_{t=1}^{T}\eta_{t}^{2}\left(\sqrt{c}\mathbb{E}_{Z^{t}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}+\sigma^{2}\right)\mathrm{Tr}\left(C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right)$
$\leq c_{4}\left(\sqrt{c}M+\sigma^{2}\right)\begin{cases}(T+t_{0})^{-\theta_{1}},&\text{ when }2\alpha>s,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0}),&\text{ when }2\alpha=s,\\ (T+t_{0})^{-(1+s-2\alpha)\theta_{1}+s-2\alpha},&\text{ when }2\alpha<s,\end{cases}$
where $c_{4}=12\sqrt{c}\left(\kappa^{2+4\alpha-2s}+((1+2\alpha-s)/(2e))^{1+2\alpha-s}\right)\mathrm{Tr}\left(C^{s}\right)\delta_{1}\bar{\eta}^{2}$.
The proof is thus complete. ∎
The next proposition is used to bound $\mathcal{T}_{4}$ in the finite-horizon setting. We define $00^{0}:=$ 0 for convenience.
Proposition 5.10.
Let $v\geq 0$, $\bar{\eta}=\eta_{1}T^{-\theta_{3}}$, $10<\theta_{3}<$ 1, $\theta_{4}>0$, and $\eta_{1}$, $\lambda_{1}$ be constants independent of $T$. Then, there exists a constant $\delta_{3}$, independent of $T$, such that for any $T\geq 2$,
$\sum_{t=0}^{T-1}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}}{1+\left(t\bar{\eta}\right)^{v}}\leq\delta_{3}\begin{cases}T^{v\theta_{3}+(1-v)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }0\leq v<1,\\ T^{\theta_{3}}\log T,&\text{ when }v=1,\\ T^{\theta_{3}},&\text{ when }v>1.\end{cases}$
Proof.
We divide the proof into three cases: $v=0$, $v>0$, and $10<v<$ 1 with $\theta_{3}+\theta_{4}<1$. The third case is an improvement upon the analysis in the second case.
Case 1: $v=0$
We first apply the inequality $1-\exp\{-x\}\geq\exp\{-2\lambda_{1}\eta_{1}\}x$ for $0\leq x\leq 2\lambda_{1}\eta_{1}$, yielding:
$\sum_{t=0}^{T-1}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}}{1+\left(t\bar{\eta}\right)^{v}}$ $=\frac{1-\exp\left\{-2\lambda_{1}\eta_{1}T^{1-\theta_{4}-\theta_{3}})\right\}}{1-\exp\left\{-2\lambda_{1}\eta_{1}T^{-\theta_{4}-\theta_{3}})\right\}}$
$\leq\frac{\exp\{2\lambda_{1}\eta_{1}\}}{2\lambda_{1}\eta_{1}}T^{\theta_{3}+\theta_{4}}\left(1-\exp\left\{-2\lambda_{1}\eta_{1}T^{1-\theta_{4}-\theta_{3}})\right\}\right).$
When $\theta_{3}+\theta_{4}\leq 1$, this term simplifies as
$\sum_{t=0}^{T-1}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}}{1+\left(t\bar{\eta}\right)^{v}}\leq\frac{\exp\{2\lambda_{1}\eta_{1}\}}{2\lambda_{1}\eta_{1}}T^{\theta_{3}+\theta_{4}}.$
When $\theta_{3}+\theta_{4}>1$, we use the inequality $1-\exp\{-x\}\leq x$ to obtain
$\sum_{t=0}^{T-1}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}}{1+\left(t\bar{\eta}\right)^{v}}\leq\exp\{2\lambda_{1}\eta_{1}\}T.$
Case 2: $v>0$
We bound the summation as
[…]
$\leq$ $1+\frac{1}{\bar{\eta}}+\frac{1}{\bar{\eta}}\begin{cases}\frac{\left(\bar{\eta}T\right)^{1-v}}{1-v},&\text{ when }0<v<1,\\ \log(\bar{\eta}T),&\text{ when }v=1,\\ \frac{1}{v-1},&\text{ when }v>1,\end{cases}$
$\leq$ $\delta_{3}\begin{cases}T^{1-v+\theta_{3}v},&\text{ when }0<v<1,\\ T^{\theta_{3}}\log T,&\text{ when }v=1,\\ T^{\theta_{3}},&\text{ when }v>1,\end{cases}$
where $\delta_{3}$ is a constant independent of $T$.
Case 3: $10<v<$ 1 with $\theta_{3}+\theta_{4}<1$
In this case, a more refined estimation can be achieved compared to Case 2. We split the summation into three parts,
$\sum_{t=0}^{T-1}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}}{1+\left(t\bar{\eta}\right)^{v}}$ (5.15)
$\leq$ $1+\sum_{t=1}^{T^{\theta_{3}+\theta_{4}}}\frac{1}{1+\left(t\bar{\eta}\right)^{v}}+\sum_{t=T^{\theta_{3}+\theta_{4}}}^{T}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}}\right\}}{1+\left(t\bar{\eta}\right)^{v}}$
=:absent:=:= : $1+\mathcal{A}_{1}+\mathcal{A}_{2}.$
We estimate $\mathcal{A}_{1}$ in the same manner as in (5.14). Noting that $\bar{\eta}T^{\theta_{3}+\theta_{4}}=\eta_{1}T^{\theta_{4}}$, we obtain
[…]
$\leq\frac{1}{\eta_{1}}T^{\theta_{3}}\left(1+\frac{\left(\eta_{1}T^{\theta_{4}}\right)^{1-v}}{1-v}\right)\leq\frac{1}{\eta_{1}}\left(1+\frac{\eta_{1}^{1-v}}{1-v}\right)T^{\theta_{4}(1-v)+\theta_{3}}.$
Now, we estimate $\mathcal{A}_{2}$. Since $T^{\theta_{3}+\theta_{4}}-1\geq kT^{\theta_{3}+\theta_{4}}$ for $T\geq 2$, where $k=1-2^{-\theta_{4}-\theta_{3}}$, it follows that
$\mathcal{A}_{2}$ $\leq\int_{kT^{\theta_{3}+\theta_{4}}}^{T}\frac{1}{1+\left(t\bar{\eta}\right)^{v}}\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}dt.$
Letting $x=tT^{-\theta_{4}-\theta_{3}}$, we rewrite the above as
[…]
Since the integral is finite and satisfies
[…]
we combine the bounds for $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, and use (5.15), to conclude that
$\sum_{t=0}^{T-1}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}}{1+\left(t\bar{\eta}\right)^{v}}\leq\left(1+\frac{1}{\eta_{1}}\left(1+\frac{\eta_{1}^{1-v}}{1-v}\right)+\eta_{1}^{-v}k^{-v}\frac{\exp\{-2\lambda_{1}\eta_{1}k\}}{2\lambda_{1}\eta_{1}}\right)T^{(1-v)\theta_{4}+\theta_{3}}.$
We then finish the proof. ∎
The following proposition concerns the finite-horizon setting. Before presenting the result, we highlight the following distinctions: in this setting, we have $t_{0}=\theta_{1}=\theta_{2}=0$, the step size is fixed as $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$, the regularization parameter is set to be $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$, where $\eta_{1}$ and $\lambda_{1}$ are constants independent of $T$. This contrasts with the setting above, where $\bar{\eta}$ and $\bar{\lambda}$ are independent of $T$.
Proposition 5.11.
Suppose that Assumption 3 holds with $0<s\leq 1$. Let $\alpha\in[0,\frac{1}{2}]$, and set the parameters $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ with $10<\theta_{3}<$ 1 and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ with $\theta_{4}>0$. Additionally, assume that $T\geq 2$, $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$ and there exists a constant $\widetilde{M}$ independent of $T$, such that for all $t\in\mathbb{N}_{T}$,
$\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}\left(=\mathbb{E}_{z^{t-1}}\left[\left\|\left(H_{t}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\right]\right)\leq\widetilde{M}.$ (5.16)
Recall that
$\mathcal{T}_{4}=6\sqrt{c}\sum_{t=1}^{T}\eta_{t}^{2}\left(\sqrt{c}\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}+\sigma^{2}\right)\mathrm{Tr}\left(C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right).$
Then, the following bound holds for $\mathcal{T}_{4}$:
$\mathcal{T}_{4}\leq\widetilde{c_{4}}\begin{cases}T^{-(1-2\alpha+s)\theta_{3}+(s-2\alpha)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }2\alpha<s\leq 1+2\alpha,\\ T^{-\theta_{3}}\log T,&\text{ when }2\alpha=s,\\ T^{-\theta_{3}},&\text{ when }2\alpha>s,\end{cases}$
where $\widetilde{c_{4}}$ is a constant independent of $T$.
Proof.
Applying the assumed condition (5.16) and Assumption 3, we get
$\mathcal{T}_{4}\leq$ $6\sqrt{c}\left(\sqrt{c}\widetilde{M}+\sigma^{2}\right)\bar{\eta}^{2}\sum_{t=1}^{T}\mathrm{Tr}\left(C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right)$
$\leq$ $6\sqrt{c}\left(\sqrt{c}\widetilde{M}+\sigma^{2}\right)\mathrm{Tr}(C^{s})\bar{\eta}^{2}\sum_{t=1}^{T}\left\|C^{1+2\alpha-s}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right\|.$
If $01+2\alpha-s>$ 0, applying Lemma 5.2 (3) for $11\leq t\leq T-$ 1, we obtain the following estimate, which also holds for $t=T$,
$\left\|C^{1+2\alpha-s}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right\|\leq\exp\left\{-2(T-t)\bar{\lambda}\bar{\eta}\right\}\frac{2(\kappa^{2(1+2\alpha-s)}+(\frac{1+2\alpha-s}{2e})^{1+2\alpha-s})}{1+\left((T-t)\bar{\eta}\right)^{1+2\alpha-s}}.$
If $01+2\alpha-s=$ 0, we derive the bound
$\left\|C^{1+2\alpha-s}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right\|\leq\kappa^{2(1+2\alpha-s)}\exp\left\{-2(T-t)\bar{\lambda}\bar{\eta}\right\}$
for any $1\leq t\leq T$. Thus, based on the above estimates, we obtain the bound for $\mathcal{T}_{4}$:
$\mathcal{T}_{4}\lesssim\bar{\eta}^{2}\sum_{t=0}^{T-1}\frac{\exp\left\{-2t\bar{\lambda}\bar{\eta}\right\}}{1+\left(t\bar{\eta}\right)^{1+2\alpha-s}},$
where we use the notation $\lesssim$ to omit constants independent of $T$ and $t$ for simplicity, indicating an inequality up to a multiplicative constant.
Since $\bar{\eta}=\eta_{1}T^{-\theta_{3}}$, and applying Proposition 5.10 with $v=1+2\alpha-s$, we obtain
$\mathcal{T}_{4}\leq\widetilde{c_{4}}\begin{cases}T^{-(1-2\alpha+s)\theta_{3}+(s-2\alpha)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }2\alpha<s\leq 1+2\alpha,\\ T^{-\theta_{3}}\log T,&\text{ when }2\alpha=s,\\ T^{-\theta_{3}},&\text{ when }2\alpha>s,\end{cases}$
where $\widetilde{c_{4}}:=12(\kappa^{2(1+2\alpha-s)}+(\frac{1+2\alpha-s}{2e})^{1+2\alpha-s})\sqrt{c}\left(\sqrt{c}\widetilde{M}+\sigma^{2}\right)\mathrm{Tr}(C^{s})\eta_{1}^{2}\delta_{3}$ is a constant independent of $T$.
The proof is then complete. ∎
5.5 Key Bounds for Estimating Prediction Error
In this subsection, we establish the key bounds for estimating the prediction error, specifically (5.13) in Proposition 5.9 and (5.16) in Proposition 5.11. The following proposition pertains to the online setting.
Proposition 5.12.
Under Assumption 2, Assumption 3 and Assumption 4, if $\theta_{1}+\theta_{2}=1$, $t_{0}\geq 1$, $\bar{\eta}\bar{\lambda}>\theta_{2}\min\{r,1\}$ and $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, then there exists a constant $M$ independent of $t$, such that
$\mathbb{E}_{z^{t-1}}\left[\left\|\left(H_{t}-H^{\dagger}\right)C^{1/2}\right\|_{\mathrm{HS}}^{2}\right]\leq M,\quad\forall t\geq 1.$ (5.17)
Proof.
The proposition is proved by induction. We have already bounded $\mathcal{T}_{1}$, $\mathcal{T}_{2}$, $\mathcal{T}_{3}$, and $\mathcal{T}_{4}$ through four propositions, where $M$ in Proposition 5.9 will share the same value during the induction process. An important fact we need to be aware of is that the bounds of $\mathcal{T}_{2}$, $\mathcal{T}_{3}$ and $\mathcal{T}_{4}$ require that $t\geq t_{0}+1$ when we bound $\mathbb{E}_{z^{t}}\left[\left\|\left(H_{t+1}-H^{\dagger}\right)C^{1/2}\right\|_{\mathrm{HS}}^{2}\right]$. Hence, we first bound $\mathbb{E}_{z^{t-1}}\left[\left\|\left(H_{t}-H^{\dagger}\right)C^{1/2}\right\|_{\mathrm{HS}}^{2}\right]$ when $t\leq\lfloor t_{0}\rfloor+1$. When $t=1$,
$\mathbb{E}_{z^{0}}\left[\left\|(H_{1}-H^{\dagger})C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\right]\leq\left\|H^{\dagger}C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\leq\kappa^{2}\left\|H^{\dagger}\right\|_{\mathrm{HS}}^{2}.$
Note that $\mathcal{T}_{1}$, $\mathcal{T}_{2}$, and $\mathcal{T}_{3}$ are deterministic and can be regarded as functions of $t$. Define a function $f:\{1,2,\cdots\}\rightarrow\mathbb{R}$ iteratively, as $f(1):=\kappa^{2}\left\|H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ and,
$f(t+1):=$ $\mathcal{T}_{1}(t+1)+\mathcal{T}_{2}(t+1)+\mathcal{T}_{3}(t+1)$
$+6\sqrt{c}\sum_{k=1}^{t}\eta_{k}^{2}\left(\sqrt{c}f(k)+\sigma^{2}\right)\mathrm{Tr}\left(C^{2}\prod_{j=k+1}^{t}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right)$
when $t>1$. Then, by the error decomposition Proposition 4.2, $\mathbb{E}_{z^{t-1}}\left[\left\|\left(H_{t}-H^{\dagger}\right)C^{1/2}\right\|_{\mathrm{HS}}^{2}\right]\leq f(t)$ for any $t\geq 1$. Choose
$M=f(\lfloor t_{0}\rfloor+1)+\frac{c_{1}\bar{\lambda}^{\min\{2r+1,2\}}+c_{2}\bar{\eta}^{-2(r+\alpha)}+c_{3}+c_{4}\sigma^{2}t_{0}^{-\theta_{1}}\log t_{0}}{1-c_{4}\sqrt{c}t_{0}^{-\theta_{1}}\log t_{0}}.$
Then, (5.17) holds for any $t\leq\lfloor t_{0}\rfloor+1$. Suppose (5.17) holds until some $t\geq\lfloor t_{0}\rfloor+1$. For $t+1$, Set $\alpha=1/2$, corresponding to prediction error $\mathbb{E}_{z^{t}}\left[\left\|\left(H_{t+1}-H^{\dagger}\right)C^{1/2}\right\|_{\mathrm{HS}}^{2}\right]$. Since Assumption 3 is satisfied with $s=1$, we set $s=1$ accordingly. Using Proposition 4.2, Proposition 5.1, Proposition 5.4, Proposition 5.8 and Proposition 5.9, we obtain that
$\mathbb{E}_{z^{t}}\left[\left\|\left(H_{t+1}-H^{\dagger}\right)C^{1/2}\right\|_{\mathrm{HS}}^{2}\right]\leq c_{1}\left(\bar{\lambda}(t+t_{0})^{-\theta_{2}}\right)^{\min\{2r+1,2\}}$
$+c_{2}\bar{\eta}^{-2(r+\alpha)}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(t+t_{0})^{-2(r+\alpha)(1-\theta_{1})-2\bar{\eta}\bar{\lambda}}$
$+c_{3}(t+t_{0})^{-2\theta_{2}\min\{r,1\}+\theta_{1}-1}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)(t+t_{0})^{-\theta_{1}}\log(t+t_{0})$
$\leq$ $c_{1}\bar{\lambda}^{\min\{2r+1,2\}}+c_{2}\bar{\eta}^{-2(r+\alpha)}+c_{3}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)(t+t_{0})^{-\theta_{1}}\log(t+t_{0})$
$\leq$ $c_{1}\bar{\lambda}^{\min\{2r+1,2\}}+c_{2}\bar{\eta}^{-2(r+\alpha)}+c_{3}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)t_{0}^{-\theta_{1}}\log t_{0},$
where the last inequality holds when $t_{0}\geq\exp\{\frac{1}{\theta_{1}}\}$. Since $c_{4}$ is independent of $t_{0}$, for sufficiently large $t_{0}$, we have
$c_{4}\sqrt{c}t_{0}^{-\theta_{1}}\log t_{0}<1.$
Recall the definition of $M$, it follows that
$\mathbb{E}_{z^{t}}\left[\left\|\left(H_{t+1}-H^{\dagger}\right)C^{1/2}\right\|_{\mathrm{HS}}^{2}\right]\leq M,$
which advances the induction.
The proof is then complete. ∎
Next, we establish a similar bound for the finite-horizon setting.
Proposition 5.13.
Under Assumption 2 and Assumption 4, if $t_{0}=\theta_{2}=\theta_{1}=0$. Suppose $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$ and
$\eta_{1}<\frac{1}{6c\kappa^{2}\left(1+\frac{1}{2e\theta_{3}}\right)}.$
Then, for any $T\geq 2$, there exists a constant $\widetilde{M}$ independent of $T$, such that
$\mathbb{E}_{z^{t-1}}\left[\left\|\left(H_{t}-H^{\dagger}\right)C^{1/2}\right\|_{\mathrm{HS}}^{2}\right]\leq\widetilde{M},$ (5.18)
for any $t\in\mathbb{N}_{T}$.
Proof.
We prove this proposition by induction. Set.
$\widetilde{M}=\kappa^{2}\|H^{\dagger}\|_{\mathrm{HS}}^{2}+\frac{c_{1}\lambda_{1}^{\min\{2r+1,2\}}+c_{2}\eta_{1}^{-(2r+1)}+6\sqrt{c}\sigma^{2}\kappa^{2}\left(1+\frac{1}{2e\theta_{3}}\right)\eta_{1}}{1-6c\kappa^{2}\left(1+\frac{1}{2e\theta_{3}}\right)\eta_{1}}.$
For $t=1$, it is clear that
$\mathbb{E}_{z^{0}}\left[\left\|(H_{1}-H^{\dagger})L_{C}^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\right]\leq\kappa^{2}\|H^{\dagger}\|_{\mathrm{HS}}^{2}\leq\widetilde{M}.$
Assume that (5.18) holds from $1$ to $t$. We now prove that it also holds for $t+1$. Using Proposition 4.2, Proposition 5.1, Proposition 5.4 with $t_{0}=0$ and $\mathcal{T}_{3}=0$ , we have
$\mathbb{E}_{z^{t}}\left[\left\|\left(H_{t+1}-H^{\dagger}\right)C^{1/2}\right\|_{\mathrm{HS}}^{2}\right]\leq c_{1}\lambda_{1}^{\min\{2r+1,2\}}+c_{2}\eta_{1}^{-(2r+1)}+\mathcal{T}_{4}.$ (5.19)
Note that Proposition 5.11 cannot be used in the induction process, because the current step size $\eta_{t}$ relies on the total number of iterations $T$. Therefore, we re-estimate $\mathcal{T}_{4}$. By the definition of $\mathcal{T}_{4}$ and the induction hypothesis,
$\mathcal{T}_{4}$ $\leq 6\sqrt{c}(\sqrt{c}\widetilde{M}+\sigma^{2})\bar{\eta}^{2}\sum_{i=1}^{t}\mathrm{Tr}\left(C^{2}(I-\bar{\eta}(C+\bar{\lambda}I)^{2(t-i)})\right)$
$\leq 6\sqrt{c}(\sqrt{c}\widetilde{M}+\sigma^{2})\mathrm{Tr}(C)\bar{\eta}^{2}\left(\kappa^{2}+\sum_{i=1}^{t-1}\left\|C(I-\bar{\eta}(C+\bar{\lambda}I)^{2i})\right\|\right)$
$\leq 6\sqrt{c}(\sqrt{c}\widetilde{M}+\sigma^{2})\mathrm{Tr}(C)\bar{\eta}^{2}\left(\kappa^{2}+\frac{1}{2e}\sum_{i=1}^{t-1}\left(\bar{\eta}i\right)^{-1}\exp\{-2\bar{\eta}\bar{\lambda}i\}\right).$
where we have used Lemma 5.2 (2) with $\beta=1$. Using $\eta_{1}\kappa^{2}\leq 1$, $\eta=\eta_{1}T^{-\theta_{3}}$ and $\sum_{i=1}^{t-1}i^{-1}\leq 1+\log T$, it follows that
$\mathcal{T}_{4}$ $\leq 6\sqrt{c}(\sqrt{c}\widetilde{M}+\sigma^{2})\mathrm{Tr}(C)\eta_{1}\left(1+\frac{1}{2e}(1+\log T)\right)T^{-\theta_{3}}$
$\leq 6\sqrt{c}(\sqrt{c}\widetilde{M}+\sigma^{2})\mathrm{Tr}(C)\left(1+\frac{1}{2e\theta_{3}}\right)\eta_{1},$
where we have used the fact that $\sup_{x>0}(1+\log x)x^{-\theta_{3}}=\frac{1}{\theta_{3}}\exp\{\theta_{3}-1\}$. Substituting this into (5.19) yields that
$\mathbb{E}_{z^{t}}\left[\left\|\left(H_{t+1}-H^{\dagger}\right)C^{1/2}\right\|_{\mathrm{HS}}^{2}\right]$ $\leq c_{1}\lambda_{1}^{\min\{2r+1,2\}}+c_{2}\eta_{1}^{-(2r+1)}+6\sqrt{c}(\sqrt{c}\widetilde{M}+\sigma^{2})\kappa^{2}\left(1+\frac{1}{2e\theta_{3}}\right)\eta_{1}$
$\leq\widetilde{M},$
which advances the induction.
The proof is thus complete. ∎
6 Convergence Analysis in Expectation
In this section, we prove the error bounds in expectation provided by Subsection 2.2.
Proof of Theorem 2.3.
Let $\theta_{1}+\theta_{2}=1$ and $\alpha=\frac{1}{2}$. If $T\geq t_{0}+1$, from Proposition 4.2, Proposition 5.1, Proposition 5.4, Proposition 5.8, Proposition 5.9 and Proposition 5.12 with $\alpha=\frac{1}{2}$ and $0<s\leq 1$, there holds
$\mathbb{E}_{z^{T}}[\mathcal{E}(H_{T+1})-\mathcal{E}(H^{\dagger})]\leq$ $c_{1}\left(\bar{\lambda}(T+t_{0})^{-\theta_{2}}\right)^{\min\{2r+1,2\}}$
$+c_{2}\bar{\eta}^{-(2r+1)}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(T+t_{0})^{-(2r+1)(1-\theta_{1})-2\bar{\eta}\bar{\lambda}}$
$+c_{3}(T+t_{0})^{-2\theta_{2}\min\{r,1\}+\theta_{1}-1}$
$+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)\begin{cases}(T+t_{0})^{-\theta_{1}},&\text{ when }s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0}),&\text{ when }s=1.\end{cases}$
We choose $\theta_{1}=\frac{2\min\{r+1/2,1\}}{1+2\min\{r+1/2,1\}}$ and $\theta_{2}=\frac{1}{1+2\min\{r+1/2,1\}}$, then
$\mathbb{E}_{z^{T}}[\mathcal{E}(H_{T+1})-\mathcal{E}(H^{\dagger})]\leq c_{1,1}\begin{cases}(T+t_{0})^{-\frac{2\min\{r+1/2,1\}}{1+2\min\{r+1/2,1\}}}\log(T+t_{0}),&\text{ when }s=1,\\ (T+t_{0})^{-\frac{2\min\{r+1/2,1\}}{1+2\min\{r+1/2,1\}}},&\text{ when }s<1,\end{cases}$ (6.1)
for any $T\geq t_{0}+1$, where the constant $c_{1,1}=c_{1}\bar{\lambda}^{\min\{2r+1,2\}}+c_{2}\bar{\eta}^{-(2r+1)}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}+c_{3}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)$ is independent of $T$. Let $c_{1,1}$ be sufficiently large such that (6.1) holds true for $11\leq T<t_{0}+$ 1.
We then finish the proof. ∎
Proof of Theorem 2.4.
Let $\theta_{1}+\theta_{2}=1$ and $\alpha=0$. If $T\geq t_{0}+1$, from Proposition 4.2, Proposition 5.1, Proposition 5.4, Proposition 5.6, Proposition 5.9 and Proposition 5.12 with $\alpha=0$ and $0<s\leq 1$, there holds
$\mathbb{E}_{z^{T}}[\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}]\leq$ $c_{1}\left(\bar{\lambda}(T+t_{0})^{-\theta_{2}}\right)^{2\min\{r,1\}}$
$+c_{2}\bar{\eta}^{-2r}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(T+t_{0})^{-2r(1-\theta_{1})-2\bar{\eta}\bar{\lambda}}$
$+c_{3}(T+t_{0})^{-2\theta_{2}\min\{r,1\}}$
$+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)(T+t_{0})^{-(1+s)\theta_{1}+s}.$
We choose $\theta_{1}=\frac{s+2\min\{r,1\}}{1+s+2\min\{r,1\}}$ and $\theta_{2}=\frac{1}{1+s+2\min\{r,1\}}$, then
$\mathbb{E}_{z^{T}}[\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}]\leq c_{1,2}(T+t_{0})^{-\frac{2\min\{r,1\}}{1+s+2\min\{r,1\}}},$ (6.2)
for any $T\geq t_{0}+1$, where $c_{1,2}=c_{1}\bar{\lambda}^{2\min\{r,1\}}+c_{2}\bar{\eta}^{-2r}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}+c_{3}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)$ is a constant independent of $T$. Let $c_{1,2}$ be sufficiently large such that (6.2) holds true for $11\leq T<t_{0}+$ 1.
The proof is complete. ∎
Proof of Theorem 2.5.
If the conditions $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$ and
$\eta_{1}<\frac{1}{6c\kappa^{2}\left(1+\frac{e^{\theta}_{3}}{2e{\theta_{3}}}\right)}$
hold. By Proposition 4.2, Proposition 5.1, Proposition 5.4, Proposition 5.11 and Proposition 5.13 with $\alpha=1/2$ and $0<s\leq 1$, we obtain
$\mathbb{E}_{z^{T}}[\mathcal{E}(H_{T+1})-\mathcal{E}(H^{\dagger})]\leq$ $c_{1}\left(\lambda_{1}T^{-\theta_{4}}\right)^{\min\{2r+1,2\}}$
$+c_{2}\eta_{1}^{-(2r+1)}T^{-(2r+1)(1-\theta_{3})}\exp\{-\tau\eta_{1}\lambda_{1}T^{1-\theta_{4}-\theta_{3}}\}$
$+\widetilde{c_{4}}\begin{cases}T^{-\theta_{3}}\log T,&\text{ when}s=1,\\ T^{-\theta_{3}},&\text{ when }s<1.\end{cases}$
We choose $\theta_{3}=\frac{2r+1}{2r+2}$ and $\theta_{4}\geq\frac{2r+1}{(2r+2)\min\{2r+1,2\}}$, then
$\mathbb{E}_{z^{T}}[\mathcal{E}(H_{T+1})-\mathcal{E}(H^{\dagger})]\leq c_{1,3}\begin{cases}T^{-\frac{2r+1}{2r+2}},&\text{ when }s<1,\\ T^{-\frac{2r+1}{2r+2}}\log T,&\text{ when }s=1,\end{cases}$
where $c_{1,3}=c_{1}\lambda_{1}^{\min\{2r+1,2\}}+c_{2}\eta_{1}^{-(2r+1)}+\widetilde{c_{4}}$ is a constant independent of $T$.
The proof is complete. ∎
Proof of Theorem 2.6.
If the conditions $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$ and
$\eta_{1}<\frac{1}{6c\kappa^{2}\left(1+\frac{e^{\theta}_{3}}{2e{\theta_{3}}}\right)}$
hold. Using Proposition 4.2, Proposition 5.1, Proposition 5.4, Proposition 5.11 and Proposition 5.13 with $\alpha=0$ and $0<s\leq 1$, there holds
$\mathbb{E}_{z^{T}}[\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}]\leq$ $c_{1}\left(\lambda_{1}T^{-\theta_{4}}\right)^{\min\{2r,2\}}$ (6.3)
$+c_{2}\eta_{1}^{-2r}T^{-2r(1-\theta_{3})}\exp\{-\tau\eta_{1}\lambda_{1}T^{1-\theta_{4}-\theta_{3}}\}$
$+\widetilde{c_{4}}T^{-(1+s)\theta_{3}+s\min\{1,\theta_{3}+\theta_{4}\}}.$
Choosing $\theta_{3}=\frac{2r+s}{1+2r+s}$ and $\theta_{4}\geq\frac{2r}{(1+2r+s)\min\{2r,2\}}$, then
$\mathbb{E}_{z^{T}}[\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}]\leq c_{1,4}T^{-\frac{2r}{1+2r+s}},$
where $c_{1,4}=c_{1}\lambda_{1}^{\min\{2r,2\}}+c_{2}\eta_{1}^{-2r}+\widetilde{c_{4}}$ is a constant independent of $T$.
The proof is complete. ∎
Remark 4.
If we choose $\theta_{3}+\theta_{4}<1$ in the above two proofs under the constant step size, then $\exp\{-\tau T^{1-\theta_{4}-\theta_{3}}\}=o(T^{-k})$ for any $k>0$. Although the second term on the right-hand side of (6.3) decays faster than any polynomial, the overall learning rate would be slower than that achieved by our results.
7 Convergence Analysis in High Probability
In this section, we derive the high-probability error bounds presented in Subsection 2.3. Our proofs are mainly based on the following proposition. This proposition is from [43, Proposition A.3], and is an extension of [39, Theorem 3.4].
Proposition 7.1.
Let $(\xi_{i})_{i\geq 1}$ be a martingale difference sequence in a Hilbert space, i.e., $\mathbb{E}_{i-1}[\xi_{i}]=0$. Suppose that $\|\xi_{i}\|\leq M_{\xi}$ and $\sum_{i=1}^{t}\mathbb{E}_{i-1}\|\xi_{i}\|^{2}\leq\tau^{2}$ almost surely for some constant $M_{\xi}>0$ and $\tau>0$. Then, for any $\delta\in(0,1)$,with probability at least $1-\delta$, the following inequality holds:
$\sup_{1\leq k\leq t}\left\|\sum_{i=1}^{k}\xi_{i}\right\|\leq 2\left(\frac{M_{\xi}}{3}+\tau\right)\log\frac{2}{\delta}.$
Additionally,
$\sup_{1\leq k\leq t}\left\|\sum_{i=1}^{k}\xi_{i}\right\|^{2}\leq 8\left(M_{\xi}^{2}+\tau^{2}\right)\log^{2}\frac{2}{\delta}.$
By Proposition 4.3, since $\mathcal{T}_{1}$, $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$ have already been bounded in Section 5, our goal is to bound the remaining term $6\left\|\sum_{t=1}^{T}\chi_{t}\right\|^{2}_{\mathrm{HS}}$ with high probability. According to Proposition 7.1, this requires the uniform bound on $\left\|\chi_{t}\right\|_{\mathrm{HS}}$ for $1\leq t\leq T$. Using (4.14), it is sufficient to bound $\|H_{t}\|_{\mathrm{HS}}$. However, $\left\|H_{t}\right\|_{\mathrm{HS}}$ may grow rapidly with increasing $t$. Therefore, we first establish a high-probability bound on $\|H_{t}\|_{\mathrm{HS}}$, which motivates the decomposition of $H_{t}-H^{\dagger}$ into $L_{t}+R_{t}$ as follows.
Let us denote $\phi(x_{t})\otimes\phi(x_{t})$ by $C_{t}$. We define two random processes $(L_{t})_{t\geq 1}$ and $(R_{t})_{t\geq 1}$ recursively by
$L_{1}=-H^{\dagger},\quad R_{1}=\mathbf{0},$
and for any $t\geq 1$,
$\begin{split}L_{t+1}&:=L_{t}\left(I-\eta_{t}(C+\lambda_{t}I)\right)-\eta_{t}\lambda_{t}H^{\dagger},\\ R_{t+1}&:=R_{t}\left(I-\eta_{t}(C_{t}+\lambda_{t}I)\right)+\eta_{t}(y_{t}-H^{\dagger}\phi(x_{t}))\otimes\phi(x_{t})+\eta_{t}L_{t}(C-C_{t}).\end{split}$ (7.1)
Note that for any $t\geq 1$, $L_{t}$ is deterministic, while $R_{t}$ depends on $z^{t-1}$ and is independent of $z_{t}$. Moreover, by induction, one can verify that $L_{t}+R_{t}=H_{t}-H^{\dagger}$ for all $t\geq 1$.
We then provide a bound on $L_{t}$ in Lemma 7.2.
Lemma 7.2.
Suppose that $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$ holds. Then, for any $t\geq 1$,
$\|L_{t}\|_{\mathrm{HS}}\leq\|H^{\dagger}\|_{\mathrm{HS}}.$
Proof.
We prove it by induction. For $t=1$, $\|L_{1}\|_{\mathrm{HS}}=\|H^{\dagger}\|_{\mathrm{HS}}$. Since $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, there holds $1-\eta_{t}(\kappa^{2}+\lambda_{t})\geq 0$ for any $t\geq 1$. Thus,
$\|L_{t+1}\|_{\mathrm{HS}}\leq\|L_{t}\|_{\mathrm{HS}}(1-\eta_{t}\lambda_{t})+\eta_{t}\lambda_{t}\|H^{\dagger}\|_{\mathrm{HS}},$
which implies that $\|L_{t}\|_{\mathrm{HS}}\leq\|H^{\dagger}\|_{\mathrm{HS}}$ for any $t\geq 1$. We then finish the proof. ∎
By the following proposition, we can, with high probability, control the increasing rate of $R_{t}$ in the online setting.
Proposition 7.3.
Under Assumption 5, suppose that $\theta_{1}+\theta_{2}=1$, $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, and $\bar{\eta}\bar{\lambda}\geq\theta_{1}$. Then, with probability at least $1-\delta$, there holds
$\|R_{t}\|_{\mathrm{HS}}\leq d_{2}(t+t_{0})^{\frac{1}{2}-\theta_{1}}\log(t+t_{0})\log\frac{2}{\delta},\quad 1\leq t\leq T,$
where $d_{2}$ is a constant independent of $t$, $T$, and $\delta$.
Proof.
Denote $(y_{t}-H^{\dagger}\phi(x_{t}))\otimes\phi(x_{t})+L_{t}(C-C_{t})$ by $K_{t}$. Then, by applying induction to (7.1), $R_{t+1}$ can be expressed as
$R_{t+1}=\sum_{i=1}^{t}\eta_{i}K_{i}\prod_{j=i+1}^{t}\left(I-\eta_{j}(C_{j}+\lambda_{j}I)\right).$ (7.2)
Since $\mathbb{E}_{z_{i}}[K_{i}]=0$, the sequence $\left(\eta_{i}K_{i}\prod_{j=i+1}^{t}\left(I-\eta_{j}(C_{j}+\lambda_{j}I)\right)\right)_{t\geq i\geq 1}$, when traversed from $i=t$ to $i=1$, forms a martingale difference sequence with respect to the increasing $\sigma$-algebra sequence $\sigma(z_{t})$, $\sigma(z_{t-1},z_{t})$, $\ldots$, $\sigma(z_{j}:j=2,3,\ldots,t)$. We apply Proposition 7.1 to bound $R_{t+1}$ for each $t$ individually.
Using Lemma 7.2 and Assumption 5, we have
$\|K_{t}\|_{\mathrm{HS}}\leq\kappa M_{\rho}+3\kappa^{2}\|H^{\dagger}\|_{\mathrm{HS}}.$ (7.3)
By Lemma 5.3,
$\left\|\prod_{j=i+1}^{t}\left(I-\eta_{j}(C_{j}+\lambda_{j}I)\right)\right\|$ $\leq\exp\left\{-\sum_{j=i+1}^{t}\eta_{j}\lambda_{j}\right\}$ (7.4)
$\leq\left(\frac{t+t_{0}+1}{i+t_{0}+1}\right)^{-\bar{\eta}\bar{\lambda}}\leq 2^{\bar{\eta}\bar{\lambda}}\left(\frac{t+t_{0}}{i+t_{0}}\right)^{-\bar{\eta}\bar{\lambda}}.$
Combining (7.2), (7.3) and (7.4), and using the assumption that $\bar{\eta}\bar{\lambda}\geq\theta_{1}$, we obtain
$\left\|\eta_{i}K_{i}\prod_{j=i+1}^{t}\left(I-\eta_{j}(C_{j}+\lambda_{j}I)\right)\right\|_{\mathrm{HS}}\leq\left(\kappa M_{\rho}+3\kappa^{2}\|H^{\dagger}\|_{\mathrm{HS}}\right)2^{\bar{\eta}\bar{\lambda}}\bar{\eta}(t+t_{0})^{-\theta_{1}},\quad 1\leq i\leq t.$
Moreover,
$\sum_{i=1}^{t}\mathbb{E}_{z_{i}}\left[\left\|\eta_{i}K_{i}\prod_{j=i+1}^{t}\left(I-\eta_{j}(C_{j}+\lambda_{j}I)\right)\right\|_{\mathrm{HS}}^{2}\right]$
$\leq$ $\left(\kappa M_{\rho}+3\kappa^{2}\|H^{\dagger}\|_{\mathrm{HS}}\right)^{2}2^{1+\bar{\eta}\bar{\lambda}}\bar{\eta}^{2}\sum_{i=1}^{t}(i+t_{0})^{-2\theta_{1}}\left(\frac{t+t_{0}}{i+t_{0}}\right)^{-2\bar{\eta}\bar{\lambda}}$
$\leq$ $\left(\kappa M_{\rho}+3\kappa^{2}\|H^{\dagger}\|_{\mathrm{HS}}\right)^{2}2^{1+\bar{\eta}\bar{\lambda}}\frac{\bar{\eta}^{2}}{2\bar{\eta}\bar{\lambda}-2\theta_{1}+1}(t+t_{0})^{1-2\theta_{1}}.$
Then, by Proposition 7.1, with probability at least $1-\delta_{t+1}$,
$\|R_{t+1}\|_{\mathrm{HS}}\leq d_{1}(t+t_{0}+1)^{\frac{1}{2}-\theta_{1}}\log\frac{2}{\delta_{t+1}},$
for some constant $d_{1}$ that is independent of $t$, $\delta_{t+1}$ and $T$.
Now, for any $\delta\in(0,1)$, choose $\delta_{t}=\delta(t+t_{0})^{-2}t_{0}$ for any $1\leq t\leq T$. Then $\sum_{t=1}^{T}\delta_{t}\leq\delta$ and
$\|R_{t}\|_{\mathrm{HS}}\leq d_{2}(t+t_{0})^{\frac{1}{2}-\theta_{1}}\log(t+t_{0})\log\frac{2}{\delta},\quad 1\leq t\leq T,$
where $d_{2}$ is a constant independent of $t$, $T$, and $\delta$.
The proof is complete. ∎
Recall that $\chi_{t}=\eta_{t}\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))$ in Proposition 4.3. Define
$\widetilde{\chi}_{t}:=\chi_{t}\mathbbm{1}_{A_{t}},$
where
$A_{t}:=\left\{\|R_{t}\|_{\mathrm{HS}}\leq d_{2}(t+t_{0})^{\frac{1}{2}-\theta_{1}}\log(t+t_{0})\log\frac{2}{\delta}\right\}.$
Then, $A_{t}$ is independent of $z_{t}$, and $\widetilde{\chi}_{t}$ depends on $z^{t}=\{z_{1},z_{2},\cdots,z_{t}\}$. Moreover, for any $t\geq 1$, we have $\mathbb{E}_{z_{t}}\left[\widetilde{\chi}_{t}\right]=\mathbbm{1}_{A_{t}}\mathbb{E}_{z_{t}}\left[\chi_{t}\right]=0$. By Proposition 7.3,
$\mathbb{P}\left(\widetilde{\chi}_{t}=\chi_{t}\text{ for any }1\leq t\leq T\right)\geq 1-\delta.$
In the next proposition, we provide bounds for $\sup_{1\leq t\leq T}\left\|\widetilde{\chi}_{t}\right\|^{2}_{\mathrm{HS}}$ and $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}$ in preparation for applying Proposition 7.1.
Proposition 7.4.
Under Assumption 3 and 5, suppose that $\theta_{1}+\theta_{2}=1$, $\alpha\in[0,\frac{1}{2}]$, $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, $\bar{\eta}\bar{\lambda}\geq\theta_{1}$ and $\bar{\eta}\bar{\lambda}\geq 2\theta_{1}-\frac{1}{2}$. Then,
  • (1)
    The Hilbert-Schmidt norm of $\widetilde{\chi}_{t}$ is uniformly bounded as follows:
    $\sup_{1\leq t\leq T}\left\|\widetilde{\chi}_{t}\right\|^{2}_{\mathrm{HS}}\leq M_{1}^{2},$
    where
    $M_{1}^{2}:=2d_{3}^{2}(T+t_{0})^{-2\theta_{1}}+2d_{3}^{2}(T+t_{0})^{1-4\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}.$
  • (2)
    The total squared Hilbert-Schmidt norm in expectation is bounded by
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq\tau_{1}^{2},$
    where $\tau_{1}$ is defined as
    $\tau_{1}^{2}=d_{5}\begin{cases}(T+t_{0})^{-\theta_{1}}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta},&1+2\alpha-s>1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0})+(T+t_{0})^{1-3\theta_{1}}\log^{3}(T+t_{0})\log^{2}\frac{2}{\delta},&1+2\alpha-s=1,\\ (T+t_{0})^{s-2\alpha-\theta_{1}(1+s-2\alpha)}+(T+t_{0})^{1+s-2\alpha-\theta_{1}(3+s-2\alpha)}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta},&0\leq 1+2\alpha-s<1,\end{cases}$
    where $d_{3}$ and $d_{5}$ are constants independent of $T$ and $\delta$.
Proof.
  • (1)
    From Proposition 4.3 and $H_{t}=L_{t}+R_{t}+H^{\dagger}$, using Lemma 7.2 yields that
    $\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}\leq$ $2\eta_{t}\kappa\left(M_{\rho}+\kappa\left\|L_{t}+R_{t}+H^{\dagger}\right\|_{L^{\infty}_{\mathrm{HS}}}\right)\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|\mathbbm{1}_{A_{t}}$ (7.5)
    $\leq$ $2\eta_{t}\kappa\left(M_{\rho}+2\kappa\|H^{\dagger}\|_{\mathrm{HS}}+\kappa d_{2}(t+t_{0})^{\frac{1}{2}-\theta_{1}}\log(t+t_{0})\log\frac{2}{\delta}\right)$
    ×$\times\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|.$
    If $\alpha>0$, by Lemma 5.2 (1) and Lemma 5.3, we obtain
    $\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|$ $\leq\exp\left\{-\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{2(\kappa^{2\alpha}+(\alpha/e)^{\alpha})}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{\alpha}}$ (7.6)
    $\leq 2(\kappa^{2\alpha}+(\alpha/e)^{\alpha})\left(\frac{T+t_{0}+1}{t+t_{0}+1}\right)^{-\bar{\eta}\bar{\lambda}}.$
    If $\alpha=0$, then
    $\left\|\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|\leq\exp\left\{-\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\leq\left(\frac{T+t_{0}+1}{t+t_{0}+1}\right)^{-\bar{\eta}\bar{\lambda}}.$ (7.7)
    Substituting (7.6) or (7.7) into (7.5) yields that
    $\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}\leq d_{3}(T+t_{0})^{-\bar{\eta}\bar{\lambda}}(t+t_{0})^{\bar{\eta}\bar{\lambda}-\theta_{1}}\left(1+(t+t_{0})^{\frac{1}{2}-\theta_{1}}\log(t+t_{0})\log\frac{2}{\delta}\right),$
    where $d_{3}$ is a constant independent of $\delta$, $T$, and $t$. If $\bar{\eta}\bar{\lambda}\geq\theta_{1}$ and $\bar{\eta}\bar{\lambda}\geq 2\theta_{1}-\frac{1}{2}$, the right-hand side of the above inequality achieve its maximum within $1\leq t\leq T$ at $t=T$. Therefore, we obtain
    $\sup_{1\leq t\leq T}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}\leq d_{3}(T+t_{0})^{-\theta_{1}}+d_{3}(T+t_{0})^{\frac{1}{2}-2\theta_{1}}\log(T+t_{0})\log\frac{2}{\delta}.$
    Thus,
    $\sup_{1\leq t\leq T}\left\|\widetilde{\chi}_{t}\right\|^{2}_{\mathrm{HS}}\leq 2d_{3}^{2}(T+t_{0})^{-2\theta_{1}}+2d_{3}^{2}(T+t_{0})^{1-4\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}.$
  • (2)
    By the definition of $\widetilde{\chi}_{t}$, we see that
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}$ $=\sum_{t=1}^{T}\eta_{t}^{2}\mathbb{E}_{z_{t}}\left[\left\|\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}^{2}\mathbbm{1}_{A_{t}}\right]$
    $\leq\sum_{t=1}^{T}\eta_{t}^{2}\mathbb{E}_{z_{t}}\left[\left\|\left(y_{t}-H_{t}\phi(x_{t})\right)\otimes\phi(x_{t})C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}^{2}\mathbbm{1}_{A_{t}}\right]$
    $=\sum_{t=1}^{T}\eta_{t}^{2}\mathbb{E}_{z_{t}}\left[\|y_{t}-H_{t}\phi(x_{t})\|_{\mathcal{Y}}^{2}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}\mathbbm{1}_{A_{t}}\right],$
    where the last inequality uses the definition of the Hilbert-Schmidt norm. By $H_{t}=L_{t}+R_{t}+H^{\dagger}$, Lemma 7.2 and the definition of $A_{t}$, it follows that
    $\|y_{t}-H_{t}\phi(x_{t})\|_{\mathcal{Y}}^{2}\mathbbm{1}_{A_{t}}\leq 2M_{\rho}^{2}+2\kappa^{2}\left(8\|H^{\dagger}\|_{\mathrm{HS}}^{2}+2d_{2}^{2}(t+t_{0})^{1-2\theta_{1}}\log^{2}(t+t_{0})\log^{2}\frac{2}{\delta}\right).$
    Then,
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $\sum_{t=1}^{T}\eta_{t}^{2}\left(2M_{\rho}^{2}+2\kappa^{2}\left(8\|H^{\dagger}\|_{\mathrm{HS}}^{2}+2d_{2}^{2}(t+t_{0})^{1-2\theta_{1}}\log^{2}(t+t_{0})\log^{2}\frac{2}{\delta}\right)\right)$ (7.8)
    ×$\times\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}.$
    By the definition of the trace of operators and using Assumption 3 and (3) in Lemma 5.2, it follows that if $02\alpha+1-s>$ 0,
    $\quad\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}$ (7.9)
    $\quad=\mathrm{Tr}\left(C^{2\alpha+1}\prod_{j=t+1}^{T}\left(I-\eta_{j}(C+\lambda_{j}I)\right)^{2}\right)$
    $\quad\leq\mathrm{Tr}\left(C^{s}\right)\left\|C^{2\alpha+1-s}\prod_{j=t+1}^{T}\left(I-\eta_{j}(C+\lambda_{j}I)\right)^{2}\right\|$
    $\quad\leq\mathrm{Tr}\left(C^{s}\right)\frac{2(\kappa^{2(2\alpha+1-s))}+((2\alpha+1-s)/(2e))^{2\alpha+1-s})}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{2\alpha+1-s}}$
    ×$\quad\quad\times\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\},$
    else if $02\alpha+1-s=$ 0,
    $\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}\leq\mathrm{Tr}\left(C^{s}\right)\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}.$ (7.10)
    Substituting (7.9) (or (7.10)) into (7.8), we deduce that there exists a constant $d_{4}$ independent of $\delta$, $t$, and $T$, such that
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $d_{4}\sum_{t=1}^{T}(t+t_{0})^{-2\theta_{1}}\left(1+(t+t_{0})^{1-2\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)$ (7.11)
    ×$\times\frac{\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{2\alpha+1-s}}.$
    Now, we apply Proposition 5.7 and Remark 3 with $v=2\alpha+1-s$ to derive the bound. Let $d_{5}$ denote a constant independent of $T$ and $\delta$.
    Case 1: If $12\alpha+1-s>$ 1, then
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq d_{5}(T+t_{0})^{-\theta_{1}}+d_{5}(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}.$
    Case 2: If $12\alpha+1-s=$ 1, then
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq d_{5}(T+t_{0})^{-\theta_{1}}\log(T+t_{0})+d_{5}(T+t_{0})^{1-3\theta_{1}}\log^{3}(T+t_{0})\log^{2}\frac{2}{\delta}.$
    Case 3: If $10<2\alpha+1-s<$ 1, then
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq d_{5}(T+t_{0})^{s-2\alpha-\theta_{1}(1+s-2\alpha)}+d_{5}(T+t_{0})^{1+s-2\alpha-\theta_{1}(3+s-2\alpha)}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}.$
    Case 4: If $02\alpha+1-s=$ 0. Since
    $\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}$ $=\exp\{-2\bar{\eta}\bar{\lambda}\sum_{j=t+1}^{T}(j+t_{0})^{-1}\}$
    $\leq\left(\frac{T+t_{0}+1}{t+t_{0}+1}\right)^{-2\bar{\eta}\bar{\lambda}},$
    substituting this into (7.11) gives
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}$ $\leq d_{5}(T+t_{0})^{-2\bar{\eta}\bar{\lambda}}\left((T+t_{0})^{2\bar{\eta}\bar{\lambda}-2\theta_{1}+1}+(T+t_{0})^{2\bar{\eta}\bar{\lambda}+2-4\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)$
    $\leq d_{5}\left((T+t_{0})^{1-2\theta_{1}}+(T+t_{0})^{2-4\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right),$
    which is consistent with the bound in Case 3.
The proof is complete. ∎
Since $\widetilde{\chi}_{t}$ is $\sigma\left(z_{1},z_{2},\cdots,z_{t}\right)$ measurable and $\mathbb{E}_{z_{t}}\left[\widetilde{\chi}_{t}\right]=0$, $\left(\widetilde{\chi}_{t}\right)_{1\leq t\leq T}$ is a martingale difference sequence. Based on Proposition 7.1, we derive the high-probability error bounds.
Proof of Theorem 2.7.
By Proposition 4.3 with $\alpha=\frac{1}{2}$,
$\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\leq\mathcal{T}_{1}+\mathcal{T}_{2}+\mathcal{T}_{3}+6\left\|\sum_{t=1}^{T}\chi_{t}\right\|^{2}_{\mathrm{HS}}.$ (7.12)
Using Proposition 7.3, there holds
$\mathbb{P}\left(\widetilde{\chi}_{t}=\chi_{t}\text{ for any }1\leq t\leq T\right)\geq 1-\delta.$ (7.13)
Choose $\theta_{1}=\frac{\min\{2r+1,2\}}{1+\min\{2r+1,2\}}$ and $\theta_{2}=\frac{1}{1+\min\{2r+1,2\}}$. Applying Proposition 5.1, Proposition 5.4 and Proposition 5.8 with $\alpha=\frac{1}{2}$ and $T\geq t_{0}+1$, we have
$\mathcal{T}_{1}\leq c_{1}\lambda_{T}^{\min\{2r+1,2\}}=c_{1}\bar{\lambda}^{\min\{2r+1,2\}}(T+t_{0})^{-\frac{\min\{2r+1,2\}}{1+\min\{2r+1,2\}}},$ (7.14)
$\mathcal{T}_{2}$ $\leq c_{2}\bar{\eta}^{-(2r+1)}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(T+t_{0})^{-(2r+1)(1-\theta_{1})-2\bar{\eta}\bar{\lambda}}$ (7.15)
$\leq c_{2}\bar{\eta}^{-(2r+1)}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(T+t_{0})^{-\frac{\min\{2r+1,2\}}{1+\min\{2r+1,2\}}},$
and
$\mathcal{T}_{3}\leq c_{3}(T+t_{0})^{-2\theta_{2}\min\{r,1\}+\theta_{1}-1}\leq c_{3}(T+t_{0})^{-\frac{\min\{2r+1,2\}}{1+\min\{2r+1,2\}}}.$ (7.16)
Using Proposition 7.1 and Proposition 7.4, we deduce that with probability at least $1-\delta$,
$\left\|\sum_{i=1}^{T}\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq 8\left(M_{1}^{2}+\tau_{1}^{2}\right)\log^{2}\frac{2}{\delta}$
with
$M_{1}^{2}$ $=2d_{3}^{2}(T+t_{0})^{-2\theta_{1}}+2d_{3}^{2}(T+t_{0})^{1-4\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}$
$\leq 2d_{3}^{2}(T+t_{0})^{-\theta_{1}}+2d_{3}^{2}(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta},$
and
$\tau_{1}^{2}$ $=d_{5}\begin{cases}(T+t_{0})^{-\theta_{1}}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta},&\text{ when }s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0})+(T+t_{0})^{1-3\theta_{1}}\log^{3}(T+t_{0})\log^{2}\frac{2}{\delta},&\text{ when }s=1.\\ \end{cases}$
Therefore,
$\left\|\sum_{i=1}^{T}\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $8(2d_{3}^{2}+d_{5})\left((T+t_{0})^{-\theta_{1}}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)$ (7.17)
×$\times\log^{2}\frac{2}{\delta}\begin{cases}1,&\text{ when }s<1,\\ \log(T+t_{0}),&\text{ when }s=1.\end{cases}$
If $T\geq t_{0}+1$, combining (7.12), (7.13), (7.14), (7.15), (7.16), and (7.17), we obtain that there exists some constant $c_{2,1}$ independent of $T$ and $\delta$, such that
$\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\leq c_{2,1}\begin{cases}(T+t_{0})^{-\theta_{1}}\log^{2}\frac{2}{\delta}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{4}\frac{2}{\delta},&s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0})\log^{2}\frac{2}{\delta}+(T+t_{0})^{1-3\theta_{1}}\log^{3}(T+t_{0})\log^{4}\frac{2}{\delta},&s=1.\\ \end{cases}$
holds with probability at least $1-2\delta$.
Since it is easy to verify that $\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}$ for $11\leq T<t_{0}+$ 1 can be bounded uniformly by a constant, we can choose $c_{2,1}$ to be sufficiently large such that the bound holds true for $11\leq T<t_{0}+$ 1.
The proof is complete. ∎
Proof of Corollary 2.8.
For any $t\geq 1$, using Theorem 2.7 with $\delta_{t}=(t+t_{0})^{-2}t_{0}\delta$, then $\sum_{t\geq 1}\delta_{t}\leq\delta$.
When $s<1$,
$\mathcal{E}(h_{t+1})-\mathcal{E}(h^{\dagger})\lesssim(t+t_{0})^{-\theta_{1}}\log^{4}\frac{2}{\delta_{t}}\lesssim(t+t_{0})^{-\theta_{1}}\log^{4}(t+t_{0})\log^{4}\frac{2}{\delta}.$
When $s=1$, similarly,
$\mathcal{E}(h_{t+1})-\mathcal{E}(h^{\dagger})\lesssim(t+t_{0})^{-\theta_{1}}\log^{5}(t+t_{0})\log^{4}\frac{2}{\delta}.$
The proof is complete. ∎
Proof of Theorem 2.9.
By Proposition 4.3 with $\alpha=0$,
$\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}\leq\mathcal{T}_{1}+\mathcal{T}_{2}+\mathcal{T}_{3}+6\left\|\sum_{t=1}^{T}\chi_{t}\right\|^{2}_{\mathrm{HS}}.$ (7.18)
According to Proposition 7.3, we have
$\mathbb{P}\left(\widetilde{\chi}_{t}=\chi_{t}\text{ for any }1\leq t\leq T\right)\geq 1-\delta.$ (7.19)
Applying Proposition 5.1, Proposition 5.4, and Proposition 5.6 with $\alpha=0$ and $T\geq t_{0}+1$, we obtain
$\mathcal{T}_{1}\leq c_{1}\lambda_{T}^{\min\{2r,2\}}=c_{1}\bar{\lambda}^{\min\{2r,2\}}(T+t_{0})^{-\min\{2r,2\}\theta_{2}},$ (7.20)
$\mathcal{T}_{2}$ $\leq c_{2}\bar{\eta}^{-2r}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(T+t_{0})^{-2r(1-\theta_{1})-2\bar{\eta}\bar{\lambda}},$ (7.21)
and
$\mathcal{T}_{3}\leq c_{3}(T+t_{0})^{-2\theta_{2}\min\{r,1\}}.$ (7.22)
Using Proposition 7.1 and Proposition 7.4, we deduce that with probability at least $1-\delta$,
$\left\|\sum_{i=1}^{T}\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq 8\left(M_{1}^{2}+\tau_{1}^{2}\right)\log^{2}\frac{2}{\delta}.$
Additionally, we have
$M_{1}^{2}$ $=2d_{3}^{2}(T+t_{0})^{-2\theta_{1}}+2d_{3}^{2}(T+t_{0})^{1-4\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta},$
and
$\tau_{1}^{2}=d_{5}(T+t_{0})^{s-\theta_{1}(1+s)}+d_{5}(T+t_{0})^{1+s-\theta_{1}(3+s)}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}.$
Thus,
$\left\|\sum_{i=1}^{T}\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $8\left(2d_{3}^{2}+d_{5}\right)\left((T+t_{0})^{s-\theta_{1}(1+s)}+(T+t_{0})^{1+s-\theta_{1}(3+s)}\right.$ (7.23)
×$\left.\times\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}.$
Combining (7.18), (7.19), (7.20), (7.21), (7.22), and (7.23), let $c_{2,2}$ denote a constant independent of $T$ and $\delta$.
  • (1)
    If $2\min\{r,1\}+s\geq 1$, i.e., $2r+s\geq 1$, choose $\theta_{1}=\frac{2\min\{r,1\}+s}{1+2\min\{r,1\}+s}$ and $\theta_{2}=\frac{1}{2\min\{r,1\}+s}$. Then
    $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ $\leq c_{2,2}\left((T+t_{0})^{-\frac{2\min\{r,1\}}{1+2\min\{r,1\}+s}}+(T+t_{0})^{-\frac{4\min\{r,1\}+s-1}{1+2\min\{r,1\}+s}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}$
    $\lesssim(T+t_{0})^{-\frac{2\min\{r,1\}}{1+2\min\{r,1\}+s}}\log^{4}\frac{2}{\delta}$
    holds with probability at least $1-2\delta$.
  • (2)
    If $12\min\{r,1\}+s<$ 1, i.e., $12r+s<$ 1, choose $\theta_{1}=\frac{1+2\min\{r,1\}+s}{3+2\min\{r,1\}+s}=\frac{1+2r+s}{3+2r+s}$ and $\theta_{2}=\frac{2}{3+2\min\{r,1\}+s}=\frac{2}{3+2r+s}$. Then,
    $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ $\leq c_{2,2}(T+t_{0})^{-\frac{4\min\{r,1\}}{3+2\min\{r,1\}+s}}\log^{2}(T+t_{0})\log^{4}\frac{2}{\delta}$
    $=c_{2,2}(T+t_{0})^{-\frac{4r}{3+2r+s}}\log^{2}(T+t_{0})\log^{4}\frac{2}{\delta}$
    holds with probability at least $1-2\delta$.
Since $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ for $11\leq T<t_{0}+$ 1 can be uniformly bounded by a constant, we can choose $c_{2,2}$ to be sufficiently large so that the bound also holds in this case.
The proof is then complete. ∎
Proof of Corollary 2.10.
Using Theorem 2.9 with $\delta_{t}=(t+t_{0})^{-2}t_{0}\delta$, we derive the desired bounds. ∎
Next, we focus on the finite-horizon setting and derive the corresponding high-probability error bounds.
Proposition 7.5.
Under Assumption 5, let $T\geq 2$, $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}=\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ and $\lambda_{t}=\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$. Suppose that $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$, there holds
$\left\|R_{t}\right\|_{\mathrm{HS}}\leq\frac{2d_{6}}{\log 2}T^{\frac{1}{2}-\theta_{3}}\log T\log\frac{2}{\delta},\quad 1\leq t\leq T.$
Proof.
The proof follows the same strategy as in Proposition 7.3. Define $K_{t}=(y_{t}-H^{\dagger}\phi(x_{t}))\otimes\phi(x_{t})+L_{t}(C-C_{t})$, then
$R_{t+1}=\sum_{i=1}^{t}\eta_{i}K_{i}\prod_{j=i+1}^{t}\left(I-\eta_{j}(C_{j}+\lambda_{j}I)\right).$
By (7.3), there holds
$\left\|\eta_{i}K_{i}\prod_{j=i+1}^{t}\left(I-\eta_{j}(C_{j}+\lambda_{j}I)\right)\right\|_{\mathrm{HS}}\leq\eta_{1}\left(\kappa M_{\rho}+3\kappa^{2}\|H^{\dagger}\|_{\mathrm{HS}}\right)T^{-\theta_{3}},$
and
$\sum_{i=1}^{t}\mathbb{E}_{z_{i}}\left[\left\|\eta_{i}K_{i}\prod_{j=i+1}^{t}\left(I-\eta_{j}(C_{j}+\lambda_{j}I)\right)\right\|_{\mathrm{HS}}^{2}\right]\leq\eta_{1}^{2}\left(\kappa M_{\rho}+3\kappa^{2}\|H^{\dagger}\|_{\mathrm{HS}}\right)^{2}T^{1-2\theta_{3}}.$
Then, by Proposition 7.1, for some constant $d_{6}$ independent of $t$, $T$, and $\delta_{t+1}$, with probability at least $1-\delta_{t+1}$, it holds that
$\left\|R_{t+1}\right\|_{\mathrm{HS}}\leq d_{6}T^{\frac{1}{2}-\theta_{3}}\log\frac{2}{\delta_{t+1}}.$
Choosing $\delta_{t}=\frac{\delta}{T}$ for any $1\leq t\leq T$, we obtain
$\left\|R_{t}\right\|_{\mathrm{HS}}\leq\frac{2d_{6}}{\log 2}T^{\frac{1}{2}-\theta_{3}}\log T\log\frac{2}{\delta},\quad 1\leq t\leq T.$
The proof is complete. ∎
We now define
$\overline{\chi}_{t}:=\chi_{t}\mathbbm{1}_{\overline{A}_{t}},$
where
$\overline{A}_{t}:=\left\{\|R_{t}\|_{\mathrm{HS}}\leq\frac{2d_{6}}{\log 2}T^{\frac{1}{2}-\theta_{3}}\log T\log\frac{2}{\delta}\right\}.$
Note that $\overline{A}_{t}$ is independent of $z_{t}$, and for any $t\in\mathbb{N}_{T}$, we have $\mathbb{E}_{z_{t}}\left[\overline{\chi}_{t}\right]=0$. Moreover, by Proposition 7.5,
$\mathbb{P}\left(\overline{\chi}_{t}=\chi_{t}\text{ for any }1\leq t\leq T\right)\geq 1-\delta.$
Proposition 7.6.
Suppose Assumption 3 and 5 hold. Let $\alpha\in[0,\frac{1}{2}]$, $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ with $10<\theta_{3}<$ 1, and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ with $\theta_{4}>0$. Assume $T\geq 2$ and $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$. Then,
  • (1)
    The Hilbert-Schmidt norm of $\overline{\chi}_{t}$ is uniformly bounded as follows:
    $\sup_{1\leq t\leq T}\left\|\overline{\chi}_{t}\right\|^{2}_{\mathrm{HS}}\leq M_{2}^{2},$
    where
    $M_{2}^{2}:=d_{7}T^{-2\theta_{3}}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right).$
  • (2)
    The total squared Hilbert-Schmidt norm in expectation is bounded by
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq\tau_{2}^{2},$
    where $\tau_{2}$ is defined as
    $\tau_{2}^{2}:=d_{8}\delta_{3}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\begin{cases}T^{-(1-2\alpha+s)\theta_{3}+(s-2\alpha)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }2\alpha<s\leq 1+2\alpha,\\ T^{-\theta_{3}}\log T,&\text{ when }2\alpha=s,\\ T^{-\theta_{3}},&\text{ when }2\alpha>s.\end{cases}$
    Here, $d_{7}$ and $d_{8}$ are constants independent of $T$ and $\delta$.
Proof.
  • (1)
    By Proposition 4.3, $H_{t}=L_{t}+R_{t}+H^{\dagger}$ and Lemma 7.2, for any $1\leq t\leq T$, we have
    $\left\|\overline{\chi}_{t}\right\|_{\mathrm{HS}}\leq$ $2\eta_{t}\kappa\left(M_{\rho}+\kappa\left\|L_{t}+R_{t}+H^{\dagger}\right\|_{L^{\infty}_{\mathrm{HS}}}\right)\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|\mathbbm{1}_{\overline{A}_{t}}$ (7.24)
    $\leq$ $2\eta_{1}\kappa^{1+2\alpha}\left(M_{\rho}+2\kappa\|H^{\dagger}\|_{\mathrm{HS}}+\kappa\frac{2d_{6}}{\log 2}T^{\frac{1}{2}-\theta_{3}}\log T\log\frac{2}{\delta}\right)T^{-\theta_{3}},$
    where the last inequality uses $\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|\leq\kappa^{2\alpha}$. Hence, there exists a constant $d_{7}$ independent of $t$, $T$, and $T$, such that
    $\sup_{1\leq t\leq T}\left\|\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq d_{7}T^{-2\theta_{3}}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right).$
  • (2)
    By the definition of $\overline{\chi}_{t}$,
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}$ $=\sum_{t=1}^{T}\eta_{t}^{2}\mathbb{E}_{z_{t}}\left[\left\|\mathcal{B}_{t}C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}^{2}\mathbbm{1}_{\overline{A}_{t}}\right]$
    $\leq\sum_{t=1}^{T}\eta_{t}^{2}\mathbb{E}_{z_{t}}\left[\left\|\left(y_{t}-H_{t}\phi(x_{t})\right)\otimes\phi(x_{t})C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}^{2}\mathbbm{1}_{\overline{A}_{t}}\right]$
    $=\sum_{t=1}^{T}\eta_{t}^{2}\mathbb{E}_{z_{t}}\left[\|y_{t}-H_{t}\phi(x_{t})\|_{\mathcal{Y}}^{2}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}\mathbbm{1}_{\overline{A}_{t}}\right],$
    where the last equality follows from the definition of the Hilbert-Schmidt norm. By Assumption 5, $H_{t}=L_{t}+R_{t}+H^{\dagger}$, Lemma 7.2 and the definition of $\overline{A}_{t}$, we obtain
    $\|y_{t}-H_{t}\phi(x_{t})\|_{\mathcal{Y}}^{2}\mathbbm{1}_{\overline{A}_{t}}\leq 2M_{\rho}^{2}+2\kappa^{2}\left(8\|H^{\dagger}\|_{\mathrm{HS}}^{2}+2\left(\frac{2d_{6}}{\log 2}\right)^{2}T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right).$
    Substituting into the earlier bound yields
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $\sum_{t=1}^{T}\eta_{t}^{2}\left(2M_{\rho}^{2}+2\kappa^{2}\left(8\|H^{\dagger}\|_{\mathrm{HS}}^{2}+2\left(\frac{2d_{6}}{\log 2}\right)^{2}T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\right)$ (7.25)
    ×$\times\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}.$
    By the definition of the trace of operators, Assumption 3, and Lemma 5.2 (3), if $02\alpha+1-s>$ 0, then
    $\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}$ (7.26)
    $=$ $\mathrm{Tr}\left(C^{2\alpha+1}\left(I-\bar{\eta}(C+\bar{\lambda}I)\right)^{2(T-t)}\right)$
    $\leq$ $\mathrm{Tr}\left(C^{s}\right)\left\|C^{2\alpha+1-s}\left(I-\bar{\eta}(C+\bar{\lambda}I)\right)^{2(T-t)}\right\|$
    $\leq$ $\mathrm{Tr}\left(C^{s}\right)\frac{2(\kappa^{2(2\alpha+1-s))}+((2\alpha+1-s)/(2e))^{2\alpha+1-s})}{1+\left((T-t)\bar{\eta}\right)^{2\alpha+1-s}}$
    ×$\times\exp\left\{-2(T-t)\bar{\eta}\bar{\lambda}\right\},$
    and if $02\alpha+1-s=$ 0, then
    $\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}\leq\mathrm{Tr}\left(C^{s}\right)\exp\left\{-2(T-t)\bar{\eta}\bar{\lambda}\right\}.$ (7.27)
    Substituting (7.26) or (7.27) into (7.25) yields
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $d_{8}T^{-2\theta_{3}}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\sum_{t=0}^{T-1}\frac{\exp\left\{-2\eta_{1}\lambda_{1}tT^{-\theta_{4}-\theta_{3}}\right\}}{1+\left(t\bar{\eta}\right)^{2\alpha+1-s}},$
    where $d_{8}$ is a constant independent of $t$, $T$, and $\delta$. Using Proposition 5.10 with $v=1+2\alpha-s$, we obtain
    $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $d_{8}\delta_{3}T^{-2\theta_{3}}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)$
    ×$\times\begin{cases}T^{(1+2\alpha-s)\theta_{3}+(s-2\alpha)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }2\alpha<s\leq 1+2\alpha,\\ T^{\theta_{3}}\log T,&\text{ when }2\alpha=s,\\ T^{\theta_{3}},&\text{ when }2\alpha>s,\end{cases}$
    $=$ $d_{8}\delta_{3}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)$
    ×$\times\begin{cases}T^{-(1-2\alpha+s)\theta_{3}+(s-2\alpha)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }2\alpha<s\leq 1+2\alpha,\\ T^{-\theta_{3}}\log T,&\text{ when }2\alpha=s,\\ T^{-\theta_{3}},&\text{ when }2\alpha>s.\end{cases}$
The proof is complete. ∎
Next, we prove the high-probability bounds for prediction and estimation errors in the finite-horizon setting.
Proof of Theorem 2.11.
By Proposition 4.3 with $\alpha=\frac{1}{2}$, we have
$\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\leq\mathcal{T}_{1}+\mathcal{T}_{2}+\mathcal{T}_{3}+6\left\|\sum_{t=1}^{T}\chi_{t}\right\|^{2}_{\mathrm{HS}},$ (7.28)
where $\mathcal{T}_{3}=0$. Using Proposition 7.5, it holds that
$\mathbb{P}\left(\overline{\chi}_{t}=\chi_{t}\text{ for any }1\leq t\leq T\right)\geq 1-\delta.$ (7.29)
Let $\theta_{3}=\frac{2r+1}{2r+2}$ and choose $\theta_{4}\geq\frac{2r+1}{(2r+2)\min\{2r+1,2\}}$. Applying Proposition 5.1 and Proposition 5.4 with $\alpha=\frac{1}{2}$, $t_{0}=\theta_{1}=\theta_{2}=0$, $\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ and $\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$, we obtain
$\mathcal{T}_{1}\leq c_{1}\lambda_{T}^{\min\{2r+1,2\}}=c_{1}\left(\lambda_{1}T^{-\theta_{4}}\right)^{\min\{2r+1,2\}}\leq c_{1}\lambda_{1}^{\min\{2r+1,2\}}T^{-\frac{2r+1}{2r+2}},$ (7.30)
and
$\mathcal{T}_{2}$ $\leq c_{2}\bar{\eta}^{-(2r+1)}T^{-(2r+1)}\exp\{-\tau\bar{\eta}\bar{\lambda}T\}$ (7.31)
$\leq c_{2}\eta_{1}^{-(2r+1)}T^{-\frac{2r+1}{2r+2}}.$
Using Proposition 7.1 and Proposition 7.6, we conclude that with probability at least $1-\delta$,
$\left\|\sum_{i=1}^{T}\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq 8\left(M_{2}^{2}+\tau_{2}^{2}\right)\log^{2}\frac{2}{\delta}.$
Additionally, for $\alpha=\frac{1}{2}$, we have
$M_{2}^{2}=$ $d_{7}T^{-2\theta_{3}}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)$
$\leq$ $d_{7}\left(T^{-\theta_{3}}+T^{1-3\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right),$
and
$\tau_{2}^{2}=d_{8}\delta_{3}\left(T^{-\theta_{3}}+T^{1-3\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\begin{cases}\log T,&\text{ when }s=1,\\ 1,&\text{ when }s<1.\end{cases}$
Therefore,
$\left\|\sum_{i=1}^{T}\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $8(d_{7}+d_{8}\delta_{3})\left(T^{-\theta_{3}}+T^{1-3\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}\begin{cases}\log T,&\text{ when }s=1,\\ 1,&\text{ when }s<1.\end{cases}$ (7.32)
Combining (7.28), (7.29), (7.30), (7.31), and (7.32), we conclude that there exists a constant $c_{2,3}$ independent of $T$, such that
$\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\leq c_{2,3}\begin{cases}T^{-\theta_{3}}\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{2}T\log^{4}\frac{2}{\delta},&\text{ when }s<1,\\ T^{-\theta_{3}}\log T\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{3}T\log^{4}\frac{2}{\delta},&\text{ when }s=1,\\ \end{cases}$
holds with probability at least $1-2\delta$.
The proof is then complete. ∎
Proof of Theorem 2.12.
By Proposition 4.3 with $\alpha=0$, we have
$\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}\leq\mathcal{T}_{1}+\mathcal{T}_{2}+\mathcal{T}_{3}+6\left\|\sum_{t=1}^{T}\chi_{t}\right\|^{2}_{\mathrm{HS}},$ (7.33)
where $\mathcal{T}_{3}=0$. According to Proposition 7.5, it holds that
$\mathbb{P}\left(\overline{\chi}_{t}=\chi_{t}\text{ for any }1\leq t\leq T\right)\geq 1-\delta.$ (7.34)
Applying Proposition 5.1 and Proposition 5.4 with $\alpha=0$, $t_{0}=\theta_{1}=\theta_{2}=0$, $\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ and $\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$, we obtain
$\mathcal{T}_{1}\leq c_{1}\lambda_{T}^{\min\{2r,2\}}=c_{1}\left(\lambda_{1}T^{-\theta_{4}}\right)^{\min\{2r,2\}},$ (7.35)
and
$\mathcal{T}_{2}$ $\leq c_{2}\bar{\eta}^{-2r}T^{-2r}\exp\{-\tau\bar{\eta}\bar{\lambda}T\}$ (7.36)
$\leq c_{2}\eta_{1}^{-2r}T^{-2r(1-\theta_{3})}.$
Using Proposition 7.1 and Proposition 7.6, we conclude that, with probability at least $1-\delta$,
$\left\|\sum_{i=1}^{T}\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq 8\left(M_{2}^{2}+\tau_{2}^{2}\right)\log^{2}\frac{2}{\delta}.$ (7.37)
Additionally, setting $\alpha=0$, we have
$M_{2}^{2}=$ $d_{7}T^{-2\theta_{3}}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right),$
and
$\tau_{2}^{2}$ $=d_{8}\delta_{3}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)T^{-(1+s)\theta_{3}+s\min\{1,\theta_{3}+\theta_{4}\}}$
$=d_{8}\delta_{3}\left(T^{-(1+s)\theta_{3}+s}+T^{-(3+s)\theta_{3}+1+s}\log^{2}T\log^{2}\frac{2}{\delta}\right)$
when $\theta_{3}+\theta_{4}\geq 1$. Therefore, if $\theta_{3}+\theta_{4}\geq 1$, then
$\left\|\sum_{i=1}^{T}\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq 8(d_{7}+d_{8}\delta_{3})\left(T^{-(1+s)\theta_{3}+s}+T^{-(3+s)\theta_{3}+1+s}\log^{2}T\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}.$ (7.38)
Let $c_{2,4}$ be a constant independent of $T$ and $\delta$. Combining (7.33), (7.34), (7.35), (7.36), (7.37), and (7.38), we obtain the following estimates:
  • (1)
    If $2r+s\geq 1$, choose $\theta_{3}=\frac{2r+s}{1+2r+s}$ and $\theta_{4}\geq\frac{r}{(1+2r+s)\min\{r,1\}}$, then
    $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ $\leq c_{2,4}\left(T^{-\frac{2r}{1+2r+s}}+T^{-\frac{4r+s-1}{1+2r+s}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}$
    $\lesssim T^{-\frac{2r}{1+2r+s}}\log^{4}\frac{2}{\delta}$
    holds with probability at least $1-2\delta$.
  • (2)
    If $12r+s<$ 1, choose $\theta_{3}=\frac{1+2r+s}{3+2r+s}$ and $\theta_{4}\geq\frac{2r}{(3+2r+s)\min\{r,1\}}=\frac{2r}{(3+2r+s)r}$, then
    $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ $\leq c_{2,4}\left(T^{-\frac{1+2r-s}{3+2r+s}}+T^{-\frac{4r}{3+2r+s}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}$
    $\lesssim T^{-\frac{4r}{3+2r+s}}\log^{2}T\log^{4}\frac{2}{\delta}$
    holds with probability at least $1-2\delta$.
The proof is then complete. ∎
Appendix
A.1 Proof of Proposition 2.1
Proof.
Since $W$ is self-adjoint and positive, $W^{1/2}$ is also self-adjoint and positive. Let $\mathcal{H}_{0}=\mathrm{span}\{K(x,\cdot)y=\mathcal{K}(x,\cdot)Wy:x\in\mathcal{X},y\in\mathcal{Y}\}$ and $\mathcal{B}_{0}=\mathrm{span}\{W^{1/2}y\otimes\phi(x):x\in\mathcal{X},y\in\mathcal{Y}\}$. Then, it is clear that $\mathcal{H}_{K}=\overline{\mathcal{H}_{0}}$ and $\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\overline{W^{1/2}\mathcal{Y}})=\overline{\mathcal{B}_{0}}$.
We define the mapping $W_{0}:\mathcal{H}_{0}\rightarrow\mathcal{B}_{0}$ by $\ \sum_{i=1}^{n}\alpha_{i}K(x_{i},\cdot)y_{i}\mapsto\sum_{i=1}^{n}\alpha_{i}W^{1/2}y_{i}\otimes\phi(x_{i})$ for any $n\in\mathbb{N}$, $x_{1},\cdots x_{n}\in\mathcal{X}$ and $y_{1},\cdots y_{n}\in\mathcal{Y}$. We see that $W_{0}$ is well-defined and linear. Moreover, for any $\sum_{i=1}^{n}\alpha_{i}K(x_{i},\cdot)y_{i}=\sum_{i=1}^{n}\alpha_{i}\mathcal{K}(x_{i},\cdot)Wy_{i}\in\mathcal{H}_{0}$, we have
$\left\|W_{0}\left(\sum_{i=1}^{n}\alpha_{i}K(x_{i},\cdot)y_{i}\right)\right\|_{\mathrm{HS}}^{2}$ $=\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}\langle W^{1/2}y_{i}\otimes\phi(x_{i}),W^{1/2}y_{j}\otimes\phi(x_{j})\rangle_{\mathrm{HS}}$
$=\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}\mathcal{K}(x_{i},x_{j})\langle Wy_{i},y_{j}\rangle_{\mathcal{Y}}$
$=\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}\langle K(x_{i},x_{j})y_{i},y_{j}\rangle_{\mathcal{Y}}$
$=\left\|\sum_{i=1}^{n}\alpha_{i}K(x_{i},\cdot)y_{i}\right\|_{\mathcal{H}}^{2}.$
By extending $W_{0}$ to $W:\mathcal{H}\rightarrow\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\overline{W^{1/2}\mathcal{Y}})$, we conclude that $\mathcal{H}$ is isometric to $\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\overline{W^{1/2}\mathcal{Y}})\subset\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$.
Next, we show that $h(x)=(Wh)(\phi(x))$ for all $h\in\mathcal{H}$. For any $y\in\mathcal{Y}$,
$\langle y,h(x)\rangle_{\mathcal{Y}}$ $=\langle K(x,\cdot)y,h\rangle_{\mathcal{H}}=\left\langle W(K(x,\cdot)y),Wh\right\rangle_{\mathrm{HS}}=\langle W^{1/2}y\otimes\phi(x),Wh\rangle_{\mathrm{HS}}$
$=\mathrm{Tr}\left(\left(W^{1/2}y\otimes\phi(x)\right)^{*}(Wh)\right)=\left\langle y,W^{1/2}(Wh)\phi(x)\right\rangle_{\mathcal{Y}},$
where the property $\langle y,h(x)\rangle_{\mathcal{Y}}=\langle K(x,\cdot)y,h\rangle_{\mathcal{H}}$ is used in the first equality. Thus, we conclude that $h(x)=W^{1/2}(Wh)\phi(x)$. The uniqueness of this representation is obvious. This completes the proof. ∎
A.2 Proof of Proposition 2.2
Lemma A.1.
Suppose that $\overline{\phi}=\mathbb{E}\left[\phi(x)\right]$. The moment condition (2.3) holds if
$\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\right]\leq c\left(\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\right)^{2}$ (A.1)
holds for some constant $c>0$.
Proof.
First, using that $\mathbb{E}\left[\phi(x)-\overline{\phi}\right]=0$, we see that
$\left(\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\right)^{2}$ $=\left(\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi}+\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\right)^{2}$ (A.2)
$=\left(\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\right)^{2}+\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}+2\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle^{2}_{\mathcal{H}_{\mathcal{K}}}\right].$
Using $\mathbb{E}\left[\phi(x)-\overline{\phi}\right]=0$ again, it holds that
$\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\right]=$ $\mathbb{E}\left[\left(\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}+\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}+2\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}\right)^{2}\right]$ (A.3)
$=$ $\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\right]+\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}+6\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle^{2}_{\mathcal{H}_{\mathcal{K}}}\right]$
$+4\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi}\right\rangle^{3}_{\mathcal{H}_{\mathcal{K}}}\right].$
By Hölder’s inequality and (2.3), we obtain that
$4\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi}\right\rangle^{3}_{\mathcal{H}_{\mathcal{K}}}\right]$ (A.4)
$\leq$ $4\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}\left(\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi}\right\rangle^{4}_{\mathcal{H}_{\mathcal{K}}}\right]\right)^{3/4}$
$\leq$ $4c^{3/4}\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}\left(\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi}\right\rangle^{2}_{\mathcal{H}_{\mathcal{K}}}\right]\right)^{3/2}$
$\leq$ $2c^{3/4}\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi}\right\rangle^{2}_{\mathcal{H}_{\mathcal{K}}}\right]+2c^{3/4}\left(\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi}\right\rangle^{2}_{\mathcal{H}_{\mathcal{K}}}\right]\right)^{2}.$
Then, taking (A.4) back into (A.3) yields that
$\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\right]\leq$ $\left(c+2c^{3/4}\right)\left(\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi}\right\rangle^{2}_{\mathcal{H}_{\mathcal{K}}}\right]\right)^{2}+\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}$ (A.5)
$+\left(6+2c^{3/4}\right)\left\langle\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle^{2}_{\mathcal{H}_{\mathcal{K}}}\right].$
Combining (A.5) with (A.2), we conclude that
$\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\right]\leq\max\left\{c+2c^{3/4},3+c^{3/4}\right\}\left(\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\right)^{2},$
which completes this proof. ∎
Proof of Proposition 2.2.
Recall that $\Sigma:=\mathbb{E}\left[\left(\phi(x)-\overline{\phi}\right)\otimes\left(\phi(x)-\overline{\phi}\right)\right]$. Since $\Sigma$ is compact and self-adjoint, it admits the spectral decomposition:
$\Sigma=\sum_{k\geq 1}\lambda_{k}\phi_{k}\otimes\phi_{k}.$
We claim that $\phi(x)-\overline{\phi}\in\overline{\mathrm{ran}\left(\Sigma\right)}$ almost surely. To prove this, for any $f\in ker\left(\Sigma\right)$, it holds that
$\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]=\left\langle\Sigma f,f\right\rangle_{\mathcal{H}_{\mathcal{K}}}=0.$
This implies that $\phi(x)-\overline{\phi}\in\ker(\Sigma)^{\perp}=\overline{\mathrm{ran}(\Sigma)}$ almost surely. Therefore, we have
$\phi(x)-\overline{\phi}=\sum_{k\geq 1}\sqrt{\lambda_{k}}\xi_{k}\phi_{k},$
where $\xi_{k}:=\frac{\left\langle\phi(x)-\overline{\phi},\phi_{k}\right\rangle}{\sqrt{\lambda_{k}}}$, $\mathbb{E}[\xi_{k}]=0$ and $\mathbb{E}\left[\|\xi_{k}\|_{\mathcal{H}_{\mathcal{K}}}^{2}\right]=1$ for all $k\geq 1$.
Now, Assume that $\{\xi_{k}\}_{k\geq 1}$ consists of independent random variables. We will show that (2.3) holds if $\left\{\mathbb{E}\left[\xi_{k}^{4}\right]\right\}_{k\geq 1}$ is uniformly bounded, i.e., there exists a constant $C>0$ such that $\mathbb{E}\left[\xi_{k}^{4}\right]\leq C$ for all $k\geq 1$. Since the $\xi_{k}$ are mean-zero and independent, it follows that
$\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\right]=\sum_{k\geq 1}\lambda_{k}^{2}\langle f,\phi_{k}\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\mathbb{E}\left[\xi_{k}^{4}\right]+6\sum_{i\not=j}\lambda_{i}\lambda_{j}\langle f,\phi_{i}\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\langle f,\phi_{j}\rangle_{\mathcal{H}_{\mathcal{K}}}^{2},$
and
$\left(\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\right)^{2}=\sum_{k\geq 1}\lambda_{k}^{2}\langle f,\phi_{k}\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}+2\sum_{i\not=j}\lambda_{i}\lambda_{j}\langle f,\phi_{i}\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\langle f,\phi_{j}\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}.$
Using $\mathbb{E}\left[\xi_{k}^{4}\right]\leq C$, we obtain
$\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\right]\leq\max\{C,\frac{1}{3}\}\left(\mathbb{E}\left[\left\langle\phi(x)-\overline{\phi},f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\right)^{2}.$
By Lemma A.1, there exists a constant $c>0$, such that
$\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\right]\leq c\left(\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\right)^{2}.$
The proof is then complete. ∎
A.3 Proof of Proposition 5.7
In this subsection, our goal is to bound
$\sum_{t=1}^{T}\exp\left\{-\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{(t+t_{0})^{-\theta}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}.$
Lemma A.2.
Let $v>0$, $p\in\mathbb{R}$, $T\geq t_{0}+1$ and $t_{0}\geq 1$. The step size $\eta_{t}$ is set as (4.4). Then, there holds
$\sum_{t=1}^{T/2}\frac{(t+t_{0})^{p}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}\leq\delta^{\prime}\begin{cases}(T+t_{0})^{-(1-\theta_{1})v+p+1},&\text{ when }p>-1,\\ (T+t_{0})^{-(1-\theta_{1})v}\log(T+t_{0}),&\text{ when }p=-1,\\ (T+t_{0})^{-(1-\theta_{1})v},&\text{ when }p<-1,\end{cases}$
where $\delta^{\prime}$ is a constant independent of $T$ and $t_{0}$.
Proof.
Since $\sum_{j=t+1}^{T}\eta_{j}\geq\bar{\eta}\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]$, it follows that
$\sum_{t=1}^{T/2}\frac{(t+t_{0})^{p}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}\leq\frac{1}{\min\{1,{\bar{\eta}}^{v}\}}\sum_{t=1}^{T/2}\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}.$
As $t+t_{0}+1\leq\frac{3}{4}(T+t_{0}+1)$ when $t\leq T/2$ and $T\geq t_{0}+1$, we obtain
$\sum_{t=1}^{T/2}\frac{(t+t_{0})^{p}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}$ $\leq\frac{\left(1-(3/4)^{1-\theta_{1}}\right)^{-v}}{\min\{1,{\bar{\eta}}^{v}\}}\sum_{t=1}^{T/2}\frac{(t+t_{0})^{p}}{(T+t_{0})^{(1-\theta_{1})v}}$
[…]
$\leq\frac{\left(1-(3/4)^{1-\theta_{1}}\right)^{-v}}{\min\{1,{\bar{\eta}}^{v}\}}(T+t_{0})^{-(1-\theta_{1})v}\begin{cases}\frac{1}{p+1}(T+t_{0})^{p+1},&\text{ when }p>-1,\\ \log(T+t_{0}),&\text{ when }p=-1,\\ \frac{t_{0}^{p+1}}{-1-p}\leq\frac{1}{-1-p},&\text{ when }p<-1,\end{cases}$
$\leq\delta^{\prime}\begin{cases}(T+t_{0})^{-(1-\theta_{1})v+p+1},&\text{ when }p>-1,\\ (T+t_{0})^{-(1-\theta_{1})v}\log(T+t_{0}),&\text{ when }p=-1,\\ (T+t_{0})^{-(1-\theta_{1})v},&\text{ when }p<-1,\end{cases}$
where
$\delta^{\prime}=\frac{\left(1-(3/4)^{1-\theta_{1}}\right)^{-v}}{\min\{1,{\bar{\eta}}^{v}\}}\begin{cases}\frac{1}{p+1},&\text{ when }p>-1,\\ 1,&\text{ when }p=-1,\\ \frac{1}{-1-p},&\text{ when }p<-1,\end{cases}$
which is independent of $T$ and $t_{0}$.
The proof is then finished. ∎
Lemma A.3.
Let $v>0$, $p\in\mathbb{R}$, $T\geq t_{0}+1$ and $t_{0}\geq 1$. The step size $\eta_{t}$ is set as (4.4). Then, there holds
$\sum_{t=T/2}^{T}\frac{(t+t_{0})^{p}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}\leq\delta^{\prime\prime}\begin{cases}(T+t_{0})^{p+\theta_{1}},&\text{ when}v>1,\\ (T+t_{0})^{p+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{p+1-v(1-\theta_{1})},&\text{ when }v<1,\end{cases}$
where $\delta^{\prime\prime}$ is a constant independent of $T$ and $t_{0}$.
Proof.
It is obvious that
$\begin{split}&\sum_{t=T/2}^{T}\frac{(t+t_{0})^{p}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}\\ &\quad\leq\frac{1}{\min\{1,{\bar{\eta}}^{v}\}}\sum_{t=T/2}^{T-1}\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}+(T+t_{0})^{p}.\end{split}$ (A.6)
Next, we bound $\sum_{t=T/2}^{T-1}\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}$. If $p>0$, it holds that
[…]
If $p\leq 0$, there holds $(t+t_{0})^{p}\leq 3^{-p}(t+t_{0}+2)^{p}$. Thus we have
[…]
Therefore,
$\sum_{t=T/2}^{T-1}\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}$
[…]
Let $\xi=(T+t_{0}+1)^{1-\theta_{1}}-(u+t_{0})^{1-\theta_{1}}$, then $d\xi=-(1-\theta_{1})(u+t_{0})^{-\theta_{1}}du$, then
$\sum_{t=T/2}^{T-1}$ $\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}$ (A.7)
[…]
Since $u+t_{0}\in\left[T/2+t_{0}+1,T+t_{0}+1\right]$, whenever $p+\theta_{1}>0$ or not, we have
$(u+t_{0})^{p+\theta_{1}}\leq 2^{\lvert p+\theta_{1}\rvert}(T+t_{0})^{p+\theta_{1}}.$ (A.8)
Hence, substituting (A.8) to (A.7) yields that
$\sum_{t=T/2}^{T-1}$ $\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}$ (A.9)
[…]
$\leq\frac{\max\{3^{-p},1\}}{1-\theta_{1}}2^{\lvert p+\theta_{1}\rvert}(T+t_{0})^{p+\theta_{1}}\begin{cases}\frac{v}{v-1},&\text{ when }v>1,\\ \left(2-\theta_{1}\right)\log(T+t_{0}+1),&\text{ when }v=1,\\ \frac{\left(1-2^{\theta_{1}-1}\right)^{1-v}}{1-v}(T+t_{0}+1)^{(1-\theta_{1})(1-v)},&\text{ when }v<1,\end{cases}$
$=\widetilde{\delta^{\prime\prime}}\begin{cases}(T+t_{0})^{p+\theta_{1}},&\text{ when }v>1,\\ (T+t_{0})^{p+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{p+1-v(1-\theta_{1})},&\text{ when }v<1,\end{cases}$
with some constant $\widetilde{\delta^{\prime\prime}}$ independent of $T$ and $t_{0}$. Combining (A.9) with (A.6) yields that there exists a constant $\delta^{\prime\prime}$ independent of $T$ and $t_{0}$, such that
$\sum_{t=T/2}^{T}\frac{(t+t_{0})^{p}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}\leq\delta^{\prime\prime}\begin{cases}(T+t_{0})^{p+\theta_{1}},&\text{ when}v>1,\\ (T+t_{0})^{p+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{p+1-v(1-\theta_{1})},&\text{ when }v<1,\end{cases}$ (A.10)
which completes the proof. ∎
Proposition A.4.
Let $v>0$, $\theta\in\mathbb{R}$, $t_{0}\geq 1$ and $T\geq t_{0}+1$. The step size $\eta_{t}$ is set as (4.4). Suppose that $\bar{\eta}\bar{\lambda}>\theta-1$ and $\theta_{1}+\theta_{2}=1$. Then, there holds
$\sum_{t=1}^{T}\exp\left\{-\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{(t+t_{0})^{-\theta}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}\leq\delta_{1}\begin{cases}(T+t_{0})^{-\theta+\theta_{1}},&\text{ when }v>1,\\ (T+t_{0})^{-\theta+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{-\theta+1-v(1-\theta_{1})},&\text{ when }v<1,\end{cases}$
where $\delta_{1}$ is a constant independent of $T$ and $t_{0}$.
Proof.
From Lemma 5.3 (2), we have
$\sum_{t=1}^{T}$ $\exp\left\{-\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{(t+t_{0})^{-\theta}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}\leq\sum_{t=1}^{T}\left(\frac{t+t_{0}+1}{T+t_{0}+1}\right)^{\bar{\eta}\bar{\lambda}}\frac{(t+t_{0})^{-\theta}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}$
$\leq 2^{\bar{\eta}\bar{\lambda}}(T+t_{0})^{-\bar{\eta}\bar{\lambda}}\sum_{t=1}^{T}\frac{(t+t_{0})^{\bar{\eta}\bar{\lambda}-\theta}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}.$
Using Lemma A.2 and Lemma A.3, we obtain
$\sum_{t=1}^{T}\exp\left\{-\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{(t+t_{0})^{-\theta}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}$
$\leq$ $2^{\bar{\eta}\bar{\lambda}}(T+t_{0})^{-\bar{\eta}\bar{\lambda}}\left(\delta^{\prime}(T+t_{0})^{-(1-\theta_{1})v+p+1}+\delta^{\prime\prime}\begin{cases}(T+t_{0})^{p+\theta_{1}},&\text{ when }v>1,\\ (T+t_{0})^{p+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{p+1-v(1-\theta_{1})},&\text{ when }v<1,\end{cases}\right)$
where $p=\bar{\eta}\bar{\lambda}-\theta>-1$. Therefore,
$\sum_{t=1}^{T}\exp\left\{-\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{(t+t_{0})^{-\theta}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}$ (A.11)
$\leq$ $2^{\bar{\eta}\bar{\lambda}}(\delta^{\prime}+\delta^{\prime\prime})\begin{cases}(T+t_{0})^{-\theta+\theta_{1}},&\text{ when }v>1,\\ (T+t_{0})^{-\theta+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{-\theta+1-v(1-\theta_{1})},&\text{ when }v<1.\end{cases}$
We finish the proof by setting $\delta_{1}=2^{\bar{\eta}\bar{\lambda}}(\delta^{\prime}+\delta^{\prime\prime})$, which is independent of $T$ and $t_{0}$. ∎

### v3

@@
 Abstract
-This paper investigates regularized stochastic gradient descent (SGD) algorithms for estimating nonlinear operators from a Polish space to a separable Hilbert space.
-We assume that the regression operator lies in a vector-valued reproducing kernel Hilbert space induced by an operator-valued kernel.
-Two significant settings are considered: an online setting with polynomially decaying step sizes and regularization parameters, and a finite-horizon setting with constant step sizes and regularization parameters.
-We introduce regularity conditions on the structure and smoothness of the target operator and the input random variables.
-Under these conditions, we provide a dimension-free convergence analysis for the prediction and estimation errors, deriving both expectation and high-probability error bounds.
-Our analysis demonstrates that these convergence rates are nearly optimal.
-Furthermore, we present a new technique for deriving bounds with high probability for general SGD schemes, which also ensures almost-sure convergence.
-Finally, we discuss potential extensions to more general operator-valued kernels and the encoder-decoder framework.
+We consider a class of statistical inverse problems involving the estimation of a regression operator from a Polish space to a separable Hilbert space, where the target lies in a vector-valued reproducing kernel Hilbert space induced by an operator-valued kernel.
+To address the associated ill-posedness, we analyze regularized stochastic gradient descent (SGD) algorithms in both online and finite-horizon settings.
+The former uses polynomially decaying step sizes and regularization parameters, while the latter adopts fixed values.
+Under suitable structural and distributional assumptions, we establish dimension-independent bounds for prediction and estimation errors.
+The resulting convergence rates are near-optimal in expectation, and we also derive high-probability estimates that imply almost sure convergence.
+Our analysis introduces a general technique for obtaining high-probability guarantees in infinite-dimensional settings.
+Possible extensions to broader kernel classes and encoder-decoder structures are briefly discussed.
 Keywords: Nonlinear operator learning, Operator-valued kernel, Regularized stochastic gradient descent, Convergence analysis
 1 Introduction
-In this paper, we consider a general model abstracted from nonlinear operator learning problems:
-$y=h^{\dagger}(x)+\epsilon.$ (1.1)
-Here, $\mathcal{X}$ is a Polish space and $\left(\mathcal{Y},\langle\cdot,\cdot\rangle_{\mathcal{Y}},\|\cdot\|_{\mathcal{Y}}\right)$ is a separable Hilbert space.
-The pair $(x,y)$ satisfying (1.1) is a random variable taking values in $\mathcal{X}\times\mathcal{Y}$, distributed according to an unknown probability measure $\rho$.
-The operator $h^{\dagger}:\mathcal{X}\rightarrow\mathcal{Y}$ is a measurable (possibly nonlinear) mapping defined by the conditional expectation $h^{\dagger}(x):=\mathbb{E}[y|x]$.
-The noise term $\epsilon$ is a centered $\mathcal{Y}$-valued random variable, assumed to be independent of $x$ and to have finite variance, i.e., $\sigma^{2}:=\mathbb{E}[\|\epsilon\|^{2}_{\mathcal{Y}}]<\infty$.
-The model (1.1) has been widely employed in surrogate approaches for structured output prediction [46, 21, 26, 7, 14, 4].
-In practice, many applications involve inputs or outputs with explicit or implicit discrete structures.
-Examples of implicitly structured data used in predictions include text, images, and videos in document processing and retrieval, as well as genes and proteins in computational biology.
-To learn models that predict outputs with structured components, surrogate methods embed the structured output into a Hilbert space, which formulates the task as operator regression with an infinite-dimensional output space.
-During prediction, a decoding step maps the output from the Hilbert space back to the original structured output space.
-Structured prediction tasks such as image completion [46], label ranking [29], and graph prediction [6] can thus be addressed through operator learning using surrogate approaches.
-Another important application of model (1.1) is functional output regression [24, 28, 25, 41].
-These problems have become increasingly relevant with the growing capacity to collect functional data, motivating a shift toward a functional perspective in modeling [25].
-This has led to the development of the now-thriving field of operator learning [1, 30], which aims to approximate operators between Hilbert (or more generally, Banach) spaces using data.
-A prominent example is the learning of solution operators for partial differential equations (PDEs), where the goal is to approximate the operator that maps a parameter space—describing the physical and geometrical constraints of the PDE—to its solution space [33, 3, 36].
-In this paper, we study operator learning algorithms designed to act directly on functions rather than on high-dimensional vectors.
-This functional perspective allows us to capture the intrinsic properties of the problem, avoiding reliance on specific discretizations or pixelizations.
+In this work, we consider the nonlinear regression model
+$y=h^{\dagger}(x)+\epsilon,$ (1.1)
+where $x$ belongs to a Polish space $\mathcal{X}$, and $y$ lies in a separable Hilbert space $\mathcal{Y}$.
+The pair $(x,y)$ is distributed according to an unknown probability measure $\rho$ on $\mathcal{X}\times\mathcal{Y}$.
+The regression operator $h^{\dagger}:\mathcal{X}\to\mathcal{Y}$ is defined as the conditional expectation
+$h^{\dagger}(x)=\mathbb{E}[y|x],$
+and the noise $\epsilon$ is a centered $\mathcal{Y}$-valued random variable with finite variance $\sigma^{2}:=\mathbb{E}[\|\epsilon\|^{2}_{\mathcal{Y}}]$, assumed independent of $x$.
+The objective is to recover $h^{\dagger}$ from independent, identically distributed samples of $(x,y)$.
+This formulation encompasses a broad class of statistical inverse problems in infinite-dimensional settings, where the forward map $h^{\dagger}$ must be recovered from noisy, indirect observations.
+Traditional approaches to inverse problems often employ explicit regularization techniques, such as Tikhonov regularization or spectral filtering, to stabilize the solution, typically under certain assumptions on the forward operator’s structure.
+In our work, we also use explicit regularization by penalizing the norm in a vector-valued reproducing kernel Hilbert space (RKHS).
+However, unlike classical deterministic frameworks, our approach studies stochastic gradient descent algorithms in infinite-dimensional function spaces with streaming data, requiring new analytical tools to establish convergence and error bounds under operator regularity conditions.
+A key motivation for this model arises from structured output prediction, where the outputs exhibit functional or combinatorial structure.
+In many practical scenarios, the output space is not finite-dimensional but consists of complex objects such as functions, sequences, or graphs.
+To address these tasks, surrogate learning approaches embed the structured output into a Hilbert space and learn a continuous operator mapping inputs to embedded outputs.
+Applications include image completion [43], label ranking [27], graph prediction [6], and protein function prediction [14].
+This framework also includes functional output regression, which has become increasingly relevant with the growing availability of functional data in science and engineering.
+Here, both the input and output may be infinite-dimensional, motivating the development of operator-theoretic regression approaches.
+These applications have spurred the growth of the operator learning paradigm, which aims to learn mappings between function spaces directly from data.
+A prominent example is learning solution operators for parameterized partial differential equations (PDEs), where the operator maps coefficient or boundary condition functions to PDE solutions [30, 3, 33].
+In this paper, we focus on learning algorithms that operate directly on function spaces rather than on finite-dimensional discretizations.
+This functional viewpoint preserves the intrinsic geometry and infinite-dimensional structure of the problem, enabling discretization-invariant analysis and broad applicability.
+Our setting naturally leads to the study of regularized learning in vector-valued RKHSs, where the regression operator is modeled within a rich nonparametric hypothesis space induced by operator-valued kernels.
 We introduce a supervised learning framework based on model (1.1).
@@
 In particular, our analysis covers most operator-valued kernels, including those of the form (1.2) with a compact operator $W$.
-Furthermore, when choosing kernels as in (1.2), it follows from [12, Example 5] that $K$ is a Mercer [resp.
+Furthermore, when choosing kernels as in (1.2), it follows from [11, Example 5] that $K$ is a Mercer [resp.
 $\mathcal{C}_{0}$] [^2] That is, the Banach space of continuous functions vanishing at infinity with the uniform norm.
@@
 The proof of Proposition 2.1 is deferred to Appendix A.1.
-By applying Proposition 2.1 with $W=I$—in which case the kernel coincides with that specified by Assumption 1—the iteration (1.5) can be equivalently expressed as
+By applying Proposition 2.1 with W=IW=I—in which case the kernel coincides with that specified by Assumption 1—the iteration (1.5) can be equivalently expressed as
 $\begin{cases}H_{1}=\mathbf{0},\\ H_{t+1}=H_{t}-\eta_{t}\left(\left(H_{t}\phi(x_{t})-y_{t}\right)\otimes\phi(x_{t})+\lambda_{t}H_{t}\right),\\ h_{t}(\cdot)=H_{t}\left(\phi(\cdot)\right).\end{cases}$ (2.1)
@@
 In this paper, Assumption 3 is introduced to derive sharper error bounds.
-When combined with Assumption 3 for some $10<s<$ 1, it leads to improved convergence rates.
+When combined with Assumption 3 for some $0<s<1$, it leads to improved convergence rates.
 This condition—commonly known as the capacity condition—was first introduced in [17] and has since been widely adopted in the literature, including [35, 4, 21, 38], as a way to capture the intrinsic complexity of infinite-dimensional learning problems.
@@
 $\mathbb{E}\left[\left\|A\phi(x)\right\|^{4}_{\mathcal{H}_{\mathcal{K}}}\right]\leq c\left(\mathbb{E}\left[\|A\phi(x)\|^{2}_{\mathcal{H}_{\mathcal{K}}}\right]\right)^{2}.$
-According to [41, Proposition 2.1], this assumption is equivalent to
+According to [38, Proposition 2.1], this assumption is equivalent to
 $\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{4}\right]\leq c\left(\mathbb{E}\left[\left\langle\phi(x),f\right\rangle_{\mathcal{H}_{\mathcal{K}}}^{2}\right]\right)^{2},\quad\forall f\in\mathcal{H}_{\mathcal{K}}.$ (2.3)
@@
 Suppose that Assumption 1, Assumption 2, Assumption 3 and Assumption 4 are satisfied.
-Define $\{h_{t}\}_{t\geq 1}$ through (2.1) with step sizes $\{\eta_{t}=\bar{\eta}(t+t_{0})^{-\theta_{1}}\}_{t\geq 1}$ and regularization parameters $\{\lambda_{t}=\bar{\lambda}(t+t_{0})^{-\theta_{2}}\}_{t\geq 1}$, where $10<\theta_{1}<$ 1, $10<\theta_{2}<$ 1 and $\bar{\eta}\bar{\lambda}>\theta_{2}\min\{r,1\}$.
+Define $\{h_{t}\}_{t\geq 1}$ through (2.1) with step sizes $\{\eta_{t}=\bar{\eta}(t+t_{0})^{-\theta_{1}}\}_{t\geq 1}$ and regularization parameters $\{\lambda_{t}=\bar{\lambda}(t+t_{0})^{-\theta_{2}}\}_{t\geq 1}$, where $0<\theta_{1}<1$, $0<\theta_{2}<1$ and $\bar{\eta}\bar{\lambda}>\theta_{2}\min\{r,1\}$.
 Additionally, let $t_{0}$ satisfy $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, $t_{0}\geq\exp\{\frac{1}{\theta_{1}}\}$, and
@@
 Then for any $T\geq 1$,
-$\mathbb{E}_{z^{T}}\left[\mathcal{E}(h_{T+1})-\mathcal{E}(h^{\dagger})\right]\leq c_{1,1}\begin{cases}(T+t_{0})^{-\theta_{1}},&\text{when }s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0}),&\text{when }s=1.\\ \end{cases}$
+[…]
 Here the constant $c_{1,1}$ is independent of $T$, and will be given in the proof.
@@
 In Theorem 2.3, since the constant $c_{4}$ is independent of $t_{0}$, one can choose $t_{0}$ sufficiently large to satisfy the required conditions.
-Compared to Assumption 3 with $s=1$, the stronger assumption with $10<s<$ 1 only removes a logarithmic factor in the convergence rate.
+Compared to Assumption 3 with $s=1$, the stronger assumption with $0<s<1$ only removes a logarithmic factor in the convergence rate.
 It is also clear that the convergence rate saturates at $r=1/2$, i.e., increasing $r$ beyond $1/2$ does not yield further improvement.
-According to [41, Theorem 2.9], the result is minimax optimal (up to a logarithmic term) when $s=1$ and $r<1/2$.
-Compared to the unregularized SGD algorithm analyzed in [41, Theorem 2.4], adding a regularization term here leads to faster convergence.
-Specifically, while the prediction error rate of unregularized SGD in [41] saturates at $r=(1-s)/2$, the regularized SGD in our work improves the saturation level to $r=1/2$.
+According to [38, Theorem 2.9], the result is minimax optimal (up to a logarithmic term) when $s=1$ and $r<1/2$.
+Compared to the unregularized SGD algorithm analyzed in [38, Theorem 2.4], adding a regularization term here leads to faster convergence.
+Specifically, while the prediction error rate of unregularized SGD in [38] saturates at r=(1s)/2r=(1-s)/2, the regularized SGD in our work improves the saturation level to $r=1/2$.
 The following theorem provides the convergence rate for the estimation error.
@@
 Here the constant $c_{1,2}$ is independent of $T$, and will be given in the proof.
-The convergence rate of the estimation error saturates at $r=1$, which improves the convergence of unregularized SGD in [41, Theorem 2.6], where the rate saturates at $r=\frac{1-s}{2}$.
-Moreover, in [22, Theorem 3] and [41, Theorem 2.6], we cannot guarantee the convergence of the estimation error with decaying step sizes for $s=1$, whereas adding a regularization term addresses this issue.
-According to [41, Theorem 2.9], the convergence rate with decaying step sizes is minimax optimal when $r<1$.
+The convergence rate of the estimation error saturates at $r=1$, which improves the convergence of unregularized SGD in [38, Theorem 2.6], where the rate saturates at $r=\frac{1-s}{2}$.
+Moreover, in [21, Theorem 3] and [38, Theorem 2.6], we cannot guarantee the convergence of the estimation error with decaying step sizes for $s=1$, whereas adding a regularization term addresses this issue.
+According to [38, Theorem 2.9], the convergence rate with decaying step sizes is minimax optimal when $r<1$.
 Next, we present the convergence rates for prediction and estimation errors with constant step sizes and regularization parameters, where both depend on the total number of iterations $T$ (i.e., the total sample size).
@@
 Then, with probability at least $1-2\delta$,
-$\mathcal{E}(h_{T+1})-\mathcal{E}(h^{\dagger})$ $\leq c_{2,3}\begin{cases}T^{-\theta_{3}}\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{2}T\log^{4}\frac{2}{\delta},&\text{ when }s<1,\\ T^{-\theta_{3}}\log T\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{3}T\log^{4}\frac{2}{\delta},&\text{ when }s=1,\\ \end{cases}$
-$\lesssim\log^{4}\frac{2}{\delta}\begin{cases}T^{-\theta_{3}},&\text{ when }s<1,\\ T^{-\theta_{3}}\log T,&\text{ when }s=1.\end{cases}$
+[…]
+[…]
 Here the constant $c_{2,3}$ is independent of $T$ and $\delta$, and will be given in the proof.
@@
   1. (1)
-    Let $\mathcal{H}$ be separable, which is true if the spaces $\mathcal{X}$ and $\mathcal{Y}$ are separable and $K$ is a Mercer kernel [11, Corollary 5.2].
-A kernel $K$ is Mercer if and only if the RKHS induced by $K$ is a subspace of the spaces of continuous operators from $\mathcal{X}$ to $\mathcal{Y}$, which in turn holds if and only if $K$ is locally bounded and $K(x,\cdot)$ is strongly continuous for any $x\in\mathcal{X}$ [11, Proposition 5.1].
+    Let $\mathcal{H}$ be separable, which is true if the spaces $\mathcal{X}$ and $\mathcal{Y}$ are separable and $K$ is a Mercer kernel [12, Corollary 5.2].
+A kernel $K$ is Mercer if and only if the RKHS induced by $K$ is a subspace of the spaces of continuous operators from $\mathcal{X}$ to $\mathcal{Y}$, which in turn holds if and only if $K$ is locally bounded and $K(x,\cdot)$ is strongly continuous for any $x\in\mathcal{X}$ [12, Proposition 5.1].
   2. (2)
@@
   3. (3)
-    We further assume that $K$ is strongly measurable. Under assumptions (1) and (2), this is equivalent to requiring that each element in $\mathcal{H}$ is a measurable function [11, Proposition 3.3].
+    We further assume that $K$ is strongly measurable. Under assumptions (1) and (2), this is equivalent to requiring that each element in $\mathcal{H}$ is a measurable function [12, Proposition 3.3].
   4. (4)
@@
 $D(h)(\cdot):=\mathop{\arg\min}_{z\in\mathcal{Z}}\left\{\|h(\cdot)-\phi(z)\|_{\mathcal{Y}}\right\}$
-for any estimator $h$, as detailed in [13, 4].
-Let $\hat{f}=D(\hat{h})$, denote the estimator for $f^{\dagger}$ obtained via the algorithm.
-The surrogate approach for structured prediction is illustrated in Figure 1.
+[…]
 According to Ciliberto et al. [13], the following properties hold:
@@
 This makes it an effective tool for structured prediction in streaming or time-dependent environments.
-$\mathcal{X}$$\mathcal{Z}$$\mathcal{Y}$$\hat{h}$$\phi$$\hat{f}=D(\hat{h})$$D$
+[…]
 Figure 1: Surrogate approach for structured prediction
@@
 The PCA algorithm applied to a random input $x$ in $\mathcal{X}$ seeks to minimize the reconstruction error $\mathbb{E}\left[\left\|(I-P)x\right\|_{\mathcal{X}}^{2}\right]$ over $\Pi_{d_{\mathcal{X}}}$, the set of all orthogonal projections $P$ with rank $d_{\mathcal{X}}$.
-Given the covariance operator of $x$ defined as $\Sigma_{x}:=\mathbb{E}[x\otimes x]$, there exist eigenvalue-eigenvector pairs $\{\lambda_{i}^{d_{\mathcal{X}}},\phi_{i}^{d_{\mathcal{X}}}\}_{i\geq 1}$ satisfying $\langle\Sigma_{x}\phi_{i}^{d_{\mathcal{X}}},\phi_{j}^{d_{\mathcal{X}}}\rangle_{\mathcal{X}}=\delta_{ij}$ and $\lambda_{1}^{d_{\mathcal{X}}}\geq\lambda_{2}^{d_{\mathcal{X}}}\geq\cdots\geq 0$, where $\delta_{ij}=1$ if $i=j$, otherwise 00.
+Given the covariance operator of $x$ defined as $\Sigma_{x}:=\mathbb{E}[x\otimes x]$, there exist eigenvalue-eigenvector pairs $\{\lambda_{i}^{d_{\mathcal{X}}},\phi_{i}^{d_{\mathcal{X}}}\}_{i\geq 1}$ satisfying $\langle\Sigma_{x}\phi_{i}^{d_{\mathcal{X}}},\phi_{j}^{d_{\mathcal{X}}}\rangle_{\mathcal{X}}=\delta_{ij}$ and $\lambda_{1}^{d_{\mathcal{X}}}\geq\lambda_{2}^{d_{\mathcal{X}}}\geq\cdots\geq 0$, where $\delta_{ij}=1$ if $i=j$, otherwise 0.
 It can be shown that the optimal PCA projection is given by
@@
 $\mathcal{D}_{d_{\mathcal{X}}}^{\mathcal{X}}(\eta):=\sum_{i=1}^{d_{\mathcal{X}}}\eta_{i}\phi_{i}^{\mathcal{X}}=\left(\mathcal{E}_{d_{\mathcal{X}}}^{\mathcal{X}}\right)^{*}(\eta).$
-It then follows that $\mathbb{E}\left[\left\|(I-P_{d_{\mathcal{X}}}^{\mathcal{X}})x\right\|_{\mathcal{X}}^{2}\right]=\sum_{i>d_{\mathcal{X}}}\lambda_{i}^{d_{\mathcal{X}}}$, see [32, Theorem 3.8].
+It then follows that $\mathbb{E}\left[\left\|(I-P_{d_{\mathcal{X}}}^{\mathcal{X}})x\right\|_{\mathcal{X}}^{2}\right]=\sum_{i>d_{\mathcal{X}}}\lambda_{i}^{d_{\mathcal{X}}}$, see [28, Theorem 3.8].
 In practice, it is usually difficult to obtain $\Sigma_{x}$ directly, so we typically use the empirical covariance operator $\Sigma_{x}^{T}=\frac{1}{T}\sum_{i=1}^{T}x_{i}\otimes x_{i}$ as a substitute, thus deriving the empirical PCA.
-Following the same procedure, we naturally obtain the empirical encoder ^$\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$ and empirical decoder $\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$.
-Similarly, by replacing $\Sigma_{x}^{T}$, the input random variable $x$, and the rank $d_{\mathcal{X}}$ with $\Sigma_{y}^{T}:=\frac{1}{T}\sum_{i=1}^{T}y_{i}\otimes y_{i}$, the output random variable $y$ and $d_{\mathcal{Y}}$, respectively, we apply empirical PCA to $y$ in $\mathcal{Y}$ with rank $d_{\mathcal{Y}}$, and obtain the empirical encoder ^$\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$ and empirical decoder $\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$.
-$\mathcal{X}$$\mathcal{Y}$$\mathbb{R}^{d_{\mathcal{X}}}$$\mathbb{R}^{d_{\mathcal{Y}}}$${\color[rgb]{0.82,0.01,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.01,0.11}h^{\dagger}}$^$\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$$\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$^$\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$$\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$$f^{\dagger}$
-Figure 2: Commutative diagram of PCA enceoder-decoder framework
+Following the same procedure, we naturally obtain the empirical encoder $\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$ and empirical decoder $\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$.
+Similarly, by replacing $\Sigma_{x}^{T}$, the input random variable $x$, and the rank $d_{\mathcal{X}}$ with $\Sigma_{y}^{T}:=\frac{1}{T}\sum_{i=1}^{T}y_{i}\otimes y_{i}$, the output random variable $y$ and $d_{\mathcal{Y}}$, respectively, we apply empirical PCA to $y$ in $\mathcal{Y}$ with rank $d_{\mathcal{Y}}$, and obtain the empirical encoder $\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$ and empirical decoder $\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$.
+$\mathcal{X}$$\mathcal{Y}$$\mathbb{R}^{d_{\mathcal{X}}}$$\mathbb{R}^{d_{\mathcal{Y}}}$$\color[rgb]{0.82,0.01,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.01,0.11}h^{\dagger}$$\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$$\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$$\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$$\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$$f^{\dagger}$
+Figure 2: Commutative diagram of PCA encoder-decoder framework
 We now formulate the estimator as $h:=\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ f\circ\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$.
 It is then natural to choose $f^{\dagger}=\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ h^{\dagger}\circ\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$.
-This formulation naturally give rise to a commutative diagram, as illustrated in Figure 2.
+This formulation naturally gives rise to a commutative diagram, as illustrated in Figure 2.
 The works [3, 29] represent $f$ using neural networks.
-In constract, we represent $f$ in an RKHS induced by a matrix-valued kernel $k:\mathbb{R}^{d_{\mathcal{X}}}\times\mathbb{R}^{d_{\mathcal{X}}}\to\mathcal{B}(\mathbb{R}^{d_{\mathcal{Y}}})$.
+In contrast, we represent $f$ in an RKHS induced by a matrix-valued kernel $k:\mathbb{R}^{d_{\mathcal{X}}}\times\mathbb{R}^{d_{\mathcal{X}}}\to\mathcal{B}(\mathbb{R}^{d_{\mathcal{Y}}})$.
 Specifically, we consider kernels of the form $k(u,v)=\phi(\|u-v\|_{\mathbb{R}^{d_{\mathcal{X}}}})$, where $\phi:[0,\infty)\to\mathbb{R}$ is a radial function, such that $k$ is positive definite for $d_{\mathcal{X}}>0$.
@@
 where $\eta$ is the step size.
-This can be interpreted as an SGD scheme based on the samples {^$\{\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}x_{t},\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}y_{t}\}_{t=1}^{T}$ in $\mathcal{H}_{k}$.
+This can be interpreted as an SGD scheme based on the samples $\{\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}x_{t},\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}y_{t}\}_{t=1}^{T}$ in $\mathcal{H}_{k}$.
 Accordingly, we define $h_{1}=0$ and
@@
 As introduced in Section 1, we consider two types of step sizes and regularization parameters. Both can be uniformly expressed in the following form:
-$\begin{cases}\eta_{t}=\bar{\eta}(t+t_{0})^{-\theta_{1}},&\\ \lambda_{t}=\bar{\lambda}(t+t_{0})^{-\theta_{2}},&\end{cases}$ (4.4)
+[…]
 where $t_{0}\geq 0$, $\eta_{t}$ is the step size with $\theta_{1}\in[0,1)$ and $\lambda_{t}$ denotes the regularization parameter with $\theta_{2}\in[0,1)$.
@@
 $\leq\sum_{t=1}^{T}\eta_{t}^{2}\left(\mathbb{E}_{z^{t-1}}\sqrt{\mathbb{E}_{x_{t}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{4}}+\sigma^{2}\right)$
-×$\qquad\times\left(\mathbb{E}_{x_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{4}\right)^{1/2}$
+$\qquad\times\left(\mathbb{E}_{x_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{4}\right)^{1/2}$
 $\leq\sqrt{c}\sum_{t=1}^{T}\eta_{t}^{2}\left(\sqrt{c}\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}+\sigma^{2}\right)\mathbb{E}_{x_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2},$
@@
 $\left\|C^{r+\alpha}(C+\lambda_{T}I)^{-1}\right\|\leq\sup_{0\leq x\leq\kappa^{2}}\frac{x^{r+\alpha}}{x+\lambda_{T}}$ (5.2)
-$\leq\begin{cases}\kappa^{2(r+\alpha-1)},&\text{ when }r+\alpha\geq 1,\\ (r+\alpha)^{r+\alpha}(1-r-\alpha)^{1-r-\alpha}\lambda_{T}^{r+\alpha-1},&\text{when }r+\alpha<1,\end{cases}$
+[…]
 combining (5.1) with (5.2), there exists a constant $c_{1}$ such that
@@
 Lemma 5.3.
-Let $10\leq\theta_{1}<$ 1, $10\leq\theta_{2}<$ 1, and $\eta_{t},\lambda_{t}$ be defined as (4.4). Then the following bounds hold for $1\leq l\leq m$ with $l\in\mathbb{N}$:
+Let $0\leq\theta_{1}<1$, $0\leq\theta_{2}<1$, and $\eta_{t},\lambda_{t}$ be defined as (4.4). Then the following bounds hold for $1\leq l\leq m$ with $l\in\mathbb{N}$:
   • (1)
@@
 We bound the summation $\sum_{t=l}^{m}\eta_{t}$ using
-[…]
+$\sum_{t=l}^{m}\eta_{t}$ $=\bar{\eta}\sum_{t=l}^{m}(t+t_{0})^{-\theta_{1}}\geq\bar{\eta}\int_{l}^{m+1}(x+t_{0})^{-\theta_{1}}dx$
 $=\frac{\bar{\eta}}{1-\theta_{1}}\left[(m+t_{0}+1)^{1-\theta_{1}}-(l+t_{0})^{1-\theta_{1}}\right].$
@@
 Next, we analyze the summation involving $\lambda_{t}$ using the same estimate as before:
+$\sum_{t=l}^{m}\eta_{t}\lambda_{t}$ $=\bar{\eta}\bar{\lambda}\sum_{t=l}^{m}(t+t_{0})^{-\theta_{1}-\theta_{2}}\geq\bar{\eta}\bar{\lambda}\int_{l}^{m+1}(x+t_{0})^{-\theta_{1}-\theta_{2}}dx$
 […]
-=$\begin{cases}\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(m+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(l+t_{0})^{1-\theta_{1}-\theta_{2}}\right],&\text{ when }\theta_{1}+\theta_{2}\not=1,\\ \bar{\eta}\bar{\lambda}\log\left(\frac{m+t_{0}+1}{t_{0}+l}\right),&\text{ when}\theta_{1}+\theta_{2}=1.\end{cases}$
 For the case $l=1$, $m=T$, and $T\geq t_{0}+1$, we obtain:
@@
   • (1)
-    $10<\theta_{1}<$ 1, $10<\theta_{2}<$ 1, and $t_{0}>0$, corresponding to the online setting;
+    $0<\theta_{1}<1$, $0<\theta_{2}<1$, and $t_{0}>0$, corresponding to the online setting;
   • (2)
@@
 Suppose that Assumption 2 holds with $S^{\dagger}\in\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\mathcal{Y})$ and $r>0$.
-Then, for any $T\geq t_{0}+1$, $t_{0}\geq 0$, $10\leq\theta_{1}<$ 1, and $10\leq\theta_{2}<$ 1, the quantity
+Then, for any $T\geq t_{0}+1$, $t_{0}\geq 0$, $0\leq\theta_{1}<1$, and $0\leq\theta_{2}<1$, the quantity
 $\mathcal{T}_{2}=6\left\|H_{\lambda_{0}}C^{\alpha}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|_{\mathrm{HS}}^{2}$
@@
 $\mathcal{T}_{2}$ $=6\left\|H_{\lambda_{0}}C^{\alpha}\prod_{t=1}^{T}\left(I-\eta_{t}(C+\lambda_{t}I)\right)\right\|_{\mathrm{HS}}^{2}$ (5.7)
-$\leq c_{2}\bar{\eta}^{-2(r+\alpha)}\begin{cases}(T+t_{0})^{-2(r+\alpha)(1-\theta_{1})}\exp\{-\tau\bar{\eta}\bar{\lambda}(T+t_{0})^{1-\theta_{1}-\theta_{2}}\},&\text{ when }0\leq\theta_{1}+\theta_{2}<1,\\ (t_{0}+1)^{2\bar{\eta}\bar{\lambda}}(T+t_{0})^{-2(r+\alpha)(1-\theta_{1})-2\bar{\eta}\bar{\lambda}},&\text{ when }\theta_{1}+\theta_{2}=1,\\ (T+t_{0})^{-2(r+\alpha)(1-\theta_{1})},&\text{ when }\theta_{1}+\theta_{2}>1,\end{cases}$
+[…]
 where $c_{2}=6\|S^{\dagger}\|_{\mathrm{HS}}^{2}\left(\frac{(r+\alpha)(1-\theta_{1})}{e(1-2^{\theta_{1}-1})}\right)^{2(r+\alpha)}$ is independent of $t_{0}$, $T$, $\bar{\eta}$, and $\bar{\lambda}$.
@@
 Therefore, it is sufficient to bound $\mathcal{T}_{3}$ under the regime of decaying step sizes and regularization parameters.
-In what follows, we focus on the setting where $10<\theta_{1}<$ 1 and $10<\theta_{2}<$ 1.
+In what follows, we focus on the setting where $0<\theta_{1}<1$ and $0<\theta_{2}<1$.
 Lemma 5.5.
@@
 $\left\|H_{\lambda_{t-1}}-H_{\lambda_{t}}\right\|_{\mathrm{HS}}\leq\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1},$
-where $\widetilde{c}_{3}$ is a constant independent of $t_{0}$, $t$, $\bar{\eta}$, and $\bar{\lambda}$.
+where $\widetilde{c}_{3}$ is a constant independent of $_{0}, tt$, $\bar{\eta}$, and $\bar{\lambda}$.
 Proof.
@@
 $\leq\|S^{\dagger}\|_{\mathrm{HS}}\left\lvert\lambda_{t}-\lambda_{t-1}\right\rvert\left\|C^{r}(C+\lambda_{t}I)^{-1}\right\|$
-$\leq\|S^{\dagger}\|_{\mathrm{HS}}\bar{\lambda}\left\lvert(t+t_{0}-1)^{-\theta_{2}}-(t+t_{0})^{-\theta_{2}}\right\rvert\begin{cases}\kappa^{2r-2},&\text{ when }r\geq 1,\\ r^{r}(1-r)^{1-r}\lambda_{t}^{r-1},&\text{ when }r<1,\end{cases}$
+[…]
 where the last inequality uses the fact that
@@
 $\leq\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1},$
-where $\widetilde{c}_{3}$ is a constant independent of $t_{0}$, $t$, $\bar{\eta}$, and $\bar{\lambda}$.
+where $\widetilde{c}_{3}$ is a constant independent of $_{0}, tt$, $\bar{\eta}$, and $\bar{\lambda}$.
 The proof is complete. ∎
@@
 Then, the following bound holds for $\mathcal{T}_{3}$:
-$\mathcal{T}_{3}=6\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|^{2}_{\mathrm{HS}}\leq c_{3}\begin{cases}1,&\text{ when }\theta_{1}+\theta_{2}>1,\\ (T+t_{0})^{-2\theta_{2}\min\{r,1\}},&\text{ when }\theta_{1}+\theta_{2}\leq 1,\end{cases}$
+[…]
 where $c_{3}=c_{3}(t_{0},\bar{\lambda},\bar{\eta})$ is a constant independent of $T$.
@@
 Thus, we obtain the following bound for the exponential term:
-$\exp\left\{-\sum_{j=t}^{T}\eta_{j}\lambda_{j}\right\}\leq\begin{cases}\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(T+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(t+t_{0})^{1-\theta_{1}-\theta_{2}}\right]\right\},&\text{ when }\theta_{1}+\theta_{2}<1,\\ (t+t_{0})^{\bar{\eta}\bar{\lambda}}(T+t_{0})^{-\bar{\eta}\bar{\lambda}},&\text{ when }\theta_{1}+\theta_{2}=1,\\ 1,&\text{ when }\theta_{1}+\theta_{2}>1.\end{cases}$
+[…]
 We next consider the three cases corresponding to $\theta_{1}+\theta_{2}>1$, $\theta_{1}+\theta_{2}=1$, and $\theta_{1}+\theta_{2}<1$.
@@
 $\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|_{\mathrm{HS}}\leq\widetilde{c}_{3}\bar{\lambda}^{\min\{r,1\}}\sum_{t=1}^{T}(t+t_{0})^{-\theta_{2}\min\{r,1\}-1}$
-×$\times\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(T+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(t+t_{0})^{1-\theta_{1}-\theta_{2}}\right]\right\}.$
+$\times\exp\left\{-\frac{\bar{\eta}\bar{\lambda}}{1-\theta_{1}-\theta_{2}}\left[(T+t_{0}+1)^{1-\theta_{1}-\theta_{2}}-(t+t_{0})^{1-\theta_{1}-\theta_{2}}\right]\right\}.$
 Now, we estimate the summation in the last inequality.
@@
 $\mathcal{T}_{3}$ $=6\left\|\sum_{t=1}^{T}(H_{\lambda_{t-1}}-H_{\lambda_{t}})\prod_{j=t}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|^{2}_{\mathrm{HS}}$
-$\leq c_{3}\begin{cases}1,&\text{ when }\theta_{1}+\theta_{2}>1,\\ (T+t_{0})^{-2\theta_{2}\min\{r,1\}},&\text{ when }\theta_{1}+\theta_{2}\leq 1,\end{cases}$
+[…]
 where $c_{3}$ is a constant independent of $T$.
@@
 Note that we cannot guarantee convergence of the estimation error when $\theta_{1}+\theta_{2}>1$.
-However, convergence of the prediction error is essured when $\theta_{1}+\theta_{2}>1$ with $10<\theta_{1},\theta_{2}<$ 1, and both the prediction and estimation errors converge when $\theta_{1}+\theta_{2}\leq 1$.
+However, convergence of the prediction error is essured when $\theta_{1}+\theta_{2}>1$ with $0<\theta_{1},\theta_{2}<1$, and both the prediction and estimation errors converge when $\theta_{1}+\theta_{2}\leq 1$.
 The proofs for the remaining cases are similar and are omitted here to avoid repetition.
@@
 Suppose that Assumption 3 holds with $0<s\leq 1$.
-If $t_{0}\geq 1$, $T\geq t_{0}+1$, $10<\theta_{1}<$ 1, $10<\theta_{2}<$ 1, $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, and the following condition holds for any $t\in\mathbb{N}_{T}$:
+If $t_{0}\geq 1$, $T\geq t_{0}+1$, $0<\theta_{1}<1$, $0<\theta_{2}<1$, $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, and the following condition holds for any $t\in\mathbb{N}_{T}$:
 $\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}\left(=\mathbb{E}_{z^{t-1}}\left[\left\|\left(H_{t}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\right]\right)\leq M,$ (5.13)
@@
 $\leq$ $2\left(\kappa^{2+4\alpha-2s}+((1+2\alpha-s)/(2e))^{1+2\alpha-s}\right)\left(\sqrt{c}M+\sigma^{2}\right)\mathrm{Tr}\left(C^{s}\right)$
-×$\times\sum_{t=1}^{T}\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{\eta_{t}^{2}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{1+2\alpha-s}}.$
+$\times\sum_{t=1}^{T}\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{\eta_{t}^{2}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{1+2\alpha-s}}.$
 Now, applying Proposition 5.7 with $\theta=2\theta_{1}$ and $v=1+2\alpha-s$, we get
 $\mathcal{T}_{4}$ $=6\sqrt{c}\sum_{t=1}^{T}\eta_{t}^{2}\left(\sqrt{c}\mathbb{E}_{Z^{t}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}+\sigma^{2}\right)\mathrm{Tr}\left(C^{1+2\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right)$
-$\leq c_{4}\left(\sqrt{c}M+\sigma^{2}\right)\begin{cases}(T+t_{0})^{-\theta_{1}},&\text{ when }2\alpha>s,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0}),&\text{ when }2\alpha=s,\\ (T+t_{0})^{-(1+s-2\alpha)\theta_{1}+s-2\alpha},&\text{ when }2\alpha<s,\end{cases}$
+[…]
 where $c_{4}=12\sqrt{c}\left(\kappa^{2+4\alpha-2s}+((1+2\alpha-s)/(2e))^{1+2\alpha-s}\right)\mathrm{Tr}\left(C^{s}\right)\delta_{1}\bar{\eta}^{2}$.
 The proof is thus complete. ∎
-The next proposition is used to bound $\mathcal{T}_{4}$ in the finite-horizon setting. We define $00^{0}:=$ 0 for convenience.
+The next proposition is used to bound $\mathcal{T}_{4}$ in the finite-horizon setting. We define $0^{0}:=0$ for convenience.
 Proposition 5.10.
-Let $v\geq 0$, $\bar{\eta}=\eta_{1}T^{-\theta_{3}}$, $10<\theta_{3}<$ 1, $\theta_{4}>0$, and $\eta_{1}$, $\lambda_{1}$ be constants independent of $T$.
+Let $v\geq 0$, $\bar{\eta}=\eta_{1}T^{-\theta_{3}}$, $0<\theta_{3}<1$, $\theta_{4}>0$, and $\eta_{1}$, $\lambda_{1}$ be constants independent of $T$.
 Then, there exists a constant $\delta_{3}$, independent of $T$, such that for any $T\geq 2$,
-$\sum_{t=0}^{T-1}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}}{1+\left(t\bar{\eta}\right)^{v}}\leq\delta_{3}\begin{cases}T^{v\theta_{3}+(1-v)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }0\leq v<1,\\ T^{\theta_{3}}\log T,&\text{ when }v=1,\\ T^{\theta_{3}},&\text{ when }v>1.\end{cases}$
+[…]
 Proof.
-We divide the proof into three cases: $v=0$, $v>0$, and $10<v<$ 1 with $\theta_{3}+\theta_{4}<1$. The third case is an improvement upon the analysis in the second case.
+We divide the proof into three cases: $v=0$, $v>0$, and $0<v<1$ with $\theta_{3}+\theta_{4}<1$. The third case is an improvement upon the analysis in the second case.
 Case 1: $v=0$
@@
 We bound the summation as
+$\sum_{t=0}^{T-1}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}})\right\}}{1+\left(t\bar{\eta}\right)^{v}}\leq$ $1+\int_{0}^{T-1}\frac{1}{1+\left(t\bar{\eta}\right)^{v}}dt\leq 1+\frac{1}{\bar{\eta}}\int_{0}^{\bar{\eta}(T-1)}\frac{1}{1+t^{v}}dt$ (5.14)
+$\leq$ $1+\frac{1}{\bar{\eta}}\left(1+\int_{1}^{\bar{\eta}(T-1)}t^{-v}dt\right)$
 […]
 […]
-$\leq$ $1+\frac{1}{\bar{\eta}}+\frac{1}{\bar{\eta}}\begin{cases}\frac{\left(\bar{\eta}T\right)^{1-v}}{1-v},&\text{ when }0<v<1,\\ \log(\bar{\eta}T),&\text{ when }v=1,\\ \frac{1}{v-1},&\text{ when }v>1,\end{cases}$
-$\leq$ $\delta_{3}\begin{cases}T^{1-v+\theta_{3}v},&\text{ when }0<v<1,\\ T^{\theta_{3}}\log T,&\text{ when }v=1,\\ T^{\theta_{3}},&\text{ when }v>1,\end{cases}$
 where $\delta_{3}$ is a constant independent of $T$.
-Case 3: $10<v<$ 1 with $\theta_{3}+\theta_{4}<1$
+Case 3: $0<v<1$ with $\theta_{3}+\theta_{4}<1$
 In this case, a more refined estimation can be achieved compared to Case 2. We split the summation into three parts,
@@
 $\leq$ $1+\sum_{t=1}^{T^{\theta_{3}+\theta_{4}}}\frac{1}{1+\left(t\bar{\eta}\right)^{v}}+\sum_{t=T^{\theta_{3}+\theta_{4}}}^{T}\frac{\exp\left\{-2\lambda_{1}\eta_{1}tT^{-\theta_{4}-\theta_{3}}\right\}}{1+\left(t\bar{\eta}\right)^{v}}$
-=:absent:=:= : $1+\mathcal{A}_{1}+\mathcal{A}_{2}.$
+$=:$ $1+\mathcal{A}_{1}+\mathcal{A}_{2}.$
 We estimate $\mathcal{A}_{1}$ in the same manner as in (5.14). Noting that $\bar{\eta}T^{\theta_{3}+\theta_{4}}=\eta_{1}T^{\theta_{4}}$, we obtain
-[…]
+$\mathcal{A}_{1}$ $\leq\int_{0}^{T^{\theta_{3}+\theta_{4}}}\frac{1}{1+\left(t\bar{\eta}\right)^{v}}dt\leq\frac{1}{\bar{\eta}}+\frac{1}{\bar{\eta}}\int_{1}^{\bar{\eta}T^{\theta_{3}+\theta_{4}}}\frac{1}{t^{v}}dt$
 $\leq\frac{1}{\eta_{1}}T^{\theta_{3}}\left(1+\frac{\left(\eta_{1}T^{\theta_{4}}\right)^{1-v}}{1-v}\right)\leq\frac{1}{\eta_{1}}\left(1+\frac{\eta_{1}^{1-v}}{1-v}\right)T^{\theta_{4}(1-v)+\theta_{3}}.$
@@
 Letting $x=tT^{-\theta_{4}-\theta_{3}}$, we rewrite the above as
-[…]
+$\mathcal{A}_{2}\leq\eta_{1}^{-v}T^{(1-v)\theta_{4}+\theta_{3}}\int_{k}^{+\infty}x^{-v}\exp\{-2\lambda_{1}\eta_{1}x\}dx.$
 Since the integral is finite and satisfies
-[…]
+$\int_{k}^{+\infty}x^{-v}\exp\{-2\lambda_{1}\eta_{1}x\}dx\leq k^{-v}\frac{\exp\{-2\lambda_{1}\eta_{1}k\}}{2\lambda_{1}\eta_{1}}<\infty,$
 we combine the bounds for $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, and use (5.15), to conclude that
@@
 Suppose that Assumption 3 holds with $0<s\leq 1$.
-Let $\alpha\in[0,\frac{1}{2}]$, and set the parameters $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ with $10<\theta_{3}<$ 1 and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ with $\theta_{4}>0$.
+Let $\alpha\in[0,\frac{1}{2}]$, and set the parameters $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ with $0<\theta_{3}<1$ and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ with $\theta_{4}>0$.
 Additionally, assume that $T\geq 2$, $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$ and there exists a constant $\widetilde{M}$ independent of $T$, such that for all $t\in\mathbb{N}_{T}$,
-$\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}\left(=\mathbb{E}_{z^{t-1}}\left[\left\|\left(H_{t}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\right]\right)\leq\widetilde{M}.$ (5.16)
+$\mathbb{E}_{z^{t-1}}\left\|(H_{t}-H^{\dagger})\phi(x_{t})\right\|_{\mathcal{Y}}^{2}\left(=\mathbb{E}_{z^{t-1}}\left[\left\|\left(H_{t}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\right]\right)\leq\widetilde{M}$.
+(5.16)
 Recall that
@@
 $\leq$ $6\sqrt{c}\left(\sqrt{c}\widetilde{M}+\sigma^{2}\right)\mathrm{Tr}(C^{s})\bar{\eta}^{2}\sum_{t=1}^{T}\left\|C^{1+2\alpha-s}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right\|.$
-If $01+2\alpha-s>$ 0, applying Lemma 5.2 (3) for $11\leq t\leq T-$ 1, we obtain the following estimate, which also holds for $t=T$,
+If $1+2\alpha-s>0$, applying Lemma 5.2 (3) for $1\leq t\leq T-1$, we obtain the following estimate, which also holds for $t=T$,
 $\left\|C^{1+2\alpha-s}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right\|\leq\exp\left\{-2(T-t)\bar{\lambda}\bar{\eta}\right\}\frac{2(\kappa^{2(1+2\alpha-s)}+(\frac{1+2\alpha-s}{2e})^{1+2\alpha-s})}{1+\left((T-t)\bar{\eta}\right)^{1+2\alpha-s}}.$
-If $01+2\alpha-s=$ 0, we derive the bound
+If $1+2\alpha-s=0$, we derive the bound
 $\left\|C^{1+2\alpha-s}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right\|\leq\kappa^{2(1+2\alpha-s)}\exp\left\{-2(T-t)\bar{\lambda}\bar{\eta}\right\}$
@@
 Since $\bar{\eta}=\eta_{1}T^{-\theta_{3}}$, and applying Proposition 5.10 with $v=1+2\alpha-s$, we obtain
-$\mathcal{T}_{4}\leq\widetilde{c_{4}}\begin{cases}T^{-(1-2\alpha+s)\theta_{3}+(s-2\alpha)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }2\alpha<s\leq 1+2\alpha,\\ T^{-\theta_{3}}\log T,&\text{ when }2\alpha=s,\\ T^{-\theta_{3}},&\text{ when }2\alpha>s,\end{cases}$
+[…]
 where $\widetilde{c_{4}}:=12(\kappa^{2(1+2\alpha-s)}+(\frac{1+2\alpha-s}{2e})^{1+2\alpha-s})\sqrt{c}\left(\sqrt{c}\widetilde{M}+\sigma^{2}\right)\mathrm{Tr}(C^{s})\eta_{1}^{2}\delta_{3}$ is a constant independent of $T$.
@@
 Define a function $f:\{1,2,\cdots\}\rightarrow\mathbb{R}$ iteratively, as $f(1):=\kappa^{2}\left\|H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ and,
-$f(t+1):=$ $\mathcal{T}_{1}(t+1)+\mathcal{T}_{2}(t+1)+\mathcal{T}_{3}(t+1)$
+$f(t+1)$= $\mathcal{T}_{1}(t+1)+\mathcal{T}_{2}(t+1)+\mathcal{T}_{3}(t+1)$
 $+6\sqrt{c}\sum_{k=1}^{t}\eta_{k}^{2}\left(\sqrt{c}f(k)+\sigma^{2}\right)\mathrm{Tr}\left(C^{2}\prod_{j=k+1}^{t}(I-\eta_{j}(C+\lambda_{j}I))^{2}\right)$
@@
 $+c_{3}(T+t_{0})^{-2\theta_{2}\min\{r,1\}+\theta_{1}-1}$
-$+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)\begin{cases}(T+t_{0})^{-\theta_{1}},&\text{ when }s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0}),&\text{ when }s=1.\end{cases}$
+[…]
 We choose $\theta_{1}=\frac{2\min\{r+1/2,1\}}{1+2\min\{r+1/2,1\}}$ and $\theta_{2}=\frac{1}{1+2\min\{r+1/2,1\}}$, then
@@
 for any $T\geq t_{0}+1$, where the constant $c_{1,1}=c_{1}\bar{\lambda}^{\min\{2r+1,2\}}+c_{2}\bar{\eta}^{-(2r+1)}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}+c_{3}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)$ is independent of $T$.
-Let $c_{1,1}$ be sufficiently large such that (6.1) holds true for $11\leq T<t_{0}+$ 1.
+Let $c_{1,1}$ be sufficiently large such that (6.1) holds true for $1\leq T<t_{0}+1$.
 We then finish the proof. ∎
@@
 for any $T\geq t_{0}+1$, where $c_{1,2}=c_{1}\bar{\lambda}^{2\min\{r,1\}}+c_{2}\bar{\eta}^{-2r}(t_{0}+1)^{2\bar{\eta}\bar{\lambda}}+c_{3}+c_{4}\left(\sqrt{c}M+\sigma^{2}\right)$ is a constant independent of $T$.
-Let $c_{1,2}$ be sufficiently large such that (6.2) holds true for $11\leq T<t_{0}+$ 1.
+Let $c_{1,2}$ be sufficiently large such that (6.2) holds true for $1\leq T<t_{0}+1$.
 The proof is complete. ∎
@@
 $+c_{2}\eta_{1}^{-(2r+1)}T^{-(2r+1)(1-\theta_{3})}\exp\{-\tau\eta_{1}\lambda_{1}T^{1-\theta_{4}-\theta_{3}}\}$
-$+\widetilde{c_{4}}\begin{cases}T^{-\theta_{3}}\log T,&\text{ when}s=1,\\ T^{-\theta_{3}},&\text{ when }s<1.\end{cases}$
+[…]
 We choose $\theta_{3}=\frac{2r+1}{2r+2}$ and $\theta_{4}\geq\frac{2r+1}{(2r+2)\min\{2r+1,2\}}$, then
@@
 Our proofs are mainly based on the following proposition.
-This proposition is from [43, Proposition A.3], and is an extension of [39, Theorem 3.4].
+This proposition is from [40, Proposition A.3], and is an extension of [36, Theorem 3.4].
 Proposition 7.1.
@@
     $\leq$ $2\eta_{t}\kappa\left(M_{\rho}+2\kappa\|H^{\dagger}\|_{\mathrm{HS}}+\kappa d_{2}(t+t_{0})^{\frac{1}{2}-\theta_{1}}\log(t+t_{0})\log\frac{2}{\delta}\right)$
-    ×$\times\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|.$
+    $\times\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\right\|.$
     If $\alpha>0$, by Lemma 5.2 (1) and Lemma 5.3, we obtain
@@
     $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $\sum_{t=1}^{T}\eta_{t}^{2}\left(2M_{\rho}^{2}+2\kappa^{2}\left(8\|H^{\dagger}\|_{\mathrm{HS}}^{2}+2d_{2}^{2}(t+t_{0})^{1-2\theta_{1}}\log^{2}(t+t_{0})\log^{2}\frac{2}{\delta}\right)\right)$ (7.8)
-    ×$\times\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}.$
-    By the definition of the trace of operators and using Assumption 3 and (3) in Lemma 5.2, it follows that if $02\alpha+1-s>$ 0,
+    $\times\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}.$
+    By the definition of the trace of operators and using Assumption 3 and (3) in Lemma 5.2, it follows that if $2\alpha+1-s>0$,
     $\quad\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}$ (7.9)
@@
     $\quad\leq\mathrm{Tr}\left(C^{s}\right)\frac{2(\kappa^{2(2\alpha+1-s))}+((2\alpha+1-s)/(2e))^{2\alpha+1-s})}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{2\alpha+1-s}}$
-    ×$\quad\quad\times\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\},$
-    else if $02\alpha+1-s=$ 0,
+    $\quad\quad\times\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\},$
+    else if $2\alpha+1-s=0$,
     $\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}\leq\mathrm{Tr}\left(C^{s}\right)\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}.$ (7.10)
@@
     $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $d_{4}\sum_{t=1}^{T}(t+t_{0})^{-2\theta_{1}}\left(1+(t+t_{0})^{1-2\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)$ (7.11)
-    ×$\times\frac{\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{2\alpha+1-s}}.$
+    $\times\frac{\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{2\alpha+1-s}}.$
     Now, we apply Proposition 5.7 and Remark 3 with $v=2\alpha+1-s$ to derive the bound. Let $d_{5}$ denote a constant independent of $T$ and $\delta$.
-    Case 1: If $12\alpha+1-s>$ 1, then
+    Case 1: If $2\alpha+1-s>1$, then
     $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq d_{5}(T+t_{0})^{-\theta_{1}}+d_{5}(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}.$
-    Case 2: If $12\alpha+1-s=$ 1, then
+    Case 2: If $2\alpha+1-s=1$, then
     $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq d_{5}(T+t_{0})^{-\theta_{1}}\log(T+t_{0})+d_{5}(T+t_{0})^{1-3\theta_{1}}\log^{3}(T+t_{0})\log^{2}\frac{2}{\delta}.$
-    Case 3: If $10<2\alpha+1-s<$ 1, then
+    Case 3: If $0<2\alpha+1-s<1$, then
     $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq d_{5}(T+t_{0})^{s-2\alpha-\theta_{1}(1+s-2\alpha)}+d_{5}(T+t_{0})^{1+s-2\alpha-\theta_{1}(3+s-2\alpha)}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}.$
-    Case 4: If $02\alpha+1-s=$ 0. Since
+    Case 4: If $2\alpha+1-s=0$. Since
     $\exp\left\{-2\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}$ $=\exp\{-2\bar{\eta}\bar{\lambda}\sum_{j=t+1}^{T}(j+t_{0})^{-1}\}$
@@
 and
-$\tau_{1}^{2}$ $=d_{5}\begin{cases}(T+t_{0})^{-\theta_{1}}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta},&\text{ when }s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0})+(T+t_{0})^{1-3\theta_{1}}\log^{3}(T+t_{0})\log^{2}\frac{2}{\delta},&\text{ when }s=1.\\ \end{cases}$
+[…]
 Therefore,
 $\left\|\sum_{i=1}^{T}\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $8(2d_{3}^{2}+d_{5})\left((T+t_{0})^{-\theta_{1}}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)$ (7.17)
-×$\times\log^{2}\frac{2}{\delta}\begin{cases}1,&\text{ when }s<1,\\ \log(T+t_{0}),&\text{ when }s=1.\end{cases}$
+[…]
 If $T\geq t_{0}+1$, combining (7.12), (7.13), (7.14), (7.15), (7.16), and (7.17), we obtain that there exists some constant $c_{2,1}$ independent of $T$ and $\delta$, such that
-$\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\leq c_{2,1}\begin{cases}(T+t_{0})^{-\theta_{1}}\log^{2}\frac{2}{\delta}+(T+t_{0})^{1-3\theta_{1}}\log^{2}(T+t_{0})\log^{4}\frac{2}{\delta},&s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0})\log^{2}\frac{2}{\delta}+(T+t_{0})^{1-3\theta_{1}}\log^{3}(T+t_{0})\log^{4}\frac{2}{\delta},&s=1.\\ \end{cases}$
+[…]
 holds with probability at least $1-2\delta$.
-Since it is easy to verify that $\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}$ for $11\leq T<t_{0}+$ 1 can be bounded uniformly by a constant, we can choose $c_{2,1}$ to be sufficiently large such that the bound holds true for $11\leq T<t_{0}+$ 1.
+Since it is easy to verify that $\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}$ for $1\leq T<t_{0}+1$ can be bounded uniformly by a constant, we can choose $c_{2,1}$ to be sufficiently large such that the bound holds true for $1\leq T<t_{0}+1$.
 The proof is complete. ∎
@@
 $\left\|\sum_{i=1}^{T}\widetilde{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $8\left(2d_{3}^{2}+d_{5}\right)\left((T+t_{0})^{s-\theta_{1}(1+s)}+(T+t_{0})^{1+s-\theta_{1}(3+s)}\right$. (7.23)
-×$\left.\times\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}.$
+$\left.\times\log^{2}(T+t_{0})\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}.$
 Combining (7.18), (7.19), (7.20), (7.21), (7.22), and (7.23), let $c_{2,2}$ denote a constant independent of $T$ and $\delta$.
@@
   • (2)
-    If $12\min\{r,1\}+s<$ 1, i.e., $12r+s<$ 1, choose $\theta_{1}=\frac{1+2\min\{r,1\}+s}{3+2\min\{r,1\}+s}=\frac{1+2r+s}{3+2r+s}$ and $\theta_{2}=\frac{2}{3+2\min\{r,1\}+s}=\frac{2}{3+2r+s}$. Then,
+    If $2\min\{r,1\}+s<1$, i.e., $2r+s<1$, choose $\theta_{1}=\frac{1+2\min\{r,1\}+s}{3+2\min\{r,1\}+s}=\frac{1+2r+s}{3+2r+s}$ and $\theta_{2}=\frac{2}{3+2\min\{r,1\}+s}=\frac{2}{3+2r+s}$. Then,
     $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ $\leq c_{2,2}(T+t_{0})^{-\frac{4\min\{r,1\}}{3+2\min\{r,1\}+s}}\log^{2}(T+t_{0})\log^{4}\frac{2}{\delta}$
@@
     holds with probability at least $1-2\delta$.
-Since $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ for $11\leq T<t_{0}+$ 1 can be uniformly bounded by a constant, we can choose $c_{2,2}$ to be sufficiently large so that the bound also holds in this case.
+Since $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ for $1\leq T<t_{0}+1$ can be uniformly bounded by a constant, we can choose $c_{2,2}$ to be sufficiently large so that the bound also holds in this case.
 The proof is then complete. ∎
@@
 Suppose Assumption 3 and 5 hold.
-Let $\alpha\in[0,\frac{1}{2}]$, $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ with $10<\theta_{3}<$ 1, and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ with $\theta_{4}>0$.
+Let $\alpha\in[0,\frac{1}{2}]$, $t_{0}=\theta_{1}=\theta_{2}=0$, $\eta_{t}\equiv\bar{\eta}=\eta_{1}T^{-\theta_{3}}$ with $0<\theta_{3}<1$, and $\lambda_{t}\equiv\bar{\lambda}=\lambda_{1}T^{-\theta_{4}}$ with $\theta_{4}>0$.
 Assume $T\geq 2$ and $\eta_{1}(\kappa^{2}+\lambda_{1})\leq 1$.
@@
     $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $\sum_{t=1}^{T}\eta_{t}^{2}\left(2M_{\rho}^{2}+2\kappa^{2}\left(8\|H^{\dagger}\|_{\mathrm{HS}}^{2}+2\left(\frac{2d_{6}}{\log 2}\right)^{2}T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\right)$ (7.25)
-    ×$\times\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}.$
-    By the definition of the trace of operators, Assumption 3, and Lemma 5.2 (3), if $02\alpha+1-s>$ 0, then
+    $\times\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}.$
+    By the definition of the trace of operators, Assumption 3, and Lemma 5.2 (3), if $2\alpha+1-s>0$, then
     $\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}$ (7.26)
@@
     $\leq$ $\mathrm{Tr}\left(C^{s}\right)\frac{2(\kappa^{2(2\alpha+1-s))}+((2\alpha+1-s)/(2e))^{2\alpha+1-s})}{1+\left((T-t)\bar{\eta}\right)^{2\alpha+1-s}}$
-    ×$\times\exp\left\{-2(T-t)\bar{\eta}\bar{\lambda}\right\},$
-    and if $02\alpha+1-s=$ 0, then
+    $\times\exp\left\{-2(T-t)\bar{\eta}\bar{\lambda}\right\},$
+    and if $2\alpha+1-s=0$, then
     $\mathbb{E}_{z_{t}}\left\|C^{\alpha}\prod_{j=t+1}^{T}(I-\eta_{j}(C+\lambda_{j}I))\phi(x_{t})\right\|_{\mathcal{H}_{\mathcal{K}}}^{2}\leq\mathrm{Tr}\left(C^{s}\right)\exp\left\{-2(T-t)\bar{\eta}\bar{\lambda}\right\}.$ (7.27)
@@
     $\sum_{t=1}^{T}\mathbb{E}_{z_{t}}\left\|\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $d_{8}\delta_{3}T^{-2\theta_{3}}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)$
-    ×$\times\begin{cases}T^{(1+2\alpha-s)\theta_{3}+(s-2\alpha)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }2\alpha<s\leq 1+2\alpha,\\ T^{\theta_{3}}\log T,&\text{ when }2\alpha=s,\\ T^{\theta_{3}},&\text{ when }2\alpha>s,\end{cases}$
+[…]
     $=$ $d_{8}\delta_{3}\left(1+T^{1-2\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)$
-    ×$\times\begin{cases}T^{-(1-2\alpha+s)\theta_{3}+(s-2\alpha)\min\{1,\theta_{3}+\theta_{4}\}},&\text{ when }2\alpha<s\leq 1+2\alpha,\\ T^{-\theta_{3}}\log T,&\text{ when }2\alpha=s,\\ T^{-\theta_{3}},&\text{ when }2\alpha>s.\end{cases}$
+[…]
 The proof is complete. ∎
@@
 and
-$\tau_{2}^{2}=d_{8}\delta_{3}\left(T^{-\theta_{3}}+T^{1-3\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\begin{cases}\log T,&\text{ when }s=1,\\ 1,&\text{ when }s<1.\end{cases}$
+[…]
 Therefore,
-$\left\|\sum_{i=1}^{T}\overline{\chi}_{t}\right\|_{\mathrm{HS}}^{2}\leq$ $8(d_{7}+d_{8}\delta_{3})\left(T^{-\theta_{3}}+T^{1-3\theta_{3}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}\begin{cases}\log T,&\text{ when }s=1,\\ 1,&\text{ when }s<1.\end{cases}$ (7.32)
+[…]
 Combining (7.28), (7.29), (7.30), (7.31), and (7.32), we conclude that there exists a constant $c_{2,3}$ independent of $T$, such that
-$\left\|\left(H_{T+1}-H^{\dagger}\right)C^{\frac{1}{2}}\right\|_{\mathrm{HS}}^{2}\leq c_{2,3}\begin{cases}T^{-\theta_{3}}\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{2}T\log^{4}\frac{2}{\delta},&\text{ when }s<1,\\ T^{-\theta_{3}}\log T\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{3}T\log^{4}\frac{2}{\delta},&\text{ when }s=1,\\ \end{cases}$
+[…]
 holds with probability at least $1-2\delta$.
@@
   • (2)
-    If $12r+s<$ 1, choose $\theta_{3}=\frac{1+2r+s}{3+2r+s}$ and $\theta_{4}\geq\frac{2r}{(3+2r+s)\min\{r,1\}}=\frac{2r}{(3+2r+s)r}$, then
+    If $2r+s<1$, choose $\theta_{3}=\frac{1+2r+s}{3+2r+s}$ and $\theta_{4}\geq\frac{2r}{(3+2r+s)\min\{r,1\}}=\frac{2r}{(3+2r+s)r}$, then
     $\left\|H_{T+1}-H^{\dagger}\right\|_{\mathrm{HS}}^{2}$ $\leq c_{2,4}\left(T^{-\frac{1+2r-s}{3+2r+s}}+T^{-\frac{4r}{3+2r+s}}\log^{2}T\log^{2}\frac{2}{\delta}\right)\log^{2}\frac{2}{\delta}$
@@
 Then, it is clear that $\mathcal{H}_{K}=\overline{\mathcal{H}_{0}}$ and $\mathcal{B}_{\mathrm{HS}}(\mathcal{H}_{\mathcal{K}},\overline{W^{1/2}\mathcal{Y}})=\overline{\mathcal{B}_{0}}$.
-We define the mapping $W_{0}:\mathcal{H}_{0}\rightarrow\mathcal{B}_{0}$ by $\ \sum_{i=1}^{n}\alpha_{i}K(x_{i},\cdot)y_{i}\mapsto\sum_{i=1}^{n}\alpha_{i}W^{1/2}y_{i}\otimes\phi(x_{i})$ for any $n\in\mathbb{N}$, $x_{1},\cdots x_{n}\in\mathcal{X}$ and $y_{1},\cdots y_{n}\in\mathcal{Y}$.
-We see that $W_{0}$ is well-defined and linear.
-Moreover, for any $\sum_{i=1}^{n}\alpha_{i}K(x_{i},\cdot)y_{i}=\sum_{i=1}^{n}\alpha_{i}\mathcal{K}(x_{i},\cdot)Wy_{i}\in\mathcal{H}_{0}$, we have
+[…]
 $\left\|W_{0}\left(\sum_{i=1}^{n}\alpha_{i}K(x_{i},\cdot)y_{i}\right)\right\|_{\mathrm{HS}}^{2}$ $=\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}\langle W^{1/2}y_{i}\otimes\phi(x_{i}),W^{1/2}y_{j}\otimes\phi(x_{j})\rangle_{\mathrm{HS}}$
@@
 $\sum_{t=1}^{T/2}\frac{(t+t_{0})^{p}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}$ $\leq\frac{\left(1-(3/4)^{1-\theta_{1}}\right)^{-v}}{\min\{1,{\bar{\eta}}^{v}\}}\sum_{t=1}^{T/2}\frac{(t+t_{0})^{p}}{(T+t_{0})^{(1-\theta_{1})v}}$
+$\leq\frac{\left(1-(3/4)^{1-\theta_{1}}\right)^{-v}}{\min\{1,{\bar{\eta}}^{v}\}}(T+t_{0})^{-(1-\theta_{1})v}\int_{0}^{\frac{T}{2}+1}(x+t_{0})^{p}dx$
 […]
-$\leq\frac{\left(1-(3/4)^{1-\theta_{1}}\right)^{-v}}{\min\{1,{\bar{\eta}}^{v}\}}(T+t_{0})^{-(1-\theta_{1})v}\begin{cases}\frac{1}{p+1}(T+t_{0})^{p+1},&\text{ when }p>-1,\\ \log(T+t_{0}),&\text{ when }p=-1,\\ \frac{t_{0}^{p+1}}{-1-p}\leq\frac{1}{-1-p},&\text{ when }p<-1,\end{cases}$
-$\leq\delta^{\prime}\begin{cases}(T+t_{0})^{-(1-\theta_{1})v+p+1},&\text{ when }p>-1,\\ (T+t_{0})^{-(1-\theta_{1})v}\log(T+t_{0}),&\text{ when }p=-1,\\ (T+t_{0})^{-(1-\theta_{1})v},&\text{ when }p<-1,\end{cases}$
+[…]
 where
@@
 Next, we bound $\sum_{t=T/2}^{T-1}\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}$. If $p>0$, it holds that
-[…]
+$\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}\leq\int_{t+1}^{t+2}\frac{(u+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(u+t_{0})^{1-\theta_{1}}\right]^{v}}du.$
 If $p\leq 0$, there holds $(t+t_{0})^{p}\leq 3^{-p}(t+t_{0}+2)^{p}$. Thus we have
-[…]
+$\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}\leq 3^{-p}\int_{t+1}^{t+2}\frac{(u+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(u+t_{0})^{1-\theta_{1}}\right]^{v}}du.$
 Therefore,
 $\sum_{t=T/2}^{T-1}\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}$
-[…]
+$\leq\max\{3^{-p},1\}\int_{T/2+1}^{T+1}\frac{(u+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(u+t_{0})^{1-\theta_{1}}\right]^{v}}du.$
 Let $\xi=(T+t_{0}+1)^{1-\theta_{1}}-(u+t_{0})^{1-\theta_{1}}$, then $d\xi=-(1-\theta_{1})(u+t_{0})^{-\theta_{1}}du$, then
 $\sum_{t=T/2}^{T-1}$ $\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}$ (A.7)
-[…]
-[…]
+$\leq\frac{\max\{3^{-p},1\}}{1-\theta_{1}}\int_{0}^{(T+t_{0}+1)^{1-\theta_{1}}-(T/2+t_{0}+1)^{1-\theta_{1}}}\frac{(u+t_{0})^{p+\theta_{1}}}{1+\xi^{v}}d\xi$
+$\leq\frac{\max\{3^{-p},1\}}{1-\theta_{1}}\int_{0}^{(T+t_{0}+1)^{1-\theta_{1}}(1-2^{\theta_{1}-1})}\frac{(u+t_{0})^{p+\theta_{1}}}{1+\xi^{v}}d\xi.$
 Since $u+t_{0}\in\left[T/2+t_{0}+1,T+t_{0}+1\right]$, whenever $p+\theta_{1}>0$ or not, we have
@@
 $\sum_{t=T/2}^{T-1}$ $\frac{(t+t_{0})^{p}}{1+\left[(T+t_{0}+1)^{1-\theta_{1}}-(t+t_{0}+1)^{1-\theta_{1}}\right]^{v}}$ (A.9)
+$\leq\frac{\max\{3^{-p},1\}}{1-\theta_{1}}2^{\lvert p+\theta_{1}\rvert}(T+t_{0})^{p+\theta_{1}}\left(1+\int_{1}^{(T+t_{0}+1)^{1-\theta_{1}}(1-2^{\theta_{1}-1})}\frac{1}{\xi^{v}}d\xi\right)$
 […]
-$\leq\frac{\max\{3^{-p},1\}}{1-\theta_{1}}2^{\lvert p+\theta_{1}\rvert}(T+t_{0})^{p+\theta_{1}}\begin{cases}\frac{v}{v-1},&\text{ when }v>1,\\ \left(2-\theta_{1}\right)\log(T+t_{0}+1),&\text{ when }v=1,\\ \frac{\left(1-2^{\theta_{1}-1}\right)^{1-v}}{1-v}(T+t_{0}+1)^{(1-\theta_{1})(1-v)},&\text{ when }v<1,\end{cases}$
-$=\widetilde{\delta^{\prime\prime}}\begin{cases}(T+t_{0})^{p+\theta_{1}},&\text{ when }v>1,\\ (T+t_{0})^{p+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{p+1-v(1-\theta_{1})},&\text{ when }v<1,\end{cases}$
+[…]
 with some constant $\widetilde{\delta^{\prime\prime}}$ independent of $T$ and $t_{0}$.
@@
 $\sum_{t=1}^{T}\exp\left\{-\sum_{j=t+1}^{T}\eta_{j}\lambda_{j}\right\}\frac{(t+t_{0})^{-\theta}}{1+\left(\sum_{j=t+1}^{T}\eta_{j}\right)^{v}}$ (A.11)
-$\leq$ $2^{\bar{\eta}\bar{\lambda}}(\delta^{\prime}+\delta^{\prime\prime})\begin{cases}(T+t_{0})^{-\theta+\theta_{1}},&\text{ when }v>1,\\ (T+t_{0})^{-\theta+\theta_{1}}\log(T+t_{0}),&\text{ when }v=1,\\ (T+t_{0})^{-\theta+1-v(1-\theta_{1})},&\text{ when }v<1.\end{cases}$
+[…]
 We finish the proof by setting $\delta_{1}=2^{\bar{\eta}\bar{\lambda}}(\delta^{\prime}+\delta^{\prime\prime})$, which is independent of $T$ and $t_{0}$. ∎